### v3

# Dimitrios M. Thilikos ${}^{,}$

Abstract
We introduce an annotated extension of treewidth that measures the contribution of a vertex set $X$ to the treewidth of a graph $G.$ This notion provides a graph distance measure to some graph property $\mathcal{P}$: A vertex set $X$ is a $k$-treewidth modulator of $G$ to $\mathcal{P}$ if the treewidth of $X$ in $G$ is at most $k$ and its removal gives a graph in $\mathcal{P}.$ This notion allows for a version of the Graph Minors Structure Theorem (GMST) that has no need for apices and vortices: $K_{k}$-minor free graphs are those that admit tree-decompositions whose torsos have $c_{k}$-treewidth modulators to some surface of Euler-genus $c_{k}.$ This reveals that minor-exclusion is essentially tree-decomposability to a “modulator-target scheme” where the modulator is measured by its treewidth and the target is surface embeddability.
We then fix the target condition by demanding that $\Sigma$ is some particular surface and define a “surface extension” of treewidth, where $\Sigma\mbox{-}{\sf tw}(G)$ is the minimum $k$ for which $G$ admits a tree-decomposition whose torsos have a $k$-treewidth modulator to being embeddable in $\Sigma.$ We identify a finite collection $\mathfrak{D}_{\Sigma}$ of parametric graphs and prove that the minor-exclusion of the graphs in $\mathfrak{D}_{\Sigma}$ precisely determines the asymptotic behavior of ${\Sigma}\mbox{-}{\sf tw},$ for every surface $\Sigma.$ It follows that the collection $\mathfrak{D}_{\Sigma}$ bijectively corresponds to the “surface obstructions” for $\Sigma,$ i.e., surfaces that are minimally non-contained in $\Sigma.$
Keywords: Treewidth, Graph minor, Surface embedding, Graph modulator, Graph Genus, Parametric graph, Universal obstruction, Graph parameter, Graph structure theorem.
1 Introduction
An underlying fact in many results in structural graph theory is that graphs that do not “contain” some graph can be seen as the result of gluing together simpler graphs of some particular property in a tree-like fashion. When the containment relation is the minor relation [^1] We say that a graph $H$ is a minor of a graph $G$ if $H$ can be obtained from some subgraph of $G$ by contracting edges. this property is typically enjoying certain topological characteristics. Moreover, in some cases, such structural theorems may serve as the combinatorial base for algorithmic applications.
Two celebrated, and highly related, theorems of this type are the Grid Theorem and the Graph Minors Structure Theorem (GMST), proven by Robertson and Seymour in [39] and  [42] respectively.
The Grid Theorem asserts that every graph which cannot be obtained by gluing together small graphs in a tree-like fashion must contain all planar graphs up to a certain size as minors. Vice-versa, if some graph $G$ does not contain a fixed planar graph $H$ as a minor, $G$ can indeed be obtained by gluing together small graphs in a tree-like fashion where the notion of “small” now depends on the excluded graph $H.$ Taking this insight as the base case, the GMST gives the most general version of similar statement. Here, the excluded minor is no longer planar, but a clique of fixed size. Any graph excluding such a large clique minor can be obtained by gluing together graphs that “almost” embed in some surface of bounded Euler-genus in, again, a tree-like fashion. This is a property that is not enjoyed by large complete graphs.
The name “Grid Theorem” stems from the way Robertson and Seymour prove the above result: They identify a single parametric graph, namely the $(n\times n)$-grid, and show that
  1. 1.
    every planar graph is a minor of the $(n\times n)$-grid for some $n\in\mathbb{N},$
  2. 2.
    the treewidth of the $(n\times n)$-grid depends on $n,$ and
  3. 3.
    there exists a function $f$ such that for every $n\in\mathbb{N},$ every graph with treewidth more than $f(n)$ contains the $(n\times n)$-grid as a minor.
So, the grid fulfils two roles: It acts as a universal obstruction for the parameter treewidth and it is an omnivore for the class of planar graphs; that is, the class of all minors of some grid is exactly the class of planar graphs. This means that embeddability on the sphere is tightly captured by the decomposability into small pieces, defined by the parameter of treewidth.
This paper is motivated by the following question:
1.1 Our contribution
Our strategy to resolve the motivating question above follows the three ingredients given for treewidth. We identify families of grid-like graphs, we call them Dyck-grids (or surface-grids if we are talking about their less streamlined version). Each Dyck-grid is a parametric graph which is an omnivore for all graphs that embed in the respective surface. We then proceed by proving that the minor exclusion of the Dyck-grids corresponding to some set of surfaces $\mathbb{S}$ leads to a “surface” generalization of treewidth, namely ${\mathbb{S}}\mbox{-}{\sf tw},$ defined via tree-decompositions where all torsos “almost embed” into some surface in $\mathbb{S}.$ Finally, we prove that, with respect to our notion of “almost embedability”, the Dyck-grids themselves cannot be decomposed in this way. With this, we obtain an asymptotic characterization of the resulting width parameters ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}}$ in form of the exclusion of parametric graphs.
Measuring modification by treewidth.
To capture what we mean by “almost embeddable” we use again treewidth. In particular, we introduce an annotated extension of treewidth as follows.
Given a graph $H$ and a set $X\subseteq V(G),$ we say that $H$ is an $X$-rooted minor of $G$ (or, simply, an $X$-minor of $G$) if there is a collection $\mathcal{S}=\{S_{v}\mid v\in V(H)\}$ of pairwise vertex-disjoint connected [^2] A set $X\subseteq V(G)$ is connected in $G$ if the induced subgraph $G[X]$ is a connected graph. subsets of $V(G),$ each containing at least one vertex of $X$ and such that, for every edge $xy\in E(H),$ the set $S_{x}\cup S_{y}$ is connected in $G.$ $H$ is a minor of $G,$ denoted by $H\leq G,$ if $H$ is a $V(G)$-minor of $G.$
Given an annotated graph $(G,X)$ where $G$ is a graph and $X\subseteq V(G),$ we define ${\sf tw}(G,X)$ as the maximum treewidth of an $X$-minor of $G.$
Notice that the annotated extension of treewidth above can be defined for every parameter (see (12) in Section 7). In the case of treewidth, this gives a measure of the contribution of $X$ to the tree decomposability of $G.$ We say that an $X\subseteq V(G)$ is a $k$-treewidth modulator of $G$ to some graph class $\mathcal{G}$ if ${\sf tw}(G,X)\leq k$ and $G-X\in\mathcal{G}.$ That way, treewidth may serve as the measure of the amount of modification caused by the removal of a modulator to some target graph class. In the case of our results, the target graph class will be the class $\mathcal{E}_{g}$ containing all graphs of Euler-genus at most $g.$
As a first step, we display how the GMST can be restated using the modulator-target duality above. For this, we consider the Hadwiger number of a graph, that is the graph parameter ${\sf hw},$ where ${\sf hw}(G)$ is the largest size of a clique minor of $G.$ We define the parameter ${\sf hw}^{\prime}$ as follows
The clique-sum closure, defined formally in Section 2, formalizes the “tree-decomposability” notion that underlines structural theorems such as the Grid Theorem and the GMST.
In this paper we provide a variant of the GMST stating that ${\sf hw}$ and ${\sf hw}^{\prime}$ are asymptotically equivalent (Theorem 8.2). That is, there exists a function $f\colon\mathbb{N}\to\mathbb{N}$ such that for all graphs $G$ we have ${\sf hw}(G)\leq f({\sf hw}^{\prime}(G))$ and ${\sf hw}^{\prime}(G)\leq f({\sf hw}(G)).$ In what follows, we denote this equivalence relation between graph parameters using the symbol $\sim$,∼ , so our result may be stated as ${\sf hw}\sim{\sf hw}^{\prime}.$ This indicates that $K_{k}$-minor-free graphs are those that are tree-decomposable into pieces which are “treewidth close” to graphs of small Euler-genus. Moreover, it restates the GMST by substituting the presence of apices and vortices [^3] See Section 3 for a formal definition of apices and vortices. with a graph modification measure: the removal of a “low-treewidth” modulator.
Dyck-grids.
Observe that the definition of ${\sf hw}^{\prime}$ in (2) is simultaneously minimizing two quantities: the treewidth of the modulator and the Euler-genus of the remaining graph. The min-max duality of the GMST is linking this entity with the exclusion of a parametric graph as a minor. In this case, the parametric graph is the complete graph on $k$ vertices.
But what about fixing the Euler-genus? What should replace cliques as the “excluded entity” to provide a min-max duality for some particular surface? In the “bottom” case of such a theorem we should position the case where the target condition is void. In this bottom case the resulting min-max theorem is the Grid Theorem asserting the equivalence between treewidth and the exclusion of grids (expressing planarity).
Figure 1: The annulus grid $\mathscr{A}_{9},$ the handle grid $\mathscr{H}_{9}$ and the crosscap grid $\mathscr{C}_{9}$ in order from left to right. Notice that both $\mathscr{H}_{9}$ and $\mathscr{C}_{9}$ contain two $(18\times 9)$ grids as vertex-disjoint subgraphs (depicted in different colors). Moreover, $\mathscr{C}_{9}$ contains a $(18\times 18)$-grid as a spanning subgraph.
To present our main result and give an answer to the question above, we define three parametric graphs $\mathscr{A}=\langle\mathscr{A}_{k}\rangle_{k\in\mathbb{N}},$ $\mathscr{H}=\langle\mathscr{H}_{k}\rangle_{k\in\mathbb{N}},$ and $\mathscr{C}=\langle\mathscr{C}_{k}\rangle_{k\in\mathbb{N}}$ as follows:
The annulus grid $\mathscr{A}_{k}$ is the $(4k,k)$-cylindrical grid [^4] An $(n\times m)$-cylindical grid is a Cartesian product of a cycle on $n$ vertices and a path on $m$ vertices. depicted in the left of Figure 1. The handle grid $\mathscr{H}_{k}$ (resp. crosscap grid $\mathscr{C}_{k}$) is obtained adding in $\mathscr{A}_{k}$ edges as indicated in the middle (resp. right) part of Figure 1. We refer to the added edges as transactions of the handle grid $\mathscr{H}_{k}$ or the crosscap grid $\mathscr{C}_{k}.$
Let now $h\in\mathbb{N}$ and $c\in[0,2].$ We define the parametric graph $\mathscr{D}^{h,c}=\langle\mathscr{D}_{k}^{h,c}\rangle_{k\in\mathbb{N}}$ by taking one copy of $\mathscr{A}_{k},$ $h$ copies of $\mathscr{H}_{k},$ and $c\in[0,2]$ copies of $\mathscr{C}_{k},$ then “cut” them along the dotted red line, as in Figure 1, and join them together in the cyclic order $\mathscr{A}_{k},\mathscr{H}_{k},\ldots,\mathscr{H}_{k},\mathscr{C}_{k},\ldots,\mathscr{C}_{k},$ as indicated in Figure 2. We call the graph $\mathscr{D}_{k}^{(h,c)}$ the Dyck-grid of order $k$ with $h$ handles and $c$ crosscaps. For technical reasons, we make the convention that $\mathscr{D}^{(-1,2)}_{k}=\mathscr{D}^{(0,0)}_{k}.$
Let $g\in\mathbb{N}\cup\{-1\}$ be an integer. If $g\geq 0,$ we define $\mathcal{E}_{g}$ as the class of graphs that are embeddable in some surface $\Sigma$ of Euler-genus at most $g.$ We also set $\mathcal{E}_{-1}$ to be the class containing only the empty graph and we use $\mathcal{G}_{\text{all}}$ to denote the class of all graphs. We define the graph parameter $g\mbox{-}{\sf tw}\colon\mathcal{G_{\text{all}}}\to\mathbb{N}$ as follows
Figure 2: The Dyck-grid of order $8$ with one handle and two crosscaps, i.e., the graph $\mathscr{D}_{8}^{1,2}.$
We stress that the only difference between (2) and (3) is that, in ($,$ the target property is fixed to $\mathcal{E}_{g}.$ Notice that $g\mbox{-}{\sf tw}(G)$ can be seen as a parametric extension of surface embeddability in the sense that if $g\mbox{-}{\sf tw}(G)\leq k$, then $G$ can be decomposed to graphs that are “$k$-treewidth close” to being embeddable in a surface of Euler-genus $g$. Observe that, for every graph $G,$ ${\sf tw}(G,V(G))={\sf tw}(G).$ Therefore, in the base case where $g=-1,$ we have that $(-1)\mbox{-}{\sf tw}={\sf tw}$.
Our next step is to define the graph parameter ${g}\mbox{-}\mathsf{bg}$ for $g\in\mathbb{N}$ as follows
${g}\mbox{-}\mathsf{bg}(G)\coloneqq\begin{cases}\max\{k\mid\mathscr{D}^{(0,0)}_{k}\leq G\}&\mbox{if $g=0$}\\ \max\{k\mid\mathscr{D}^{(h+1,0)}_{k}\leq G\mbox{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\mathscr{D}^{(h,1)}_{k}\leq G\}&\mbox{if $g=2h+1\geq 1$}\\ \max\{k\mid\mathscr{D}^{(h+1,0)}_{k}\leq G\mbox{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\mathscr{D}^{(h,2)}_{k}\leq G\}&\mbox{if $g=2h+2\geq 2$}\end{cases}$ (4)
Notice that, in case $g=0$, ${g}\mbox{-}\mathsf{bg}(G)$ is the maximum $k$ for which $G$ contains the annulus $\mathscr{A}_{k}$ as a minor. This is asymptotically equivalent to the size of the biggest grid minor of $G$. This motivated the choice of $\mathsf{bg}$ for the name of this parameter.
We are now ready to state our main result that is the following min-max equivalence between the parameters defined in (3) and (4).
Theorem 1.1.
For every $g\in\mathbb{N}\cup\{-1\},$ the parameters ${g}\mbox{-}\mathsf{tw}$ and ${(g+1)}\mbox{-}\mathsf{bg}$ are asymptotically equivalent. That is, ${g}\mbox{-}\mathsf{tw}\sim{(g+1)}\mbox{-}\mathsf{bg}.$
One important contribution of Theorem 1.1 to the general theory of graph minors is, that it makes the dependencies on the size $|V(H)|$ of the excluded minor $H$, and its Euler-genus visible and utilizable. As a consequence of our combinatorial results we are able to decuce a number of improvements on deep algorithmic results in the realm of $H$-minor-free graphs. In particular, Theorem 1.1 allows to replace the dependency on the size of $H$ in the multiplicative constants of the linear grid theorem in $H$-minor-free graphs [29], the constant-factor polynomial-time approximation algorithm for treewidth of $H$-minor-free graphs [8], and the balanced separator theorem for $H$-minor-free graphs [30], by the Euler-genus of $H$.
1.2 Surfaces and their obstructions
Theorem 1.1 is formulated in terms of the Euler-genus, so it does not give an answer to our initial question in (1) on the parameter $\mathsf{p}_{\Sigma}.$ However, it follows from a recent result of Gavoille and Hilaire [23] (Subsection 1.2) that Dyck-grids are exactly the minor-universal graphs for graphs embeddable in their respective surfaces. Hence, if a graph contains a large enough Dyck-grid for some surface $\Sigma$ as a minor, it also contains every graph embeddable in $\Sigma$ up to a fixed number of vertices as a minor. In fact, we obtain a much more general result that allows us to answer this question precisely. To explain our result in full generality, we shift our focus to (families of) surfaces.
Surfaces.
Given a pair $(h,c)\in\mathbb{N}\times[0,2]$ we define $\Sigma^{(h,c)}$ to be the two-dimensional surface without boundary created from the sphere by adding $h$ handles and $c$ crosscaps. If $c=0,$ the surface $\Sigma^{(h,c)}$ is an orientable surface, otherwise it is a non-orientable surface. By Dyck’s theorem [16, 21], two crosscaps are equivalent to a handle in the presence of a (third) crosscap. This implies that the notation $\Sigma^{(h,c)}$ is sufficient to denote all two-dimensional surfaces without boundary. The Euler-genus, denoted by ${\sf eg}(\Sigma),$ of the surface $\Sigma=\Sigma^{(h,c)}$ is $2h+c.$ Notice that if a surface has odd Euler-genus $g=2t+1$ then it is always homeomorphic to the non-orientable surface $\Sigma^{(t,1)},$ while there are two surfaces, up to homeomorphism, of even, say $2t,$ Euler-genus: the orientable one, namely $\Sigma^{(t,0)},$ and the non-orientable one, that is $\Sigma^{(t-1,2)}.$
We say that a surface $\Sigma$ is contained in some surface $\Sigma^{\prime}$ if $\Sigma^{\prime}$ can be obtained from $\Sigma$ by adding handles and crosscaps. We denote this containment relation by $\Sigma^{\prime}\preceq\Sigma.$ We use $\Sigma^{\varnothing}$ for the empty surface and, as a convention, we agree that $\Sigma^{\varnothing}$ is contained in all other surfaces. Clearly, only the empty graph is embeddable in $\Sigma^{\varnothing}.$ We also fix the Euler-genus of the empty surface $\Sigma^{\varnothing}$ to be $-1.$
Let $\mathbb{S}$ be a set of surfaces. We say that $\mathbb{S}$ is closed under containment, or simply closed, if for every $\Sigma\in\mathbb{S},$ every surface contained in $\Sigma$ also belongs to $\mathbb{S}.$ The correspondence between surfaces and Dyck graphs is given by the following theorem of Gavoille and Hilaire.
[…]
For every pair $(h,c)\in\mathbb{N}\times[0,2],$ a graph is embeddable in $\Sigma^{(h,c)}$ if and only if it is a minor of $\mathscr{D}^{(h,c)}_{k},$ for some $k\in\mathbb{N}.$
More precisely, every graph which is a minor of $\mathscr{D}^{(h,c)}_{k}$ for some $k$ has Euler-genus at most $2c+c$, and for every $n\in\mathbb{N}$, $\mathscr{D}^{(h,c)}_{(2h+c)^{2}(2h+c+n)^{2}}$ contains every $n$-vertex graph which has an embedding in the surface $\Sigma^{(h,c)}$.
Surface obstructions.
Let us fix some finite and closed set $\mathbb{S}$ of surfaces. The surface obstruction set of $\mathbb{S}$, denoted by $\mathsf{sobs}(\mathbb{S}),$ is the set of all $\preceq$-minimal surfaces that do not belong to $\mathbb{S}.$ In other words,
For some examples, notice that $\mathsf{sobs}(\emptyset)=\{\Sigma^{\varnothing}\},$ $\mathsf{sobs}(\{\Sigma^{\varnothing}\})=\{\Sigma^{0,0}\},$ and $\mathsf{sobs}(\{\Sigma^{\varnothing},\Sigma^{(0,0)}\})=\{\Sigma^{(1,0)},\Sigma^{(0,1)}\}$.
See Figure 3 for the lattice corresponding to the surface containment relation and some indicative examples of finite and closed sets of surfaces and their obstruction sets. It is important to observe that every orientable surface $\Sigma$ is contained in some non-orientable surface, while the opposite is not correct. By $\mathcal{E}_{\mathbb{S}}$ we denote the class of graphs that are embeddable in some surface in $\mathbb{S}.$ We stress that $\mathcal{E}_{\{\Sigma^{\varnothing}\}}$ is the graph class containing only the empty graph, while $\mathcal{E}_{\emptyset}=\emptyset.$ Notice that there is a one-to-one correspondence between surfaces and Dyck-grids. For this, given that $\Sigma=\Sigma^{(h,c)},$ we define $\mathscr{D}^{\Sigma}=\langle\mathscr{D}^{\Sigma}_{k}\rangle_{k\in\mathbb{N}}$ as the parametric Dyck-grid $\mathscr{D}^{(h,c)}=\langle\mathscr{D}^{(h,c)}_{k}\rangle_{k\in\mathbb{N}}.$ This permits us to see $\mathscr{D}^{\Sigma}_{k}$ as a parameterization capturing all graphs of $\mathcal{E}_{\{\Sigma\}},$ via minor-containment, as indicated in Subsection 1.2. For completeness, we define $\mathscr{D}^{\Sigma_{\varnothing}}$ as the sequence containing only empty graphs.
Figure 3: The lattice of surface containment, along with some indicative closed sets of surfaces $\mathbb{S}$ (in grey) and their surface obstructions $\mathsf{sobs}(\mathbb{S})$ (depicted in bold frames). Notice that, in case (c), $\mathbb{S}$ is the set of surfaces contained in the Klein bottle $\Sigma^{(0,2)}$ and its the surface obstruction set contains only the torus, i.e., $\mathsf{sobs}(\mathbb{S})=\{\Sigma^{(1,0)}\}.$ For each set of surfaces, the orange box indicates the corresponding prevalent surface
Figure 3: The lattice of surface containment, along with some indicative closed sets of surfaces $\mathbb{S}$ (in grey) and their surface obstructions $\mathsf{sobs}(\mathbb{S})$ (depicted in bold frames). Notice that, in case (c), $\mathbb{S}$ is the set of surfaces contained in the Klein bottle $\Sigma^{(0,2)}$ and its the surface obstruction set contains only the torus, i.e., $\mathsf{sobs}(\mathbb{S})=\{\Sigma^{(1,0)}\}.$ For each set of surfaces, the orange box indicates the corresponding prevalent surface
We are now ready to state our result in its most general and compact form. We define the parameter ${\mathbb{S}}\mbox{-}\mathsf{tw}\colon\mathcal{G_{\text{all}}}\to\mathbb{N}$ such that
${\mathbb{S}}\mbox{-}{\sf tw}(G)\coloneqq\min\{k\mid\mbox{$G$ is in the clique-sum closure of the graphs with a $k$-treewidth modulator to $\mathcal{E}_{\mathbb{S}}\}$}\!\!\!$ (6)
Notice that ${\{\Sigma^{\varnothing}\}}\mbox{-}{\sf tw}{=}{\sf tw}$ and, as a convention, we agree that -${\emptyset}\mbox{-}{\sf tw}=\infty.$ Let $\mathbb{S}^{\prime}$ be a set of surfaces which is not necessarily closed. Next we define the graph parameter ${\mathbb{S}^{\prime}}\mbox{-}\mathsf{bg}\colon\mathcal{G_{\text{all}}}\to\mathbb{N}$ such that
$\mathbb{S}^{\prime}\mbox{-}\mathsf{bg}(G)$$\coloneqq$ $\max\{k\mid\mbox{$G$ contains $\mathscr{D}^{\Sigma}_{k}$ as a minor for some $\Sigma\in\mathbb{S}^{\prime}$}\}.$ (7)
We also agree that ${\{\Sigma^{\varnothing}\}}\mbox{-}\mathsf{bg}=\infty,$ while -${\emptyset}\mbox{-}{\sf bg}$ is the empty function. [^5] The empty function is the function corresponding to the empty set, i.e., is the empty relation. In the following, we are mostly concerned with the case where $\mathbb{S}^{\prime}=\mathsf{sobs}(\mathbb{S}).$
The main result of this paper relates the three concepts defined in (5), (6), and (7) as follows.
Theorem 1.2.
For every finite and closed set of surfaces $\mathbb{S},$ the parameters ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}}$ and ${\mathsf{sobs}(\mathbb{S})}\mbox{-}{\sf bg}$ are asymptotically equivalent, i.e., $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}\sim\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}.$
The above equivalence indicates that min-max duality between surface versions of treewidth and biggest grid is bijectively mapped to the surface obstruction-duality.
In the context of Theorem 1.2, we may view the parametric family $\mathfrak{D}_{\mathsf{sobs}(\mathbb{S})}\coloneqq\{\mathscr{D}^{\Sigma}\mid\Sigma\in\mathsf{sobs}(\mathbb{S})\}$ as a universal obstruction for ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}},$ in the sense that the behaviour of the parameter ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}}$ is asymptotically determined by excluding some graph from each of the parametric graphs in $\mathfrak{D}_{\mathsf{sobs}(\mathbb{S})}$ as a minor (see Subsection 2.1 for the formal definition of universal obstructions).
A graph parameter $\mathsf{p}\colon\mathcal{G}_{\text{all}}\to\mathbb{N}$ is minor-monotone if for every graph $G$ and every minor $H$ of $G$ we have $\mathsf{p}(H)\leq\mathsf{p}(G).$ We say that a graph minor-monotone parameter ${\sf p}\colon\mathcal{G}_{\text{all}}\to\mathbb{N}$ is bounded in $\mathcal{G}$ if the set $\{{\sf p}(G)\mid G\in\mathcal{G}\}$ is finite. By Theorem 1.2, $\mathsf{sobs}(\mathbb{S})$ precisely describes the asymptotic behaviour of ${\mathbb{S}}\mbox{-}\mathsf{tw}$ in minor-closed graph classes, as indicated by the following result (which, essentially, is a restatement of Theorem 1.2).
Theorem 1.3.
Let $\mathbb{S}$ be a finite and closed set of surfaces. Given some minor-closed graph class $\mathcal{G},$ the parameter ${\mathbb{S}}\mbox{-}\mathsf{tw}$ is bounded in $\mathcal{G}$ if and only if, for every $\Sigma\in\mathsf{sobs}(\mathbb{S}),$ $\mathcal{G}$ excludes some graph embeddable in $\Sigma$ as a minor.
Proof.
Suppose that, for every $\Sigma\in\mathsf{sobs}(\mathbb{S}),$ the class $\mathcal{G}$ excludes some graph $H_{\Sigma}$ that is embeddable in $\Sigma.$ Because of Subsection 1.2, there is some $k$ such that $\mathcal{G}$ excludes the graph $\mathscr{D}^{\Sigma}_{k},$ for every $\Sigma\in\mathsf{sobs}(\mathbb{S}).$ Because of (7), this implies that $\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}(G)\leq k$. Therefore, by (1.2), for every graph $G$ in $\mathcal{G},$ it holds that $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}(G)\leq f(k)$, for some function $f\colon\mathbb{N}\to\mathbb{N}$. Hence, ${\mathbb{S}}\mbox{-}\mathsf{tw}$ is bounded in $\mathcal{G}.$ The inverse direction follows the same steps but the in opposite direction. ∎
Derivation of Theorem 1.1.
By the definition of the Euler-genus of a surface, it follows that ${g}\mbox{-}\mathsf{tw}=\mathbb{S}_{g}\mbox{-}{\sf tw}$ where
$\mathbb{S}_{g}\coloneqq\begin{cases}\{\Sigma^{\varnothing}\}&\mbox{if $g=-1$}\\ \{\Sigma\mid\Sigma\preceq\Sigma^{(h,0)}\mbox{\leavevmode\nobreak\ or\leavevmode\nobreak\ }\Sigma\preceq\Sigma^{(h-1,2)}\}&\mbox{if $g=2h\geq 0$}\\ \{\Sigma\mid\Sigma\preceq\Sigma^{(h,1)}\}&\mbox{if $g=2h+1\geq 1$}\end{cases},$
where, as we did for Dyck-grids, we make the convention that $\Sigma^{(-1,2)}=\Sigma^{(0,0)}.$ In other words, $\mathbb{S}_{g}$ is the set of all surfaces of Euler genus $g$, agreeing that $\mathbb{S}_{-1}$ contains only the “empty” surface $\Sigma^{\varnothing}$.
Using the definition of the surface containment relation, it is straightforward to verify the following (see for instance the cases (a), (b), and (f) in Figure 3, corresponding to $g=0,1,$ and $2$ respectively):
$\mathsf{sobs}(\mathbb{S}_{g})=\begin{cases}\{\Sigma^{(0,0)}\}&\mbox{if $g=-1$}\\ \{\Sigma^{(t+1,0)},\Sigma^{(t,1)}\}&\mbox{if $g=2t\geq 0$}\\ \{\Sigma^{(t,2)},\Sigma^{(t+1,0)}\}&\mbox{if $g=2t+1\geq 1.$}\end{cases}$
Recall that ${\mathsf{sobs}(\mathbb{S}_{g})}\mbox{-}\mathsf{bg}$ is the maximum $k$ such that $G$ contains the $k$th member of one of the parametric graphs in $\mathfrak{D}_{\mathsf{sobs}(\mathbb{S}_{g})}$ as a minor. From the the definition of ${(g+1)}\mbox{-}\mathsf{bg},$ it follows that ${(g+1)}\mbox{-}\mathsf{bg}={\mathsf{sobs}(\mathbb{S}_{g})}\mbox{-}\mathsf{bg}.$ Therefore, Theorem 1.1 follows directly from Theorem 1.2, as ${g}\mbox{-}\mathsf{tw}={\mathbb{S}_{g}}\mbox{-}\mathsf{{\sf tw}}\sim{\mathsf{sobs}(\mathbb{S}_{g})}\mbox{-}{\sf bg}={(g+1)}\mbox{-}\mathsf{bg}.$
Prevalent surfaces.
Let $\mathbb{S}$ be a finite and closed set of surfaces. We define the prevalent surface of $\mathbb{{S}},$ denoted by $\Sigma_{\mathbb{S}},$ as the maximal element with respect to the surface containment relation of the set
$\{\Sigma^{\prime}\mid\forall\Sigma\in\mathsf{sobs}(\mathbb{S}):\Sigma^{\prime}\preceq\Sigma\}.$ (8)
[…]
The set in (8) has only one maximal element and this element $\Sigma_{\mathbb{S}}$ is always an orientable surface.
In Figure 3, for each choice of $\mathbb{S},$ the prevalent surface is surrounded by an orange box. For the most simple cases, it holds that $\Sigma_{\emptyset}=\{\Sigma^{\varnothing}\}$, $\Sigma_{\{\Sigma^{\varnothing}\}}=\{\Sigma^{(0,0)}\}$, and $\Sigma_{\{\Sigma^{\varnothing},\Sigma^{(0,0)}\}}=\Sigma^{(0,0)}$. One may observe the following (see Figure 3).
[…]
Let $\mathbb{S}$ be a finite and closed set of surfaces and let $g\coloneqq\max\{{\sf eg}(\Sigma)\mid\Sigma\in\mathbb{S}\}.$ The surface obstruction set $\mathsf{sobs}(\mathbb{S})$ has exactly one obstruction if and only if one of the following conditions holds:
  • all surfaces of Euler-genus $g$ in $\mathbb{S}$ are non-orientable and $g$ is even and positive, or
  • $\mathbb{S}\subseteq\{\Sigma^{\varnothing}\}.$
Moreover, $\Sigma_{\mathbb{S}}\in\mathsf{sobs}(\mathbb{S})$ if and only if $\mathsf{sobs}(\mathbb{S})=\{\Sigma_{\mathbb{S}}\}.$
The next result indicates that all graphs embedded in the prevalent surface of $\mathbb{S}$ appear as minors of a graph $G,$ with ${\mathbb{S}}\mbox{-}{\sf tw}(G)$ large enough.
Theorem 1.4.
For every finite and closed set of surfaces $\mathbb{S}$ there exists a function $f_{\mathbb{S}}\colon\mathbb{N}\to\mathbb{N}$ such that for every $t\in\mathbb{N},$ every graph $G$ with ${\mathbb{S}}\mbox{-}{\sf tw}(G)\geq f_{\mathbb{S}}(t)$ contains every graph on at most $t$ vertices that is embeddable in the prevalent surface $\Sigma_{\mathbb{S}}$ as a minor.
Proof.
Let $H$ be a graph embeddable in the prevalent surface $\Sigma_{\mathbb{S}}$ and let $|H|=t.$ By the definition of $\Sigma_{\mathbb{S}},$ $\Sigma_{\mathbb{S}}$ is contained in every surface $\Sigma^{\prime}\in\mathsf{sobs}(\mathbb{S}),$ therefore $H$ is also embeddable in $\Sigma^{\prime}.$ By Subsection 1.2, there is some $k_{t}$ such that $H$ is a minor of $\mathscr{D}^{\Sigma^{\prime}}_{k_{t}}.$ If now a graph $G$ is $H$-minor free, then it is also $\mathscr{D}^{\Sigma^{\prime}}_{k_{t}}$-minor free, therefore, from (7), $\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}(G)\leq k_{t}.$ This, by Theorem 1.2, implies that $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}(G)\leq f(k_{t})$ (for some function $f\colon\mathbb{N}\to\mathbb{N}$). ∎
In light of our motivating question 1, Theorem 1.4 provides a slightly different perspective. That is, let $H$ be some graph and $\mathbb{S}_{H}$ be a finite and closed set of surfaces such that the prevalent surface of $\mathbb{S}_{H}$ is a surface of smallest Euler-genus where $H$ embedds. Then $\mathbb{S}_{H}\text{-}\mathsf{tw}$ is bounded on all $H$-minor-free graphs.
1.3 The Graph Minors Structure Theorem for fixed surfaces
Let us now go back to the original statement of the GMST. For this, we define the parameter ${\sf hw}^{\prime\prime}\colon\mathcal{G}_{\text{all}}\to\mathbb{N}$ where
The precise definitions of the concepts of “$\Sigma$-decomposition”, “breath”, and “depth” are given in Section 3. Intuitively, a graph $H$ has a $\Sigma$-decomposition with $k$ vortices if $H=H^{(0)}\cup H^{(1)}\cup\cdots\cup H^{(k)}$ where $H^{(0)}$ is a graph embedded in $\Sigma$ and each vortex $H^{(i)}$ is a graph “attached around” the vertices of some face of the embedding of $H^{(0)}.$ The GMST implies that ${\sf hw}^{\prime\prime}\sim{\sf hw}$ (see Theorem 8.2). As we already mentioned, we prove that ${\sf hw}^{\prime\prime}$ can be replaced by ${\sf hw}^{\prime}$ in this relation. Our results have also their counterpart in the setting of ${\sf hw}^{\prime\prime}.$ For this we define, given a finite and closed set of surfaces $\mathbb{S},$ the parameter ${\mathbb{S}}\mbox{-}{\sf hw}\colon\mathcal{G}_{\text{all}}\to\mathbb{N}$ where
Our results imply that the above surface-oriented variant of ${\sf hw}^{\prime\prime}$ is asymptotically equivalent to the corresponding surface variants of ${\sf hw}$ and ${\sf hw}^{\prime}.$
Theorem 1.5.
For every finite and closed set of surfaces $\mathbb{S},$ it holds that the parameters ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}},$ ${\mathbb{S}}\mbox{-}\mathsf{{\sf hw}}$ and ${\mathsf{sobs}(\mathbb{S})}\mbox{-}{\sf bg}$ are asymptotically equivalent, i.e., $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}\sim{\mathbb{S}\mbox{-}\mathsf{{\sf hw}}}\sim{\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}.$
Theorem 1.5 is stated in a more extensive form (as Theorem 8.1) in Section 8, where we make further assumptions on the way the vortices are intersecting the adhesions of the clique-sums. Such assumptions do not (asymptotically) alter the parameters in question and are useful in applications of the decomposition theorems emerging from Theorem 1.5.
1.4 Consequences of our result
All of our proofs are using as a starting point the results of Kawarabayashi, Thomas, Wollan in [31, 32]. As these results are algorithmic, this is also the case for Theorem 1.1 and its more general version, that is Theorem 1.2. This can be stated as follows:
Theorem 1.6.
For every finite and closed set of surfaces $\mathbb{S},$ there exist functions $f_{\mathbb{S}},g_{\mathbb{S}}:\mathbb{N}\to\mathbb{N}$ and an algorithm that, given a graph $G$ and a $k\in\mathbb{N}$, outputs either a Dyck-grid in $\mathfrak{D}_{\mathsf{sobs}(\mathbb{S})}\coloneqq\{\mathscr{D}^{\Sigma}\mid\Sigma\in\mathsf{sobs}(\mathbb{S})\}$ of order $k$ as a minor of $G$ or a tree-decomposition of $G$ where each torso has an $f_{{\mathbb{S}}}(k)$-treewidth modulator to $\mathcal{E}_{\mathbb{S}}$. Moreover, the algorithm runs in time [^6] We use the term $\mathsf{poly}(k)$ as a shortcut for $k^{\mathcal{O}(1)}$. $\mathcal{O}(g_{{\mathbb{S}}}(k)\cdot|V(G)|^{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}3}\log(|V(G)|)),$ where $f_{{\mathbb{S}}}(k)=2^{\mathsf{poly}(k)}$ and $g_{{\mathbb{S}}}(k)=2^{2^{\mathsf{poly}(k)}}$.
One may view the above algorithm as a parameterized approximation algorithm with single-exponential optimality gap. Our combinatorial results together with Theorem 1.6 have a series of algorithmic and graph-structural implications.
It was shown by Demaine and Hajiaghayi [9] that there exists a constant $c_{H}\in\mathbb{N}$ such that for every graph $H$ and every $H$-minor-free graph $G$ and $k\in\mathbb{N}$, if $\mathsf{tw}(G)\geq c_{H}\cdot k$, then $G$ contains the $(k\times k)$-grid as a minor. The bounds on the constant $c_{H}$ were considerably improved by Kawarabayashi and Kobayashi [29]. They showed that $c_{H}=|V(H)|^{\mathcal{O}(|E(H)|)}$. A similar result for every minor-closed class $\mathcal{C}$ of graphs of Euler-genus at most $g$ was proven by Demaine, Hajiaghayi, and Thilikos [10]. They show that the factor in the function relating treewidth and the biggest grid minor only depends on the maximum Euler-genus $g$ of the graphs in $\mathcal{C}$ (but not on the size of the obstructions of $\mathcal{C}$).
The orientable (resp. non-orientable) Euler-genus of a graph $G$ is the minimum Euler-genus of an orientable (resp. non-orientable) surface where $G$ embeds. Given a graph $G$, let us denote by $\mathsf{meg}(G)$ the maximum over the orientable Euler-genus of $G$ and the non-orientable Euler-genus of $G$.
Using known algorithmic techniques from [8] and [48], we can make use of our combinatorial results to obtain the following refinement of the results of [9] and [29]. Since this is not the main focus of our paper and just an illusttrative application of our results we provide only a sketch of the proof here.
Theorem 1.7.
For every graph $H$ there exists a constant $g_{H}=\mathcal{O}(\mathsf{meg}(H))$ and a constant $a_{H}$ such that: For every integer $k$ and every $H$-minor-free graph $G$, if $\mathsf{tw}(G)\geq g_{H}\cdot k+a_{H},$ then $G$ contains a $(k\times k)$-grid as a minor.
Proof sketch.
We denote by $\mathbb{S}_{H}$ the set of surfaces where $H$ cannot be embedded. Observe that $\mathsf{meg}(G)$ is the maximum Euler genus of a surface in $\mathsf{sobs}(\mathbb{S}_{H})$. Moreover, $\mathbb{S}_{H}$ is finite and closed. It follows from Theorem 1.3 that ${\mathbb{S}_{H}}\text{-}\mathsf{tw}$ is bounded on the class of $H$-minor-free graphs. Hence, $G$ has a tree-decomposition whose adhesions are bounded in size by a constant depending only on $|V(H)|$.
We now apply the strategy from the proof of Theorem 3.10 from [8] and use the algorithm from [5] to further separate the torsos of bags which are larger than ${\mathbb{S}_{H}}\text{-}\mathsf{tw}(G)$ by cutting along contractible nooses [^7] Given a $\Sigma$-embedding of a graph $G$, a noose of $G$ is some closed curve that does not meet any edge of the embedding. The order of a noose is the number of vertices that it intersects. of their underlying embedding.
For this purpose, let $t\coloneqq\mathbb{S}_{H}\text{-}\mathsf{tw}(G)$ and let $k$ be the treewidth of $G$. Let $G^{\prime}$ be the torso of some node of a tree-decomposition of $G$ witnessing that $\mathbb{S}_{H}\text{-}\mathsf{tw}(G)=t$ and let $X\subseteq V(t)$ be the modulator of $G^{\prime}$ of annotated treewidth at most $t$. Using the insights of our proof, specifically Subsection 6.2 as in the proof for Theorem 8.2, we may assume here that $X$ is partitioned into a small number of “apex” vertices and the vertices drawn inside the “vortices” of $G^{\prime}$. Notice also that $G^{\prime}-X$ is $\Sigma$-embedded for some surface $\Sigma\in\mathbb{S}_{H}$. Then we modify this embedding by repeatedly identifying non-contractible curves of order at most $k,$ by using the algorithm from [5]. Whenever we find such a noose, we add the vertices that it meets to a set $Y$. Moreover, we find at most $\mathsf{meg}(G^{\prime}-X)$ many such curves. Hence, in particular $|Y|\leq\mathsf{meg}(G^{\prime}-X)\cdot k\leq\mathsf{meg}(H)\cdot k$ which implies that ${\sf tw}(G^{\prime},X\cup Y)\leq t+\mathsf{meg}(H)\cdot k$. Once no such curve can be found any more, it follows that separations of order at most $k$ must correspond to contractible curves in the remaining $\Sigma$-embedding of $G^{\prime}-(X\cup Y)$.
At this point we can use the proof strategy for Theorem 6.5 from [48], together with the results on slopes and respectful tangles from [43] and [45] to deduce that the difference between the treewidth of $G^{\prime}$ and the treewidth of $G^{\prime}-(X\cup Y)$ is at most $\mathcal{O}(\mathsf{meg}(H)\cdot k)+f^{\prime}(\mathsf{tw}(G,X))$ for some computable function $f^{\prime}\colon\mathbb{N}\to\mathbb{N}$. If $k$ is the maximum over the treewidth of all torsos, it follows that either $\mathsf{tw}(G)\leq\mathbb{S}\text{-}\mathsf{tw}(G)$ or there exists a node whose torso witnesses the large treewidth of $G$. In the latter case, our assertion now follows from the main result of [10]. ∎
Theorem 1.7, when combined with Theorem 1.6, immediately improves the dependencies on the constant involved in the approximation algorithm [8] for the treewidth on $H$-minor-free graphs. Specifically in the case where $H$ is rather sparse and, in particular, when the sparsity of $H$ is quantified by its Euler genus. To obtain the following result, notice that the $k$ in the proof above can be found by iterating over all possible $k\leq|V(G)|$ which adds at most a factor of $|V(G)|$ to the running time. As we need cubic time to obtain the tree decomposition and quadratic time to process it the running time remains cubic.
[…]
For every graph $H$ there exist a constant $g_{H}$ and an algorithm that, for every $H$-minor-free graph $G$ produces an $\mathcal{O}(\mathsf{meg}(H))$-approximation for the treewidth of $G$, running in time $\mathcal{O}(g_{H}\cdot|V(G)|^{3}\log(|V(G)|))$.
Note that the dependency on the Euler-genus of $H$ in the approximation factor of Subsection 1.4 can probably not be removed, at least not under certain complexity assumptions from the theory of approximation algorithms. To see this, consider the result of Wu, Per, Toniann, and David [50] which states that, assuming the Small Set Expansion Conjecture (see for example [33, 37]), there cannot be a polynomial-time constant-factor approximation algorithm for treewidth. Suppose there exists an algorith $\mathbf{A}$ and a universal constant $c$ which computes a tree-decomposition of width at most $c\cdot\mathsf{tw}(G)$ for every graph $G$ embedded in a surface of Euler-genus $G$ in time $(g\cdot|V(G)|)^{\mathcal{O}(1)}$. Then one could use the polynomial time algorithm to find a maximum genus-embedding of a graph [22] to turn the entire algorithm into one that runs in polynomial time since the maximum genus of a graph is at most a polynomial in $|V(G)|$. Hence, the existence of such an algorithm $\mathbf{A}$ would disprove the Small Set Expansion conjecture.
Another consequence of Theorem 1.7 is the following general upper bound on the treewidth of an $H$-minor-free graph. For the proof, just combine Theorem 1.7 with the fact that $G$ excludes a $(s\times s)$-grid as a minor, where $s=\lceil|V(G)|^{\nicefrac{{1}}{{2}}}\rceil+1$.
[…]
For every graph $H$ there exist a constant $a_{H}$ such that, for every $H$-minor-free graph $G$, it holds that $\mathsf{tw}(G)\leq\mathcal{O}(\mathsf{meg}(H)\cdot|V(G)|^{\nicefrac{{1}}{{2}}})+a_{H}$.
In turn, Subsection 1.4 implies the following refinement of the seminal result of Kawarabayashi and Reed [30] who proved that every $H$-minor-free graph $G$ has a balanced separator of order $\mathcal{O}(|V(H)|\cdot|V(G)|^{\nicefrac{{1}}{{2}}})$, thereby improving the seminal result of Alon, Seymour, and Thomas [3], where the bound was at most $|V(H)|^{\nicefrac{{3}}{{2}}}\cdot|V(G)|^{\nicefrac{{1}}{{2}}}$.
[…]
For every graph $H$ there exist a constant $a_{H}$ such that every $H$-minor-free graph $G$ has a balanced separator of size at most $\mathcal{O}(\mathsf{meg}(H)\cdot|V(G)|^{\nicefrac{{1}}{{2}}})+a_{H}$.
Notice that when $H$ is a clique, our results above do not improve the previously known results. However, the strength in our results lies in the fact that we decouple the dependencies on the size and the Euler-genus of the excluded minor. In particular, the leading factor in all results above only depends on the later. This means that, whenever the Euler-genus of $H$ is smaller than $|V(H)|$, our results beat all those known so far. Moreover, we believe that the Euler-genus is the more natural parameter since our bounds, as argued before, match the known lower bounds, specifically for graphs of bounded Euler-genus.
We finish with the following corollary of Theorem 1.7 that, we think, is of independent interest.
[…]
For every graph $H$ there is some constant $a_{H}$ such that if a graph $G$ excludes both $H$ and the $(k\times k)$-grid as a minor, then ${\sf tw}(G)\in\mathcal{O}(|V(H)|^{2}\cdot k)+a_{H}$.
For the proof of the above, it is enough to take into account that the Euler genus of a graph $H$ is $O(|V(H)|^{2})$.
Notice that Subsection 1.4 yields a major asymptotic improvement over the result of Kawarabayashi and Kobayashi [29] where the constant is $|V(H)|^{\mathcal{O}(|E(H)|)}$.
1.5 Comparison with other types of annotated extensions of treewidth
Given an annotated graph $(G,X),$ we define $\mathsf{torso}(G,X)$ as the graph obtained by adding, for each connected component $C$ of $G-X,$ all edges between the vertices of $N_{G}(V(C))$ in $G[X]$. Another way to define an annotated version of treewidth was proposed Eiben, Ganian, Hamm, and Kwon [17] where ${\sf tw}^{\prime}(G,X)={\sf tw}(\mathsf{torso}(G,X)).$ This version had interesting algorithmic applications in [27, 2, 17]. Notice that every $X$-rooted minor of $G$ is a minor of $\mathsf{torso}(G,X)$. This fact permits a comparison between ${\sf tw}$ and ${\sf tw}^{\prime},$ as follows.
[…]
For every annotated graph $(G,X),$ ${\sf tw}(G,X)\leq{\sf tw}^{\prime}(G,X).$
We stress that ${\sf tw}(G,X)$ can be much smaller than ${\sf tw}^{\prime}(G,X).$ Let $\Gamma_{k}$ be the $(k\times k)$-grid and let $X$ be the vertices of its perimeter. It holds that ${\sf tw}(G,X)\in\mathcal{O}(1),$ while ${\sf tw}^{\prime}(G,X)\in\Omega(k^{2}).$
1.6 Overview of the proof and the structure of the paper
Towards proving Theorem 1.5 we prove (in reverse order) that ${\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}\preceq\mathbb{S}\mbox{-}\mathsf{{\sf tw}}\preceq\mathbb{S}\mbox{-}\mathsf{{\sf hw}}\preceq{\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}.$
We refer to the relation $\mathbb{S}\mbox{-}\mathsf{{\sf hw}}\preceq{\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}$ as “the upper bound” and its proof spans Section 5. In Section 3 we give a brief introduction to the toolset of graph minors and we present the results that are necessary for the proofs of Section 5. The results of Section 5 are based on the fact that every graph embedded in a surface $\Sigma$ can be seen as the minor of some Dyck-grid $\mathscr{D}^{(h,c)}_{k}$ where $h\in\mathbb{N},c\in\{0,1,2\}$ depend on the choice of $\Sigma$. This is proved in Section 2.
The fact that $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}\preceq\mathbb{S}\mbox{-}\mathsf{\text{${\sf hw}$}}$ follows from Subsection 6.2, which is the main outcome of Section 6.
We finally refer to the relation ${\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}\preceq\mathbb{S}\mbox{-}\mathsf{{\sf tw}}$ Subsection 7.2 as “the lower bound” and we dedicate Section 7 to its proof.
2 Graph parameters and Dyck-grids
We denote by $\mathbb{Z}$ the set of integers and by $\mathbb{R}$ the set of reals. Given two integers $a,b\in\mathbb{Z}$ we denote the set $\{z\in\mathbb{Z}\mid a\leq z\leq b\}$ by $[a,b].$ In case $a>b$ the set $[a,b]$ is empty. For an integer $p\geq 1,$ we set $[p]=[1,p]$ and $\mathbb{N}_{\geq p}=\mathbb{N}\setminus[0,p-1].$
All graphs considered in this paper are undirected, finite, and without loops or multiple edges. We use standard graph-theoretic notation, and we refer the reader to [11] for any undefined terminology. Recall that we use $\mathcal{G}_{\text{all}}$ for the set of all graphs. An annotated graph is a pair $(G,X),$ where $X\subseteq V(G).$
We continue by revisiting the notions and ideas touched upon in the introduction in a more formal way.
2.1 Graph parameters
A graph parameter is a function mapping $\mathcal{G}_{\text{all}}$ to non-negative integers. Let ${\sf p}$ and ${\sf p}^{\prime}$ be two graph parameters. We say that ${\sf p}$ and ${\sf p}^{\prime}$ are asymptotically equivalent, which we denote by ${\sf p}\sim{\sf p}^{\prime}$, if there is a function $f\colon\mathbb{N}\to\mathbb{N}$ such that, for every graph $G,$ it holds that ${\sf p}(G)\leq f({\sf p}^{\prime}(G))$ and ${\sf p}^{\prime}(G)\leq f({\sf p}(G)).$ If the function $f$ is a polynomial (resp. linear) function we say that ${\sf p}_{1}$ and ${\sf p}_{2}$ are polynomially equivalent (resp. linearly equivalent) and we denote this by ${\sf p}\sim_{\mathsf{P}}{\sf p}^{\prime}$ (resp. ${\sf p}\sim_{\mathsf{L}}{\sf p}^{\prime}$).
Next, we introduce tree-decompositions and the concept of clique-sum closures, which can be seen as another way to say that a graph has a tree-decomposition where the torso of every bag satisfies some property.
Tree-decompositions
Let $G$ be a graph. A tree-decomposition of $G$ is a tuple $\mathcal{T}=(T,\beta)$ where $T$ is a tree and $\beta\colon V(T)\to 2^{V(G)}$ is a function, whose images are called the bags of $\mathcal{T},$ such that
  1. 1.
    $\bigcup_{t\in V(T)}\beta(t)=V(G),$
  2. 2.
    for every $e\in E(G)$ there exists $t\in V(T)$ with $e\subseteq\beta(t),$ and
  3. 3.
    for every $v\in V(G)$ the set $\{t\in V(T)\mid v\in\beta(t)\}$ induces a subtree of $T.$
We refer to the vertices of $T$ as the nodes of the tree-decomposition $\mathcal{T}.$
For each $t\in V(T),$ we define the adhesions of $t$ as the sets in $\{\beta(t)\cap\beta(d)\mid d\text{ adjacent with }t\}$ and the maximum size of them is called the adhesion of $t$. The adhesion of $\mathcal{T}$ is the maximum adhesion of a node of $\mathcal{T}.$ The torso of $\mathcal{T}$ on a node $t$ is the graph, denoted by $\textsf{torso}_{G}(t),$ obtained by adding edges between every pair of vertices of $\beta(t)$ which belongs to a common adhesion of $t.$ When $G$ is clear from the context we use $G_{t}$ instead of $\textsf{torso}_{G}(t).$ The clique-sum closure of a graph class $\mathcal{G}$ is the graph class containing every graph that has a tree-decomposition whose torsos belong to $\mathcal{G}.$
The width of a $(T,\beta)$ is the value $\max_{t\in V(T)}|\beta(t)|-1.$ The treewidth of $G,$ denoted by $\mathsf{tw}(G),$ is the minimum width over all tree-decompositions of $G.$ In other words a graph $G$ has treewidth at most $k$ iff $G$ belongs in the clique-sum closure of the class of graphs containing at most $k+1$ vertices.
Parametric graphs and universal obstructions.
Next, we formalize the concept of parametric graphs that can act as asymptotic obstructions to graph parameters.
A parametric graph is a sequence of graphs $\mathscr{H}=\langle\mathscr{H}_{i}\rangle_{i\in\mathbb{N}}$ indexed by non-negative integers. We say that $\mathscr{H}$ is minor-monotone if for every $i\in\mathbb{N}$ we have $\mathscr{H}_{i}\leq\mathscr{H}_{i+1}.$ All parametrized graphs considered in this paper are minor-monotone.
Let $\mathscr{H}^{1}=\langle\mathscr{H}_{k}^{1}\rangle_{k\in\mathbb{N}}$ and $\mathscr{H}^{2}=\langle\mathscr{H}_{k}^{2}\rangle_{k\in\mathbb{N}}$ be two parametric graphs. We write $\mathscr{H}^{1}\lesssim\mathscr{H}^{2}$ if there is a function $f\colon\mathbb{N}\to\mathbb{N}$ such that, for every $k\in\mathbb{N},$ $\mathscr{H}_{k}^{1}\leq\mathscr{H}_{f(k)}^{2}.$ If $f$ is a polynomial (resp. linear) function, then we write $\mathscr{H}^{1}\lesssim_{\mathsf{P}}\mathscr{H}^{2}$ (resp. $\mathscr{H}^{1}\lesssim_{\mathsf{L}}\mathscr{H}^{2}$). We say that $\mathscr{H}^{1}$ and $\mathscr{H}^{2}$ are asymptotically equivalent if $\mathscr{H}^{1}\lesssim\mathscr{H}^{2}$ and $\mathscr{H}^{2}\lesssim\mathscr{H}^{1}$, and we denote this by $\mathscr{H}_{t}^{1}\approx\mathscr{H}^{2}.$ The notion of polynomial (resp. linear) equivalence is denoted $\mathscr{H}_{t}^{1}\approx_{\mathsf{P}}\mathscr{H}^{2}$ (resp. $\mathscr{H}^{1}\approx_{\mathsf{L}}\mathscr{H}^{2}$) and is defined in the obvious way.
For instance the parametric graphs of the grids or of the walls (defined in Section 3) are both linearly equivalent with to annulus grids.
A minor-parametric family is a finite collection =$\mathfrak{H}=\{\mathscr{H}^{(j)}\mid j\in[r]\}$ of (minor-monotone) parametric graphs with $\mathscr{H}^{(j)}=\langle\mathscr{H}^{(j)}_{i}\rangle_{i\in\mathbb{N}}$ for all $j\in[k],$ and for all distinct $i,j\in[r],$ $\mathscr{H}^{(i)}\not\lesssim\mathscr{H}^{(j)}$ and $\mathscr{H}^{(j)}\not\lesssim\mathscr{H}^{(i)}.$ All parametric families considered in this paper are minor-monotone. The notions $\lesssim_{\mathsf{L}},$ $\lesssim_{\mathsf{P}},$ $\equiv_{\mathsf{L}},$ and $\equiv_{\mathsf{P}}$ are defined analogously.
We define the parameter ${\sf p}_{\mathfrak{H}}$ such that ${\sf p}_{\mathfrak{H}}(G)$ is the maximum $k$ for which, there exists some $j\in[r],$ $\mathscr{H}^{j}_{k}\leq G.$
We say that a minor-parametric family =$\mathfrak{H}=\{\mathscr{H}^{(j)}\mid j\in[r]\}$ is a minor-universal obstruction (or universal obstruction if it is clear from the context that we are working with the minor relation) of some graph parameter if ${\sf p}\sim{\sf p}_{\mathfrak{H}}.$ Using this notion we may restate the main result of this paper (Theorem 1.2) as follows:
Recall that in Subsection 1.2 we defined the minor-parametric family $\mathfrak{D}_{\mathsf{sobs}(\mathbb{S})}=\{\mathscr{D}^{\Sigma}\mid\Sigma\in\mathsf{sobs}(\mathbb{S})\}.$ Given the above notation, Theorem 1.2 can be restated as follows.
[…]
For every finite and closed set of surfaces $\mathbb{S},$ the set $\mathfrak{D}_{\mathsf{sobs}(\mathbb{S})}$ is a minor-universal obstruction for ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}}.$
For more on the foundation of universal obstructions of parameters, see [35] (see also [36] for a recent survey).
2.2 Mixed surface grids and Dyck-grids.
The next step is to introduce several classes of grid-like parametric graphs that capture the behaviour of surfaces.
Let $m,n$ be positive integers where $m,n\geq 3$. Given an $(m,n)$-cylindrical grid, let $C_{1},\dots,C_{m}$ be the $m$ disjoint cycles numbered in such a way that $C_{i}$ separates $C_{i-1}$ from $C_{i+1}$ for all $i\in[2,m-1].$ Moreover, let us number the vertices of $C_{i}$ as $v^{i}_{1},\dots,v^{i}_{n}$ such that $v^{1}_{j}v^{2}_{j}\dots v^{m-1}_{j}v^{m}_{j}$ induces is a path for every $j\in[n].$
Consider the $(m,n\cdot 4m)$-cylindrical grid. By adding a handle at position $i\in[n]$ we mean adding the edges
$\{v^{1}_{4m(i-1)+j}v^{1}_{4m(i-1)+3m-j+1}\mid j\in[m]\}\cup\{v^{1}_{4m(i-1)+m+j}v^{1}_{4m(i-1)+4m+j-1}\mid j\in[m]\},$
and subdivide them an arbitrary number of times. We call the paths created by this subdivisions handle paths.
By adding a crosscap at position $i\in[n]$ we mean adding the edges
$\{v^{1}_{4m(i-1)+j}v^{1}_{4m(i-1)+2m+j}\mid j\in[2m]\}$
and subdivide them an arbitrary number of times. We call the paths created by this subdivisions crosscap paths.
Mixed surface grids.
A mixed surface grid of order $k$ with $c$ crosscaps and $h$ handles is a graph constructed as follows. Let $I_{\mathsf{c}},I_{\mathsf{h}}$ be a partition of the set $[2,c+h+1]$ into two sets where $|I_{\mathsf{c}}|=c$ and $|I_{\mathsf{h}}|=h.$ Consider the $(k,4(c+h+1)\cdot k)$-cylindrical grid $H$ and let $H_{I_{\mathsf{c}},I_{\mathsf{h}}}$ be obtained from $H$ by adding a crosscap at every position $i\in I_{\mathsf{c}}$ and adding a handle at every position $i\in I_{\mathsf{h}}.$
[…]
Let $h,c\in\mathbb{N}.$ The $(h,c)$-Dyck-grid of order $k\in\mathbb{N}$ is the graph $\mathscr{D}_{k}^{(h,c)}=H_{[2,h+1],[h+2,c+h+1]}.$ We denote the corresponding parametric graph by $\mathscr{D}^{(h,c)}\coloneqq\langle\mathscr{D}^{(h,c)}_{k}\rangle_{k\in\mathbb{N}}.$ A Dyck-grid $\mathscr{D}^{(h,c)}_{k}$ is said to be orientable if $c=0$ and non-orientable otherwise. We also make the convention that $\mathscr{D}^{(-1,2)}=\mathscr{D}^{(0,0)}.$ The Euler-genus of the Dyck-grid $\mathscr{D}^{(h,c)}_{k}$ is $2h+c.$ We refer to the cycles $C_{k-b-1},\ldots,C_{k}$ as the $b$ outermost cycles of $\mathscr{D}_{k}^{(h,c)}$.
Notice that $\mathscr{D}^{0,0}\sim\mathscr{A}.$ Also, $\mathscr{H}\sim\mathscr{D}^{1,0}$ and $\mathscr{C}\sim\mathscr{D}^{0,1}.$ For a drawing of $\mathscr{D}_{8}^{1,2}$ see of Figure 2.
We say that a surface $\Sigma$ corresponds to a Dyck-grid $\mathscr{D}^{(h,c)}_{k},$ and vice versa, if $\Sigma$ is homeomorphic to a surface that can be obtained by adding $h$ handles and $c$ crosscaps to a sphere, i.e., when $\mathscr{D}^{(h,c)}=\mathscr{D}^{\Sigma}.$
The next theorem is the main result of this section. It formalizes the asymptotic equivalence of mixed surface grids and Dyck-grids.
Theorem 2.1.
Let $g,h,c$ be integers such that $g=2h+c>0$ and $c>0.$ In case $c=2t$ we set $h_{0}\coloneqq t-1$ and $c_{0}\coloneqq 2,$ and in case $c=2t+1$ we set $h_{0}\coloneqq t$ and $c_{0}\coloneqq 1.$ Then the following hold:
  • every mixed surface grid of order $162^{2g}k$ with $c$ crosscaps and $h$ handles contains $\mathscr{D}^{(h+h_{0},c_{0})}_{k}$ as a minor,
  • $\mathscr{D}^{(h+h_{0},c_{0})}_{162^{2g}k}$ contains the mixed surface grid of order $k$ with $c$ crosscaps and $h$ handles as a minor.
In particular, for every minor-monotone parametric graph $\mathscr{X}^{(h,c)}=\langle\mathscr{X}^{(h,c)}_{k}\rangle_{t\in\mathbb{N}},$ where $\mathscr{X}^{(h,c)}_{k}$ denotes a mixed surface grid of order $t$ with $c$ crosscaps and $h$ handles, we have that $\mathscr{X}^{(h,c)}\approx_{{\mathsf{L}}}\mathscr{D}^{(h+h_{0},c_{0})}.$
Figure 4: Swapping the position of a crosscap and a handle. The colors show how the handle and the crosscap on the right are routed through the crosscap and the handle on the left.
At this point, we wish to mention that, for orientable Dyck-grids, the following observation can be made. This observation reflects the fact that a non-orientable surface can never be contained in an orientable one.
[…]
For every $t\in\mathbb{N},$ $t\geq 2,$ and every $k\in\mathbb{N}$ it holds that $\mathscr{D}^{(0,1)}_{t}\not\leq\mathscr{D}^{(t-1,0)}_{k}.$
The rest of this section is dedicated to prove Theorem 2.1. In the following, we will say that a handle transaction is of order $k$ if each of the two transactions involved in the handle are of order $k.$
The next lemma allows us to swap a handle with a neighboring crosscap within a mixed surface grid by sacrificing a portion of its order.
[…]
Let $k,h,c\geq 1$ be integers. Let $H^{\prime}$ be the $(9k,4(c+h+1)\cdot 9k)$-cylindrical grid, let $I^{\prime}_{c},I^{\prime}_{h}$ be a partition of $[2,c+h+1]$ where $|I^{\prime}_{c}|=c$ and $|I^{\prime}_{h}|=h.$ Moreover, let $H$ be the $(k,4(c+h+1)\cdot k)$-cylindrical grid. Finally, let $i\in I^{\prime}_{c}$ and $i+1\in I^{\prime}_{h}$ as well as $j\in I^{\prime}_{h}$ and $j+1\in I^{\prime}_{c}$ and set
$I_{c}\coloneqq(I^{\prime}_{c}\setminus\{i\})\cup\{i+1\},$ $I_{h}\coloneqq(I^{\prime}_{h}\setminus\{i+1\})\cup\{i\},$
$J_{c}\coloneqq(I^{\prime}_{c}\setminus\{j+1\})\cup\{j\},$ $J_{h}\coloneqq(I^{\prime}_{h}\setminus\{j\})\cup\{j+1\}.$
Then the mixed surface grid $H^{\prime}_{I^{\prime}_{c},I^{\prime}_{h}}$ of order $5k$ with $c$ crosscaps and $h$ handles contains as a minor
  • the mixed surface grid $H_{I_{c},I_{h}}$ of order $k$ with $c$ crosscaps and $h$ handles and
  • the mixed surface grid $H_{J_{c},J_{h}}$ of order $k$ with $c$ crosscaps and $h$ handles.
Proof.
To see that the claim of this lemma is true consider Figure 4. Since the construction is completely symmetric it suffices to consider the case where $i\in I^{\prime}_{c}$ and $i+1\in I^{\prime}_{h}$ as depicted in the figure. Notice that the handle transaction at $i+1$ partitions the set $[4\cdot 9k\cdot i+1,8\cdot 9k\cdot i]$ into four intervals, namely $[4\cdot 9k\cdot i+1,4\cdot 9k\cdot i+9k],$ $[4\cdot 9k\cdot i+9k+1,4\cdot 9k\cdot i+2\cdot 9k],$ $[4\cdot 9k\cdot i+2\cdot 9k+1,4\cdot 9k\cdot i+3\cdot 9k],$ and $[4\cdot 9k\cdot i+3\cdot 9k+1,4\cdot 9k\cdot(i+1)].$ The first and third interval form the endpoints of one of the two transactions that form the handle, the second transaction involved in the handle has its endpoints in the second and fourth interval. We fix the two intervals $[4\cdot 9k\cdot i+32k+1,4\cdot 9k\cdot i+34k]$ and $[4\cdot 9k\cdot i+34k+1,4\cdot 9k\cdot(i+1)]$ to host the starting and endpoints of a crosscap transaction of order $2k$ which is necessary for a mixed surface grid of order $k.$ As depicted in Figure 4, we may route the two pairs of $2k$ disjoint paths, from each of the two terminal intervals, through the handle. By doing so, each handle is traversed two times per pair which takes up $8k$ paths of each of the two transactions of the handle in total. Thus, as depicted in Figure 4, for each of the two transactions involved in the handle we may keep $k$ paths that allow us to maintain a handle transaction of order $k.$ Moreover, the may route the crosscap in such a way that it avoids the cycles $C_{8k+1},\dots,C_{9k}$ since we are only routing $8k$ paths in total. By combining these cycles with $4k$ paths from the grid-structure at position $0,$ a sub-handle, sub crosscap of each other handle or crosscap transaction of order $k$ each, and the newly routed handle and crosscap transactions, we obtain the desired mixed surface grid of order $k.$
2.3 A graph-minor analogue of Dyck’s lemma
The next step is to convert three consecutive crosscaps into a handle followed by a crosscap. Moreover, we show that the reversal of this exchange is possible. The numbers in the following lemma are not necessarily optimal, they are adjusted to make for a nicer presentation.
We remark that Subsection 2.3 can be seen as a graph-theoretical analogue of Dyck’s theorem, stated in terms of graph minors.
Figure 5: Three crosscaps in a row.
[…]
Let $k,h,c\geq 0$ be integers such that $k\geq 1$ and $c\geq 3.$ Let $H^{\prime}$ be the $(18k,4(c+h+1)\cdot 18k)$-cylindrical grid and $H^{\prime\prime}$ be the $(18k,4(c-2+h+1+1)\cdot 18k)$-cylindrical grid. Let $I^{\prime}_{c},I^{\prime}_{h}$ be a partition of $[2,c+h+1]$ where $|I^{\prime}_{c}|=c$ and $|I^{\prime}_{h}|=h$ and there exists some integer $i\in I^{\prime}_{c}$ such that $i+1,i+2\in I^{\prime}_{c},$ moreover, let $I^{\prime\prime}_{c},I^{\prime\prime}_{h}$ be a partition of $[2,c+h]$ such that $i\in I^{\prime\prime}_{h}$ and $i+1\in I^{\prime\prime}_{c}.$ Finally, let $H^{1}$ be the $(k,4(c+h)\cdot k)$-cylindrical grid and $H^{2}$ be the $(k,4(c+h+1)\cdot k)$-cylindrical grid.
Then the following hold:
  1. 1.
    the mixed surface grid $H^{\prime}_{I^{\prime}_{c},I^{\prime}_{h}}$ of order $18k$ with $c$ crosscaps and $h$ handles contains the mixed surface grid $H^{1}_{I^{\prime\prime}_{c},I^{\prime\prime}_{h}}$ of order $k$ with $c-2$ crosscaps and $h+1$ handles as a minor, and
  2. 2.
    the mixed surface grid $H^{\prime\prime}_{I^{\prime\prime}_{c},I^{\prime\prime}_{h}}$ of order $18k$ with $c-2$ crosscaps and $h+1$ handles contains the mixed surface grid $H^{2}_{I^{\prime}_{c},I^{\prime}_{h}}$ of order $k$ with $c$ crosscaps and $h$ handles as a minor.
Proof.
We divide the proof of this lemma into two steps for the proof of 1 (visualized in Figure 5, Figure 6, and Figure 7) and one for the proof of the reverse direction 2 (visualized in Figure 8, Figure 9, and Figure 10).
Figure 6: The result of the construction from Figure 5.
Proof of 1. We begin by routing three sets of paths through the three consecutive crosscaps. Notice that each of the three crosscap transactions contains a total of $2\cdot 18k=36k$ paths, however, each of our path-bundles will only contain $12k$ paths. This allows us to avoid the cycles $C_{12k+1},\dots,C_{18k}$ which we need to form the annulus of the resulting mixed surface grid. The first bundle of paths starts in the crosscap at position $i$ on the cycle $C_{18k},$ traverses this crosscap, traverses the crosscap at position $i+1$ and then ends again on the cycle $C_{18k}.$ The second bundle starts left (with respect to the ordering of the vertices on $C_{18k}$ as induced by the definition of mixed surface grids) of the endpoints of the first bundle. It traverses the crosscap at position $i+1$ and then the crosscap at position $i+2.$ Finally, the third bundle starts left of the end of the second bundle in the crosscap at position $i+2,$ traverses only this one crosscap, and ends on the right of the second bundle. See Figure 5 for an illustration of this first construction.
Now keep these three bundles, each of order $12k,$ keep $12k$ paths of each other crosscap transaction and keep $6k$ paths of each of the two transactions involved in every other handle transaction. Moreover, keep $12k$ paths from the grid structure at position 00 together with the cycles $C_{12k+1},\dots,C_{18k}$ The resulting graph almost resembles a mixed surface grid with the exception of positions $i,$ $i+1,$ and $i+2,$ here we have two planar and one crosscap transaction intertwined. See Figure 6 for an illustration of the situation.
Figure 7: The result of the construction from Figure 6. The colors indicate the different bundles of paths used to route the transactions.
The final step is to untangle the crosscap at position $i+2$ from the newly created handle. Notice that “position $i+2$” refers only to the bundle of paths forming the crosscap transactions as we are not in a mixed surface grid any more. Indeed, by merging the other two crosscaps into a single handle, we have lost a position. Similar to the proof of Subsection 2.2 we fix two disjoint intervals of endpoints, one within the right most part of the crosscap transaction on the cycle $C_{18k},$ and the other in the right most part of the middle transaction that is part of the handle. Both intervals are of order $2k.$ We then route a crosscap transaction from the left of these intervals to the right. As depicted in Figure 6 we may traverse the handle transaction which allows us to traverse the crosscap exactly once. Hence, the resulting transaction is, indeed, a crosscap. Since each of the two transactions involved in the handle are of order $12k,$ and we only spend $4k$ of their paths to route the crosscap, there are enough paths left to choose $k$ paths each in order to form a handle transaction of order $k.$ By restricting all other handles and crosscaps accordingly and keeping the cycles $C_{17k+1},\dots,C_{18k},$ we obtain a mixed surface grid of order $k$ as desired. See Figure 7 for an illustration.
Figure 8: A handle and a crosscap in a row together with the routing for the crosscap that allows to recreate the situation from Figure 6.
Proof of 2. The reverse transformation works along similar lines but is not completely analogous. For the sake of completeness we include illustrations of the three steps that decompose a handle and a crosscap into three distinct crosscaps.
The first step is to recreate the entangled situation from Figure 6 by routing the crosscap backwards through the handle as depicted in Figure 8. By doing so we divide the order of the original crosscap transaction by three as we need to route $6k$ paths which will later be divided into three disjoint crosscaps, but we also need to keep $6k$ paths from each of the two transactions of order $18k$ that form the handle. As we only route $6k$ paths in this step, we are able to keep the cycles $C_{6k+1},\dots,C_{18k}$ for the next step.
Figure 9: The result of the construction from Figure 8.
With this, we have now established a situation similar to the one from Figure 6 as depicted in Figure 9. We may now split the crosscap transaction of order $6k$ into three, each of them of order $2k.$ Finally, we make use of the entangled handle to create three distinct crosscap transactions of order $2k$ as depicted in Figure 10.
Since we route $4k$ paths through the annulus, we may, in particular, avoid the cycles $C_{17k+1},\dots,C_{18k}$ and thus, by restricting the grid structure at position 00 and all other handle and crosscap transactions accordingly, we obtain the desired mixed surface grid and our proof is complete. ∎
Figure 10: The result of the construction from Figure 9. The colors indicate the bundles of paths used to create the three consecutive crosscaps.
2.4 Proof of Theorem 2.1
We are finally ready for the proof of Theorem 2.1.
Proof of Theorem 2.1.
Let $H$ be a mixed surface grid of order $162^{h_{0}+c+h}\cdot k$ with $h$ handles and $c$ crosscaps.
For each of the $c$ crosscaps we need to call Subsection 2.2 at most $h$ times to achieve a situation where all handles are consecutive and thus obtain a surface grid. That is, if we aim for a surface grid of order $w$ with $h$ handles and $c$ crosscaps, we may start out with a mixed surface grid of order $9^{c+h}\cdot w.$ Moreover, this is enough to move any crosscap essentially anywhere along the cycle $C_{1}.$ That means that also the surface grid of order $9^{c+h}\cdot w$ contains every mixed surface grid of order $w$ with $h$ handles and $c$ crosscaps as a minor.
In particular, if $H^{\prime}$ is the surface grid of order $18^{h_{0}}\cdot k,$ then $H^{\prime}$ is a minor of $H$ by the reasoning above.
Now, to obtain a Dyck-grid of order $w$ with $h+h_{0}$ handles and $c_{0}$ crosscaps from a surface grid with $h$ handles and $c$ crosscaps we may proceed as follows: Choose three consecutive crosscaps at positions $i,$ $i+1,$ and $i+2$ such that the transaction at position $i-1$ is either a handle, or $i-1=0.$ Then apply Subsection 2.3 to these three crosscaps. The resulting minor we find is still a surface grid, but now on $h+1$ handles and $c-2$ crosscaps and its order has decreased by a factor of $18$ As we have to apply this exactly $h_{0}$ times to reduce the number of crosscaps below three, by starting out with a surface grid of order $18^{h_{0}}\cdot w$ with $h$ handles and $c$ crosscaps, we obtain the Dyck-grid of order $w$ and Euler-genus $2(h+h_{0})+c_{0}=2h+c$ as a minor. Notice that Subsection 2.3 states that reversing this operation also costs a factor of $18,$ that means that the (orientable in case $c=0$ and otherwise non-orientable) Dyck-grid of order $18^{h_{0}}\cdot w$ contains the surface grid of order $w$ with $h$ handles and $c$ crosscaps as a minor. In fact, for any $h^{\prime}\leq h_{0},$ it also contains the surface grid of order $w$ with $h+(h_{0}-h^{\prime})$ handles and $c-2(h_{0}-h^{\prime})$ crosscaps as a minor.
In particular, our graph $H^{\prime}$ from above contains a Dyck-grid of order $k$ and Euler-genus $2h+c$ as a minor.
Now let $g\coloneqq 2h+c$ as in the statement of our theorem and notice the following inequalities:
$c$ $\leq g,$
$2h$ $\leq g,\text{ and}$
$2h_{0}$ $\leq c,$
It follows that
$162^{h_{0}+c+h}\cdot k$ $\leq 162^{\frac{1}{2}g+g+\frac{1}{2}g}\cdot k$
$\leq 162^{2g}\cdot k.$
This is true for any choice of $h^{\prime}$ and $c^{\prime}$ such that $2h^{\prime}+c^{\prime}=g.$ Hence, by the discussion above, that any mixed surface grid of order $162^{2g}\cdot k$ with $h^{\prime}$ handles and $c^{\prime}$ crosscaps where $2h^{\prime}+c^{\prime}=g$ contains a Dyck-grid of order $k$ and Euler-genus $g$ as a minor. Moreover, the (orientable in case $c^{\prime}=0$ and otherwise non-orientable) Dyck-grid of order $162^{2g}k$ contains every mixed surface grid of order $k$ with $h^{\prime}$ handles and $c^{\prime}$ crosscaps as a minor. Hence, our proof is complete. ∎
3 A brief introduction to the toolset of graph minors
To prove the upper bound in Theorem 1.2, that is ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}}\preceq{\mathsf{sobs}(\mathbb{S})}\mbox{-}{\sf bg},$ we make use of the powerful tools developed around the Graph Minors Series of Robertson and Seymour. In particular, we will be using a theorem developed by Kawarabayashi, Thomas, and Wollan in their proof of the GMST [32] which makes the extraction of some surface grid a very achievable task. However, to be able to even state this result we need to introduce a lot of machinery. The purpose of this section is the introduction and collection of the various technical definitions from [32] we need.
3.1 Walls an tangles
Walls.
An $(n\times m)$-grid is the graph $G_{n,m}$ with vertex set $[n]\times[m]$ and edge set
$\{\{(i,j),(i,j+1)\}\mid i\in[n],j\in[m-1]\}\cup\{\{(i,j),(i+1,j)\}\mid i\in[n-1],j\in[m]\}.$
We call the path $(i,1)(i,2)\dots(i,m)$ the $i$th row and the path $(1,j)(2,j)\dots(n,j)$ the $j$th column of the grid. An elementary $k$-wall $W_{k}$ for ${k\geq 3},$ is obtained from the ${(k\times 2k)}$-grid $G_{k,2k}$ by deleting every odd edge in every odd column and every even edge in every even column, and then deleting all degree-one vertices. The rows of $W_{k}$ are the subgraphs of $W_{k}$ induced by the rows of $G_{k,2k},$ while the $j$th column of $W_{k}$ is the subgraph induced by the vertices of columns ${2j-1}$ and ${2j}$ of $G_{k,2k}.$ We define the perimeter of $W_{k}$ to be the subgraph induced by ${\{(i,j)\in V(W_{k})\mid j\in\{1,2,2k,2k-1\}\text{ and }i\in[k]\textnormal{, or }i\in\{1,k\}\text{ and }j\in[2k]\}}.$ A $k$-wall $W$ is a graph isomorphic to a subdivision of $W_{k}.$ The vertices of degree three in $W$ are called the branch vertices. In other words, $W$ is obtained from a graph $W^{\prime}$ isomorphic to $W_{k}$ by subdividing each edge of $W^{\prime}$ an arbitrary (possibly zero) number of times. The perimeter of $W,$ denoted by ${\sf Perimeter}{W},$ is the subgraph isomorphic to the subgraph of $W^{\prime}$ induced by the vertices of the perimeter of $W_{k}$ together with the subdivision vertices of the edges of the perimeter of $W_{k}.$ We define rows and columns of $k$-walls analogously to their definition for elementary walls. A wall is a $k$-wall for some $k.$
An $h$-wall $W^{\prime}$ is a subwall of some $k$-wall $W$ where $h\leq k$ if every row (column) of $W^{\prime}$ is contains in a row (column) of $W.$
The following is a statement of the Grid Theorem. While we will not explicitly use the Grid Theorem, we will, later on, make use of the existence of a function that forces the existence of a large wall in graph with large enough treewidth.
[…]
There exists a function $\mathsf{p}\colon\mathbb{N}\to\mathbb{N}$ with $\mathsf{p}(n)\in\mathcal{O}(n^{9}\operatorname{polylog}(n))$ such that for every $k\in\mathbb{N}$ and every graph $G,$ if ${\sf tw}(G)>\mathsf{p}(k)$ then $G$ contains a $k$-wall as a subgraph.
We stress that the above result holds if we replace walls by grids as these two parametric graphs are linearly equivalent.
Tangles.
In [41] Robertson and Seymour introduced tangles as the max-min counterpart of the parameter of branchwidth, that is linearly equivalent to treewidth. Tangles have been important in the proofs of the Graph Minors series of Robertson and Seymour. Also, they play an important role in an abstract theory of connectivity, see for example [26, 14, 15, 13, 12].
[…]
A separation in a graph $G$ is a pair $(A,B)$ of vertex sets such that $A\cup B=V(G)$ and there is no edge in $G$ with one endpoint in $A\setminus B$ and the other in $B\setminus A.$ The order of $(A,B)$ is $|A\cap B|.$ Let $G$ be a graph and $k$ be a positive integer. We denote by $\mathcal{S}_{k}(G)$ the collection of all tuples $(A,B)$ where $A,B\subseteq V(G)$ and $(A,B)$ is a separation of order $<k$ in $G.$ An orientation of $\mathcal{S}_{k}(G)$ is a set $\mathcal{O}$ such that for all $(A,B)\in\mathcal{S}_{k}(G)$ exactly one of $(A,B)$ and $(B,A)$ belongs to $\mathcal{O}.$ A tangle of order $k$ in $G$ is an orientation $\mathcal{T}$ of $\mathcal{S}_{k}(G)$ such that
for all $(A_{1},B_{1}),(A_{2},B_{2}),(A_{3},B_{3})\in\mathcal{T},$ it holds that $A_{1}\cup A_{2}\cup A_{3}\neq V(G).$ (11)
Let $W$ be an $r$-wall in a graph $G$ and assume that the horizontal and vertical paths are specified. Let $\mathcal{T}_{W}$ be the orientation of $\mathcal{S}_{r}$ such that for every $(A,B)\in\mathcal{T}_{W},$ $B\setminus A$ contains the vertex set of both a horizontal and a vertical path of $W$, we call $B$ the $W$-majority side of $(A,B)$. Then $\mathcal{T}_{W}$ is the tangle induced by $W$.
Similarly, let $h\in\mathbb{N}$ and $c\in[0,2]$ and let $D$ be an $(h,c;t)$-Dyck-wall in $G$. Then also $D$ defines a majority side for any separation of order less than $r$. In this case the majority side is the sida that contains a full cycle and a column of the annulus wall of $D$. In the same way as above this allows us to define the tangle $\mathcal{T}_{D}$ induced by $D$.
Well-linked-sets and tangles.
Let $α\in[2/3,1),$ $G$ be a graph and $S\subseteq V(G)$ be a vertex set. A set $X\subseteq V(G)$ is said to be an $α$-balanced separator for $S$ if for every component $C$ of $G-X$ it holds that $|C(C)\cap S|\leq α|S|.$ Let $q\in\mathbb{N}.$ We say that $S$ is a $(q,α)$-well-linked set of $G$ if there is no $α$-balanced separator of size at most $q$ for $S$ in $G.$ Given a $(q,α)$ well-linked set of $G$ we define
$\mathcal{T}_{S}\coloneqq\{(A,B)\in\mathcal{S}^{q+1}\mid|S\cap B|>α|S|\}.$
It is not hard to see that $\mathcal{T}_{S}$ is a tangle of order $q$ in $G$.
Since it is easier to represent well-linked sets of obunded size with a bounded amount of memory, we opt to work with well-linked sets instead of tangles throughout this paper. Please note that, with the observation above, this still allows us to make use of all results of graph minors that need a large oder tangle as part of the input.
Let $G$ be a graph and $v\in V(G)$ be a vertex of degree at least four. We say that $G^{\prime}$ is obtained from $G$ by splitting $v$ if there exist an edge $x_{1}x_{2}\in E(G^{\prime})$ such that $G-v=G^{\prime}-x_{1}-x_{2}$ and $N_{G}(v)=N_{G^{\prime}}(x_{1})\cup N_{G^{\prime}}(x_{2})$ where $N_{G^{\prime}}(x_{1})\cap N_{G^{\prime}}(x_{2})\neq\emptyset$. If there exists a vertex $v$ such that a graph $G^{\prime}$ is obtained from a graph $G$ by splitting $v$, we say that $G^{\prime}$ is obtained from $G$ by splitting a vertex. A pair $(H,T)$ is said to be an expansion of a graph $G$, or a $G$-expansion, if $H$ is a subdivision of a graph $G^{\prime}$ obtained from $G$ by a sequence of vertex-splits, every vertex $v$ with $deg_{H}(v)\geq 3$ belongs to $T\subseteq V(H)$, and $G^{\prime}$ is the graph obtained from $H$ by dissolving every vertex of $V(H)\setminus T$. We call $T$ the branch vertices of $(H,T)$.
Let $G$ and $H$ be graphs and let $p\coloneqq|V(H)|.$ Let $\mathcal{T}$ be a tangle in $G.$ An $H$-expansion in $G$ is controlled by $\mathcal{T}$ if for every separation $(A,B)\in\mathcal{T}$ with $|A\cap B|<p,$ there does not exist a branch vertex of the expansion contained in $A\setminus B.$ For a set $Z\subseteq V(G)$ we denote by $\mathcal{T}-Z$ the tangle in $G-Z$ formed by the separations $(A,B)$ in $G-Z$ such that $(A\cup Z,B\cup Z)\in\mathcal{T}.$ Given a surface $\Sigma$ and a $\Sigma$-decomposition $\delta$ of $G,$ we say that the decomposition is $\mathcal{T}$-central if for every separation $(A,B)\in\mathcal{T},$ there does not exist a cell $c\in C(\delta)$ such that $B\subseteq V(\Sigma_{\delta}(c)).$
Next we need a tool that allows us to relate a given $\Sigma$-decomposition to some tangle. We add a small bit of information to the statement of the lemma which is part of the proof of Lemma 15.3 in [32] but not made explicit in their formulation of the statement.
A tangle $\mathcal{T}^{\prime}$ in $G$ of order at most $k$ is a restriction of $\mathcal{T}$ if $\mathcal{T}^{\prime}\subseteq\mathcal{T}$.
[…]
Let $G$ be a graph, and let $k,\ell,r\geq 0$ such that $r\geq k(2\ell+1)+6$. Let $W$ be an $r$-wall in $G$. Let $\mathcal{T}$ be a tangle of order at least $k(2\ell+1)+6$ such that $\mathcal{T}_{W}$ is a restriction of $\mathcal{T}$. Assume that there exists a $\Sigma$-decomposition $\delta$ of $G$ of breadth $k$ and depth $\ell$ such that $W$ is flat in $\delta$. Then there exists $Z\subseteq V(G),$ $|Z|\leq k(2\ell+1)$ and a $\Sigma$-decomposition $\delta^{\prime}$ of $G-Z$ of breadth at most $k$ and depth at most $\ell$ which is $(\mathcal{T}-Z)$-central.
Moreover, the set $Z$ consists of at most $12\ell+$ 1 vertices from each vortex of $\delta$ and every vertex of $Z$ belongs of some vortex of $\delta^{\prime}$.
3.2 Growing a wall from a well-linked set
In this subsection we describe how, given a $(s,\alpha)$-well-linked set $S$ for large enough $s$ we can find a large wall $W$ in FPT-time such that $\mathcal{T}_{W}$ is a truncation of $\mathcal{T}_{S}.$ This will give us the right to equip our local structure theorems with an additional property that will allow us to localise it with respect to some tangle which will be encoded using the set $S$ instead of the collection of separations which can be very large. Towards this goal, we present an algorithmic version of Lemma 14.6 from [32]. That is, we prove the following theorem.
Theorem 3.1.
Let $k\geq 3$ be an integer, $\alpha\in[2/3,1).$ There exist universal constants $c_{1},c_{2}\in\mathbb{N}\setminus\{0\},$ and an algorithm that, given a graph $G$ and an $(36c_{1}k^{20}+3,\alpha)$-well-linked set $S\subseteq V(G)$ computes in time $2^{\mathcal{O}(k^{c_{2}})}|V(G)|^{2}|E(G)|\log(|V(G)|)$ a $k$-wall $W\subseteq G$ such that $\mathcal{T}_{W}$ is a truncation of $\mathcal{T}_{S}.$
Our proof for Theorem 3.1 combines results from three different papers. For the sake of completeness, since we need to modify the proofs of most of these results slightly, we present a complete and self contained proof here.
The first step is to show that, given our well-linked set $S$ we can compute from it a set with slightly stronger properties which also induces a tangle and whose tangle is a truncation of $\mathcal{S}.$
To stress that all results presented here are algorithmic, we phrase them mostly using the well-linked set $S$ and only make use of abstract properties of $\mathcal{T}_{S}$ without requiring the tangle to be given.
Let $\alpha\in[2/3,1)$ be fixed, $k$ be a positive integer, and $S$ be a $(k,\alpha)$-well-linked set in a graph $G.$ A set $F$ is said to be $S$-free in $G$ if for every separation $(A_{1},A_{2})\in\mathcal{S}_{|F|}$ it holds that, if $F\subseteq A_{i}$ for $i\in[2],$ then $|A_{i}\cap S|>\alpha|S|.$
The proof of the following lemma is an adaptation of Lemma 14.1 from [32].
[…]
Let $k$ be a positive integer, $\alpha\in[2/3,1),$ $G$ be a graph, and $S\subseteq V(G)$ be $(k,\alpha)$-well-linked. Then there exists an $S$-free set $F$ of size $k-1.$
Moreover, there exists an algorithm that, given $G$ and $S,$ computes $F$ in time $\mathcal{O}(k|V(G)|^{2}).$
Proof.
First let us observe that, if $G$ is disconnected, there must exist some component $G^{\prime}$ of $G$ such that $|V(G^{\prime})\cap S|>\alpha|S|$ since $S$ is well-linked. Hence, if fix any vertex $x_{0}\in V(G^{\prime})$ and set $F_{0}\coloneqq x_{0},$ then $F_{0}$ is $S$-free.
As $S$ is $(k,\alpha)$-well-linked it follows that $|S|\geq 3k+1$ since otherwise we could delete $k$ vertices from $S$ and any component in the resulting graph has at most $\frac{2}{3}|S|\leq\alpha|S|$ vertices of $S.$
Assume we have already constructed an $S$-free set $F_{i}$ of size $i+1$ for $i\in[0,k-2].$
Now let $(A,B)$ be a separation of order $i+1$ such that $F_{i}\subseteq A$ and $|B\cap S|>\alpha|S|.$ Such a separation exists since $(F_{i},V(G))$ satisfies this property by the size of $S.$ We now iteratively “push” $(A,B)$ towards $S$ by asking, in each step, for a separation of size at most $i+1$ between $A\cap B$ and $B\cap S.$ This is, in order to minimize $|B|.$ Once this minimization step concludes, that is after at most $|V(G)|$ iterations, each of which takes $\mathcal{O}(k)|V(G)|$ time, we select $x_{i+1}\in B\setminus A$ to be any vertex.
Suppose, $F_{i+1}\coloneqq F_{i}\cup\{x_{i+1}\}$ is not $S$-free. Then there exists a separation $(X,Y)$ of order at most $i+1$ where $F_{i+1}\subseteq X$ and $|Y\cap S|>\alpha|S|.$ Moreover, such a separation can be found in time $\mathcal{O}(k)|V(G)|.$ It follows from the assumption that $F_{i+1}$ is $S$-free that $|X\cap Y|=i+1.$
Now, consider the separation $(A\cap X,B\cup Y).$ Since $|Y\cap S|>\alpha|S|$ it follows that $|(B\cup Y)\cap S|>\alpha|S|.$ From the assumption that $F_{i}$ is $S$-free this means that $(A\cap X,B\cup Y)$ is of order at least $i+1.$ Thus, by the submodularity of separations in undirected graphs, it follows that $(A\cup X,B\cap Y)$ is of order at most $i+1.$ Sine $A$ is a proper subset of $A\cup X,$ our construction of $(A,B)$ implies that $|(A\cup X)|>\alpha|S|.$ However, this means that $A\cup X\cup(B\cap Y)=V(G)$ while $(A,B),(X,Y),(B\cap Y,A\cup X)\in\mathcal{T}_{S}.$ This is a clear vialation of the tangle axioms and therefore impossible. Hence, a separation like $(X,Y)$ cannot exist and this means that $F_{i+1}$ is, in fact, $S$-free. ∎
Free sets are themselves highly connected. Indeed, they exhibit a slightly stronger property than just being $(k,\alpha)$-well-linked.
Let $G$ be a graph and $S\subseteq V(G)$ be a set of vertices. We say that $S$ is strongly linked [^8] In [32] this property is called “well-linked” but we decide to change it here to not overload the term. if for every partition $S_{1},S_{2}$ of $S$ into two sets there does not exist a separation $(A_{1},A_{2})$ of $G$ such that $S_{i}\subseteq A_{i}$ for both $i\in[2],$ and $|A_{1}\cap A_{2}|<\min\{|S_{1}|,|S_{2}|\}.$ Notice that every strongly linked set of size $13k+$ 1 is also $(k,2/3)$-well-linked.
The following is a corollary of Lemma 14.2 from [32] phrased in terms of well-linked set instead of tangles.
[…]
Let $\alpha\in[2/3,1),$ $k\geq 1$ be an integer, $G$ be a graph and $S\subseteq V(G)$ be $(k,\alpha)$-well-linked.
If $F\subseteq V(G)$ is $S$-free with $|F|<k$ then $F$ is strongly linked.
We believe that it is possible to prove the following sequence of lemmas purely in terms of $(k,\alpha)$-well-linked sets, but for the sake of brevity we stick to the setting of Kawarabayashi, Thomas, and Wollan [32].
[…]
Let $\alpha\in[2/3,1),$ $k\geq 1$ be an integer, $G$ be a graph and $S\subseteq V(G)$ be $(3k+1,\alpha)$-well-linked.
If $F\subseteq V(G)$ is $S$-free with $|F|=3k$ then $\mathcal{T}_{F}\coloneqq\{(A,B)\in\mathcal{S}_{k}\mid|B\cap F|>2k\}$ is a tangle of order $k$ which is a truncation of $\mathcal{T}_{S}.$
Proof.
First let us show that for every $(A,B)\in\mathcal{S}_{k}$ we have $\{(A,B),(B,A)\}\cap\mathcal{T}_{F}\neq\emptyset.$ Since $|F|=3k$ this implies that $\mathcal{M}$ contains exactly one of $(A,B)$ and $(B,A)$ for evry $(A,B)\in\mathcal{S}_{k}.$
So let $(A,B)\in\mathcal{S}_{k}.$ This means $|A\cap B|<k.$ Hence, since $F$ is strongly linked by Subsection 3.2, this means that $\min\{|A\cap F|,|B\cap F|\}<k.$ With $|F|=3k$ this implies that $\max\{|A\cap F|,|B\cap F|\}>2k.$
Moreover, let $(A_{1},B_{1}),(A_{2},B_{2}),(A_{3},B_{3})\in\mathcal{T}_{F}.$ Then $|A_{i}\cap F|\leq k-1$ for all $i\in[3].$ Therefore, $|(A_{1}\cup A_{2}\cup A_{3})\cap F|\leq 3k-3<3k$ and thus, the union of three small sides cannot cover all of $V(G)$ and so $\mathcal{T}_{F}$ is indeed a tangle of order $k.$
Finally, we have to show that $\mathcal{T}_{F}$ is a truncation of $\mathcal{T}_{S}.$ To see this suppose there is some $(A,B)\in\mathcal{T}_{F}$ such that $(B,A)\in\mathcal{T}_{S}.$ Notice that this means that $|A\cap B|<k$ and $|B\cap F|>2k$ by definition of $\mathcal{T}_{F}.$ Hence, $|A\cap F|\leq k-1.$ Thus, $(A,B\cup F)$ is a separation of order less than $2k.$ As $F\subseteq B\cup F$ it follows from the definition of $S$-free sets that $|(B\cup F)\cap S|>\alpha|S|$ and thus, by definition of $\mathcal{T}_{S}$ we have $(A,B\cup F)\in\mathcal{T}_{S}.$ However, as $(B,A)\in\mathcal{T}_{S}$ by assumption, we now have that two small sides, namely $A$ and $B,$ of separations in $\mathcal{T}_{S}$ cover all of $V(G).$ This is clearly a contradiction to $\mathcal{T}_{S}$ being a tangle and thus we must ahve that $\mathcal{T}_{F}\subseteq\mathcal{T}_{S}.$
An interesting, separate corollary of the proof of Subsection 3.2 which is reflected in Lemma 14.4 from [32] is the following.
[…]
Let $k\geq 1$ be an integer, $G$ be a graph and $F\subseteq V(G)$ be a strongly linked set of size $3k$ in $G.$ Then $\mathsf{tw}(G)\geq k.$
We will also need the following observation on grid-minors. This observation will be helpful in determining if a grid we have found yields a tangle that agrees with $\mathcal{T}_{S}.$
[…]
Let $k\geq 2$ be an integer, $G$ be a graph and $(M,T)$ be an expansion of the $(k\times k)$-grid in $G.$
For every set $S$ obtained by selecting a single row or column $Q$ of the $(k\times k)$-grid and then, for each vertex $v\in V(Q)$ selecting a single vertex $x_{v}\in X_{v}$ which is of degree at least three in $M$ it holds that $S$ is strongly linked.
If the set $S$ as above is taken from the $i$th row (column), we call $S$ a representative set of the $i$th row (column) of $(M,T).$ Since every $k$-wall naturally correspondonds to an expansion of the $(k\times k)$-grid.
Next, we adapt Lemma 14.5 from [32] for our purposes. That is, we present almost the same proof as the one found in the work of Kawarabayashi, Thomas, and Wollan while stressing that it is algorithmic.
[…]
Let $k\geq 1$ be an integer, $G$ be a graph and $X,Y\subseteq V(G)$ be strongly linked sets such that $|X|\geq k$ and $|Y|\geq 3k.$ Exactly one of the following two statement holds.
  1. 1.
    There exist $k$ pairwise vertex-disjoint $X$-$Y$-paths in $G,$ or
  2. 2.
    if $(A,B)$ is a separation of order $<k$ such that $X\subseteq A$ and $|Y\cap B|\geq k,$ then either
    1. (a)
      there exists a separation $(A^{\prime},B^{\prime})$ of order $<k$ such that $X\subseteq A^{\prime},$ $|Y\cap B^{\prime}|\geq k,$ $A\subseteq A^{\prime},$ and $B^{\prime}\subsetneq B,$ or
    2. (b)
      there exists an edge $e$ of $G[B]$ it holds that $X$ is strongly linked in $G-e.$
Moreover, there exists an algorithm that, in time $2^{\mathcal{O}(k)}\cdot|V(G)|,$ finds either the paths, the separation $(A^{\prime},B^{\prime}),$ or the edge $e$ as above.
Proof.
Given the sets $X$ and $Y$ we may determine in time $\mathcal{O}(k\cdot|V(G)|)$ if there exist $k$ pairwise vertex-disjoint paths between $X$ and $Y$ or find some separation $(A,B)$ of order less than $k$ such that $X\subseteq A$ and $Y\subseteq B.$ In particular, it holds that $|B\cap Y|\geq k.$
So from now on we assume that we are given some separation $(A,B)$ of order at most $k-1$ such that $X\subseteq A,$ and $|Y\cap B|\geq k.$
Fix an edge $xy=e\in E(G[B]).$ We can test in time $2^{\mathcal{O}(k)}|V(G)|$ if there exists a partition of $X$ into sets $X_{1}$ and $X_{2}$ such that there exists a separation $(L,R)$ of $G-e$ with $X_{1}\subseteq L,$ $X_{2}\subseteq R,$ and $|L\cap R|<p\coloneqq\min\{|X_{1}|,|X_{2}|\}.$ If such a partition cannot be found we have successfully determined that $X$ is strongly linked in $G-e$ and can output the $e.$ Otherwise we continue to work with the separation $(L,R).$
Without loss of generality we may assume $x\in L.$ Indeed, it must now hold that $x\in L\setminus R,$ and $y\in R\setminus L$ since otherwise $(L,R)$ is witness that $X$ is not strongly linked in $G$ which contradicts our assumption. Moreover, $x$ has at least one neighbour in $L\setminus R$ as otherwise it can be moved to $R$ to form a separation of $G$ which vialoates the strong linkedness of $X.$ Similarly, $y$ has some neighbour in $R\setminus L.$
Observe that $(A\cap L,B\cup R)$ is a separation of $G$ with $X_{1}\subseteq A\cap L$ and $X_{2}\subseteq B\cup R.$ Hence, it must be at least of order $p+1.$ Let $q\coloneqq|A\cap B|<k.$ It follows from the submodularity of separations in undirected graphs that $(A\cup L,B\cap R)$ has order at most $q-1.$ Notice that, by choice of $e,$ we have $x\notin B\cap R.$ Hence, $(A\cup L,(B\cap R)\cup\{x\})$ is a separation of $G$ of order at most $q<k.$ Moreover, $(B\cap R)\cup\{x\}\subsetneq B.$
In the case where $|(B\cap R)\cup\{x\}\cap Y|\geq k.$ We have found our desired separation and may terminate. Hence, we may assume $|(B\cap R)\cup\{x\}\cap Y|\leq k-1.$
Following a symmetric argument we may analyse the separations $(A\cap R,B\cup L)$ and $(A\cup R,B\cap L)$ and reach the same two possible outcomes. Either $(A\cup R,(B\cap L)\cup\{y\})$ is the separation we were looking for, or we also obtain $|(B\cap L)\cup\{y\}|\leq k-1.$
Now if none of the two separations we have found is stuitable for our needs we have $|(B\cap R)\cup\{x\}\cap Y|\leq k-1$ and $|(B\cap L)\cup\{y\}|\leq k-1.$ It now follows that $|B\cap Y|\leq 2k-2.$ This, however, would imply that $|A\cap Y|\geq k+2$ which contradicts our initial assumption of $(A,B).$ This completes the proof. ∎
The last piece towards the proof of Theorem 3.1 is an adaptation of a theorem of Fomin, Lokshtanov, Panolan, Saurabh, and Zehavi [19, 20], that is Proposition 3.6 in [20].
Moreover, we need a way to compute a tree decomposition of bounded width if one exists. To this end, we use Korhonen’s [34] recent result.
[…]
There exists an algorithm that, given a graph $G$ and an integer $k$ outputs, in time $2^{\mathcal{O}(k)}|V(G)|,$ a tree decomposition for $G$ of width at most $12k+$ 1 or concludes that $\mathsf{tw}(G)>k.$
[…]
There exists a universal constant $c\geq 1$ and an algorithm that, given a graph $G$ and a $(3\cdot c\cdot k^{10}+1,\alpha)$-well-linked set $S\subseteq V(G)$ for some integer $k$ and some $\alpha\in[2/3,1)$ computes in time $2^{\mathcal{O}(k^{10}\log k)}|V(G)|\log(|V(G)|)$ a $k$-wall $G.$
Proof.
Our goal is to find a subgraph $G^{\prime}$ of $G$ whose treewidth is at least $c_{1}k^{10}$ for some constant $c_{1}$ and at most $c_{2}k^{10}$ for another constant $c_{2}\geq c_{1}.$ Once we have found this subgraph, we may use Subsection 3.2 to compute a tree decomposition of bounded width for $G$ and then proceed with dynamic programming to check if $G^{\prime}$ contains the $(k\times k)$-grid as a minor (using i.e., the algorithm of [1]). Since $G^{\prime}$ satisfies $\mathsf{tw}(G^{\prime})\geq c_{1}k^{10},$ by choosing $c_{1}$ to be the constant from Subsection 3.1, we are guaranteed to find the desired $k$-wall.
In the following we describe how to find the graph $G^{\prime}.$ This idea is taken directly from [20] and we only modify two details. Note that the highly well-linked set ensures that we can never run into the case where we determine that $G$ has small treewidth.
Let us fix $c$ to be the constant hidden in the $\mathcal{O}$-notation of Subsection 3.1 and set $t\coloneqq c\cdot k^{10}.$ We initialise as follows: Set $A_{0}\coloneqq V(G)$ and $F_{0}\coloneqq\emptyset.$ Notice that the pair $(A_{0},F_{0})$ satisfies the following conditions for $i=0,$ in particular since $S$ is $(3t+1,\alpha)$-well linked:
  • $\mathsf{tw}(G[F_{i}])\leq 2t,$ and
  • $\mathsf{tw}(G[A_{i}\cup F_{i}])>t.$
The process will stop as soon as it determines that $\mathsf{tw}(G[A_{i}\cup F_{i}])\leq 2t.$
Suppose we have already constructed the pair $(A_{i-1},F_{i-1})$ for some $i\geq 1$ satisfying the two conditions above while $\mathsf{tw}(G[A_{i-1}\cup F_{i-1}])>2t.$ Let $S_{i}\subseteq A_{i-1}$ be an arbitrary set of size $\lceil\frac{A_{i-1}}{2}\rceil.$ There are two cases to consider.
Case 1:
$\mathsf{tw}(G[F_{i-1}\cup S_{i}])\leq 2t$, in this case we set $F_{i}\coloneqq F_{i-1}\cup S_{i}$ and $A_{i}\coloneqq A_{i-1}\setminus S_{i}.$ Notice that this preserves the two conditions above.
Case 2:
$\mathsf{tw}(G[F_{i-1}\cup S_{i}])>t$, here we set $F_{i}\coloneqq F_{i-1}$ and $A_{i}\coloneqq S_{i}$ and still satisfy our conditions.
Clearly, this procedure must stop after at most $\log(|V(G)|)$ many steps as either we reach a situation where $G[A_{i}\cup F_{i}]$ has treewidth at most $2t,$ or $A_{i}=\emptyset.$ It follows from our conditions above that in the later case we also have that $t<\mathsf{tw}(G[F_{i}\cup A_{i}])\leq 2t.$ Hence, this algorithm is always successfull. Moreover, in each step we call Subsection 3.2 along with minor checking on bounded treewidth graphs which can be done in time $2^{\mathcal{O}(t)}(k^{2})^{2t}2^{\mathcal{O}(k^{2})}\cdot|V(G)|,$ using, e.g., the minor-checking algorithm of [1]. This leads to the final running time. ∎
Notice that Subsection 3.2 can be combined with Reed’s algorithm for finding either a tree decomposition of boundd width or a $(k,2/3)$-well-linked set [38] to produce an updated version of the original result in [20, Proposition 3.6].
[…]
There exists a universal constant $c>1$ and an algorithm that, given an integer $k$ and a graph $G,$ either computes a tree decomposition of width at most $ck^{10}$ for $G,$ or a $k$-wall in $G$ in time $2^{\mathsf{poly}(k)}|V(G)|\log(|V(G)|).$
We are now ready for the proof of Theorem 3.1. This is an algorithmic version of Lemma 14.6 from [32].
Proof of Theorem 3.1.
Let us set $c_{1}$ to be the constant from Theorem 3.1 and set $t\coloneqq 6c_{1}k^{20}.$ Then $S$ is a $(3t+3,\alpha)$-well-linked set in $G.$ By using Subsection 3.2 we can find an $S$-free set $F$ of size $13t+$ 1 in time $\mathcal{O}(t|V(G)|^{2}).$ Moreover, by Subsection 3.2 $F$ is strongly linked, and by Subsection 3.2 $\mathcal{T}_{F}$ is a tangle of order $13t+$ 1 which is a truncation of $\mathcal{T}_{S}.$ Indeed, we can also observe that $F$ is $(3t+1,2/3)$-well-linked in $G.$
We will now start iterating the following process working on the graph $G_{i},$ where $G_{0}\coloneqq G$ while maintaining the invariant that $F$ is strongly linked in $G_{i}$:
  1. 1.
    We use Subsection 3.2 to find a $6k^{2}$-wall $W$. in $G_{i}.$
  2. 2.
    We then select a representative set $X$ of the first column of $W,$ let us denote this path by $P.$ Indeed, we may make sure to select $X$ to only come from the first column of $W.$ Then, we ask for a linkage of size $2k^{2}$ from $F$ to $X$ in $G_{i}.$ This request has one of two initial outcomes:
    1. (a)
      We find the linkage, or
    2. (b)
      we find a separation $(A,B)$ of order at most $12k^{2}-$ 1 such that $F\subseteq A$ and $X\subseteq B.$
    In the first outcome we are happy, return $W$ and terminate, but in the second outcome we may call for Subsection 3.2 to either “push” $(A,B)$ closer to $X,$ or find an edge $e$ of $G-i$ such that $F$ is still strongly linked in $G_{i}-e.$ By iterating at most $|E(G_{i})|$ many times, Subsection 3.2 will eventually produce such an edge. We now set $G_{i+1}\coloneqq G_{i}-e$ and start the next iteration.
Notice that the above procedure stops after at most $|E(G)|$ iterations, justifying our running time.
Next, we show that, if the above process produces a $2k^{2}$-linkage from $F$ to $X$ this certifies that the tangle $\mathcal{T}_{W}$ is indeed a truncation of $\mathcal{T}_{F}$ which, in turn, is a truncation of $\mathcal{T}_{S}$ as desired.
Let $\mathcal{P}$ be the $F$-$X$-linkage of order $2k^{2}$ and let $X^{\prime}\subseteq X$ be the set of endpoints of the paths in $\mathcal{P}.$ Recall that $P$ is the first column of $W$ and $X^{\prime}\subseteq X\subseteq V(P).$
We now partition $L_{1}\coloneqq P$ into $k$ pairwise disjoint subpaths $R_{1},\dots,R_{k}$ such that $|V(R_{i})\cap X^{\prime}|\geq 2k$ for each $i\in[k].$ For each $i\in[k]$ let $r_{i}$ be the first vertex of $R_{i}$ encountered when traversing along $L_{1}$ and let $Q_{i}$ be the unique row of $W$ containing $r_{i}.$ Finally we choose $L_{2},\dots,L_{k}$ to be $k-1$ arbitrary and distinct colums of $W$ which are all distinct from $L_{1}.$ Let $W^{\prime}$ be the $k$-subwall of $W$ with rows $Q_{1},\dots,Q_{k}$ and columns $L_{1},\dots,L_{k}.$
Now all we have to do is to show that $\mathcal{T}_{W^{\prime}}$ is a truncation of $\mathcal{T}_{F}.$ Let $(A,B)\in\mathcal{T}_{F}$ be a separation of order at most $k-1.$ It follows that $|A\cap F|\leq|A\cap B|\leq k-1.$ Moreover, there exists some $j\in[k]$ such that $Q_{j}\cup R_{j}$ is vertex-disjoint from $A\cap B.$ Since there are $2k$ pairwise vertex-disjoint paths from $R_{j}$ to $F$ by the choice of $R_{j}$ it follows from $(A,B)\in\mathcal{T}_{F}$ that $V(Q_{j}\cup R_{j})\subseteq B\setminus A.$ Moreover, there exists some $j^{\prime}\in[k]$ such that $L_{j^{\prime}}$ is vertex-disjoint from $A\cap B.$ Since $L_{j^{\prime}}$ has a non-empty intersection with $Q_{j}$ we must have that $V(Q_{j}\cup L_{j^{\prime}})\subseteq B\setminus A$ and thus it follows that $(A,B)\in\mathcal{T}_{W^{\prime}}.$
To complete the proof, observe that eventually the process above must find such a linkage as otherwise we run out of edges to delete. ∎
3.3 $\Sigma$-decompositions
We start by introducing the notions we will be using to describe “almost embeddings”’ in surfaces.
[…]
Let $\Omega$ be a cyclic permutation of the elements of some set which we denote by $V(\Omega).$ A society is a pair $(G,\Omega),$ where $G$ is a graph and $\Omega$ is a cyclic permutation with $V(\Omega)\subseteq V(G).$ A cross in a society $(G,\Omega)$ is a pair $(P_{1},P_{2})$ of disjoint paths [^9] When we say two paths are disjoint we mean that their vertex sets are disjoint. in $G$ such that $P_{i}$ has endpoints $s_{i},t_{i}\in V(\Omega)$ and is otherwise disjoint from $V(\Omega),$ and the vertices $s_{1},s_{2},t_{1},t_{2}$ occur in $\Omega$ in the order listed.
[…]
(Drawing in a surface) A drawing (with crossings) in a surface $\Sigma$ is a triple $\Gamma=(U,V,E)$ such that
  • $V$ and $E$ are finite,
  • $V\subseteq U\subseteq\Sigma,$
  • $V\cup\bigcup_{e\in E}e=U$ and $V\cap(\bigcup_{e\in E}e)=\emptyset,$
  • for every $e\in E,$ $e=h((0,1)),$ where $h\colon[0,1]_{\mathbb{R}}\to U$ is a homeomorphism onto its image with $h(0),h(1)\in V,$ and
  • if $e,e^{\prime}\in E$ are distinct, then $|e\cap e^{\prime}|$ is finite.
We call the set $V,$ sometimes referred to by $V(\Gamma),$ the vertices of $\Gamma$ and the set $E,$ referred to by $E(\Gamma),$ the edges of $\Gamma$. If $G$ is graph and $\Gamma=(U,V,E)$ is a drawing with crossings in a surface $\Sigma$ such that $V$ and $E$ naturally correspond to $V(G)$ and $E(G)$ respectively, we say that $\Gamma$ is a drawing of $G$ in $\Sigma$ (possibly with crossings). In the case where no two edges in $E(\Gamma)$ have a common point, we say that $\Gamma$ is a drawing of $G$ in $\Sigma$ without crossings. In this last case, the connected components of $\Sigma\setminus U,$ are the faces of $\Gamma.$
[…]
Let $\Sigma$ be a surface. A $\Sigma$-decomposition of a graph $G$ is a pair $\delta=(\Gamma,\mathcal{D}),$ where $\Gamma$ is a drawing of $G$ is $\Sigma$ with crossings, and $\mathcal{D}$ is a collection of closed disks, each a subset of $\Sigma$ such that
  1. 1.
    the disks in $\mathcal{D}$ have pairwise disjoint interiors,
  2. 2.
    the boundary of each disk in $\mathcal{D}$ intersects $\Gamma$ in vertices only,
  3. 3.
    if $\Delta_{1},\Delta_{2}\in\mathcal{D}$ are distinct, then $\Delta_{1}\cap\Delta_{2}\subseteq V(\Gamma),$ and
  4. 4.
    every edge of $\Gamma$ belongs to the interior of one of the disks in $\mathcal{D}.$
Let $N$ be the set of all vertices of $\Gamma$ that do not belong to the interior of the disks in $\mathcal{D}.$ We refer to the elements of $N$ as the nodes of $\delta.$ If $\Delta\in\mathcal{D},$ then we refer to the set $\Delta-N$ as a cell of $\delta$, i.e., a cell is an closed disk where we removed at most three nodes from its boundary. We denote the set of nodes of $\delta$ by $N(\delta)$ and the set of cells by $C(\delta).$ For a cell $c\in C(\delta)$ the set of nodes that belong to the closure of $c$ is denoted by $\widetilde{c}.$ Notice that this means that the cells $c$ of $\delta$ with $\widetilde{c}\neq\emptyset$ form the edges of a hypergraph with vertex set $N(\delta)$ where $\widetilde{c}$ is the set of vertices incident with $c.$
For a cell $c\in C(\delta)$ we define the graph $\Sigma_{\delta}(c),$ or $\Sigma(c)$ if $\delta$ is clear from the context, to be the subgraph of $G$ consisting of all vertices and edges drawn in the closure of $c.$
We define $\pi_{\delta}\colon N(\delta)\to V(G)$ to be the mapping that assigns to every node in $N(\delta)$ the corresponding vertex of $G.$
[…]
Let $G$ be a graph, $\Sigma$ be a surface and $\delta=(\Gamma,\mathcal{D})$ be a $\Sigma$-decomposition of $G.$ A cell $c\in C(\delta)$ is called a vortex if $|\widetilde{c}|\geq 4.$ Moreover, we call $\delta$ vortex-free if no cell in $C(\delta)$ is a vortex.
[…]
Let $(G,\Omega)$ be a society, and let $\Sigma$ be a surface with one boundary component $B.$ A rendition in $\Sigma$ of $G$ is a $\Sigma$-decomposition $\rho$ of $G$ such that the image under $\pi_{\rho}$ of $N(\rho)\cap B$ is $V(\Omega),$ mapping one of the two cyclic orders of $B$ to the order of $\Omega.$
Let $\rho=(\Gamma,\mathcal{D})$ be a rendition of $(G,\Omega)$ in a disk. If there exists a cell $c_{0}\in C(\delta)$ such that no cell in $C(\delta)\setminus\{c_{0}\}$ is a vortex we call the tripe $(\Gamma,\mathcal{D},c_{0})$ a cylindrical rendition of $(G,\Omega)$.
Let $\rho=(\Gamma,\mathcal{D})$ be a rendition of a society $(G,\Omega)$ in a surface $\Sigma.$ For every cell $c\in C(\rho)$ with $|\widetilde{c}|=2$ we select one of the components of $\mathsf{bd}(c)-\widetilde{c}.$ This selection is called a tie-breaker in $\rho$, and we assume every rendition to come equipped with a tie-breaker. Let $Q\subseteq G$ be either a cycle or a path that uses no edge of $\Sigma(c)$ for every vortex $c\in C(\rho).$ We say that $Q$ is grounded in $\rho$ if either $Q$ is a non-zero length path with both endpoints in $\pi_{\rho}(N(\rho)),$ or $Q$ is a cycle, and it uses edges of $\Sigma(c_{1})$ and $\Sigma(c_{2})$ for two distinct cells $c_{1},c_{2}\in C(\rho).$ A $2$-connected subgraph $H$ of $G$ is said to be grounded if every cycle in $H$ is grounded.
If $Q$ is grounded we define the trace of $Q$ as follows. Let $P_{1},\dots,P_{k}$ be distinct maximal subpaths of $Q$ such that $P_{i}$ is a subgraph of $\Sigma(c)$ for some cell $c$ and $Q=\bigcup_{i\in[k]}P_{k}$. Fix an index $i.$ The maximality of $P_{i}$ implies that its endpoints are $\pi(n_{1})$ and $\pi(n_{2})$ for distinct nodes $n_{1},n_{2}\in N(\rho).$ If $|\widetilde{c}|=2,$ define $L_{i}$ to be the component of $\mathsf{bd}(c)-\{n_{1},n_{2}\}$ selected by the tie-breaker, and if $|\widetilde{c}|=3,$ define $L_{i}$ to be the component of $\mathsf{bd}(c)-\{n_{1},n_{2}\}$ that is disjoint from $\widetilde{c}.$ Finally, we define $L_{i}^{\prime}$ by slightly pushing $L_{i}$ to make it disjoint from all cells in $C(\rho).$
We define such a curve $L_{i}^{\prime}$ for all $i$ while ensuring that the curves intersect only at a common endpoint. The trace of $Q$ is defined to be $\bigcup_{i\in[k]}L_{i}^{\prime}.$ So the trace of a cycle is the homeomorphic image of the unit circle, and the trace of a path is an arc in $\Sigma$ with both endpoints in $N(\rho).$
Paths, linkages and transactions.
If $P$ is a path and $x$ and $y$ are vertices on $P,$ we denote by ${xPy}$ the subpath of $P$ with endpoints $x$ and $y.$ Moreover, if $s$ and $t$ are the endpoints of $P,$ and we order the vertices of $P$ by traversing $P$ from $s$ to $t,$ then ${xP}$ denotes the path ${xPt}$ and ${Px}$ denotes the path ${sPx}.$ Let $P$ be a path from $s$ to $t$ and $Q$ be a path from $q$ to $p.$ If $x$ is a vertex in ${V(P)\cap V(Q)}$ such that $Px$ and $xQ$ intersect only in $x$ then ${PxQ}$ is the path obtained from the union of $Px$ and $xQ.$ Let ${X,Y\subseteq V(G)}.$ A path is an $X$-$Y$-path if it has one endpoint in $X$ and the other in $Y$ and is internally disjoint from ${X\cup Y},$ Whenever we consider $X$-$Y$-paths we implicitly assume them to be ordered starting in $X$ and ending in $Y,$ except if stated otherwise. An $X$-path is an $X$-$X$-path of length at least one. In a society $(G,\Omega),$ we write $\Omega$-path as a shorthand for a $V(\Omega)$-path.
Let $(G,\Omega)$ be a society. A segment of $\Omega$ is a set ${S\subseteq V(\Omega)}$ such that there do not exist ${s_{1},s_{2}\in S}$ and ${t_{1},t_{2}\in V(\Omega)\setminus S}$ such that ${s_{1},t_{1},s_{2},t_{2}}$ occur in $\Omega$ in the order listed. A vertex ${s\in S}$ is an endpoint of the segment $S$ if there is a vertex ${t\in V(\Omega)\setminus S}$ which immediately precedes or immediately succeeds $s$ in the order $\Omega.$ For vertices ${s,t\in V(\Omega)},$ if $t$ immediately precedes $s$ we define $s\Omega t$ to be the trivial segment $V(\Omega),$ and otherwise we define $s\Omega t$ to be the uniquely determined segment with first vertex $s$ and last vertex $t.$
[…]
Let $G$ be a graph. A linkage in $G$ is a set of pairwise vertex-disjoint paths. In slight abuse of notation, if $\mathcal{L}$ is a linkage, we use $V(\mathcal{L})$ and $E(\mathcal{L})$ to denote ${\bigcup_{L\in\mathcal{L}}V(L)}$ and ${\bigcup_{L\in\mathcal{L}}E(L)}$ respectively. Given two sets $A$ and $B$ we say that a linkage $\mathcal{L}$ is an $A$-$B$-linkage if every path in $\mathcal{L}$ has one endpoint in $A$ and one endpoint in $B.$
We call $\lvert\mathcal{L}\rvert$ the order of $\mathcal{L}.$
[…]
Let ${(G,\Omega)}$ be a society. A transaction in ${(G,\Omega)}$ is an $A$-$B$-linkage for disjoint segments $A,B$ of $\Omega.$ We define the depth of ${(G,\Omega)}$ as the maximum order of a transaction in ${(G,\Omega)}.$
Let $(G,\Omega)$ be a society and let $\mathcal{P}$ be a transaction on $(G,\Omega).$ We say that $\mathcal{P}$ is a
  • crosscap transaction if the members of $\mathcal{P}$ can be numbered $P_{1},P_{2},\dots,P_{n}$ and the endpoints of $P_{i}$ denoted by $u_{i}$ and $v_{i}$ in such a way that
    $u_{1},u_{2},\dots,u_{n},v_{1},v_{2},\dots,v_{n}$
    appear in $\Omega$ in the order listed. We say that $\mathcal{P}$ is a crosscap transaction of thickness $n$.
  • handle transaction on $(G,\Omega)$ if the members of $\mathcal{P}$ can be numbered $P_{1},P_{2},\dots,P_{2n}$ and the endpoints of $P_{i}$ can be denoted by $u_{i}$ and $v_{i}$ in such a way that
    $u_{1},u_{2},\dots,u_{2n}v_{n},v_{n-1},\dots,v_{1},v_{2n},v_{2n-1},\dots,v_{n+1}$
    appear in $\Omega$ in the order listed. We say that $\mathcal{P}$ is a handle transaction of thickness $n$.
Notice that a handle transaction of thickness $1$ is also a crosscap transaction of thickness $1$ We call such a transaction a cross.
Theorem 3.2 (Two Paths Theorem, [28, 44, 46, 49, 40]).
A society $(G,\Omega)$ has no cross if and only if it has a vortex-free rendition in a disk.
Linear decompositions.
In the original Graph Minors Series by Robertson and Seymour, the depth of vortices was not described in terms of their largest transaction, but in terms of another type of decomposition that allows to break the graph inside a vortex into smaller pieces, all of which attach only via a small set of vertices.
Let $(G,\Omega)$ be a society. A linear decomposition of $(G,\Omega)$ is a sequence
$\mathcal{L}=\langle X_{1},X_{2},\dots,X_{n},v_{1},v_{2},\dots,v_{n}\rangle$
such that $v_{1},v_{2},\dots,v_{n}\in V(\Omega)$ occur in that order on $\Omega$ and $X_{1},X_{2},\dots,X_{n}$ are subsets of $V(G)$ such that
  • for all $i\in[n],$ $v_{i}\in X_{i},$
  • $\bigcup_{i\in[n]}X_{i}=V(G),$
  • for every $uv\in E(G),$ there exists $i\in[n]$ such that $u,v\in X_{i},$ and
  • for all $x\in V(G),$ the set $\{i\in[n]\mid x\in X_{i}\}$ forms an interval in $[n].$
The adhesion of a linear decomposition is $\max_{i\in[n-1]}|X_{i}\cap X_{i+1}|,$ its width is $\max_{i\in[n]}|X_{i}|.$
[…]
If a society $(G,\Omega)$ has depth at most $θ,$ then it has a linear decomposition of adhesion at most $2θ.$ Moreover there exists an algorithm that, given $(G,\Omega)$, either finds a transaction of order >$𝜃$ or outputs such a decomposition in time $\mathcal{O}(θ|V(G)|^{2}).$
[…]
Let $\Sigma$ be a surface and $G$ be a graph. Let ${\delta=(\Gamma,\mathcal{D})}$ be a $\Sigma$-decomposition of $G.$ Every vortex $c$ defines a society ${(\Sigma(c),\Omega)},$ called the vortex society of $c,$ by saying that $\Omega$ consists of the vertices $\pi_{\delta}(n)$ for ${n\in\widetilde{c}}$ in the order given by $\Gamma.$ (There are two possible choices of $\Omega,$ namely $\Omega$ and its reversal. Either choice gives a valid vortex society.) The breadth of $\delta$ is the number of cells ${c\in C(\delta)}$ which are a vortex and the depth of $\delta$ is the maximum depth of the vortex societies ${(\Sigma(c),\Omega)}$ over all vortex cells ${c\in C(\delta)}.$
3.4 Local structure theorems
To proceed, we need to combine our definition of flat walls with the idea of $\Sigma$-decompositions. This is a necessary step to be able to relate flat renditions of walls to the drawings provided by $\Sigma$-decompositions and thus to impose additional structure onto these drawings. Using this, we will introduce the general concept of a local structure theorem.
Let $G$ be a graph and ${W}$ be a wall in $G.$ We say that $W$ is flat in a $\Sigma$-decomposition $\delta$ of $G$ if there exists a closed disk ${\Delta\subseteq\Sigma}$ such that
  • the boundary of $\Delta$ does not intersect any cell of $\delta,$
  • ${\pi(N(\delta)\cap{\sf Boundary}{\Delta})\subseteq V({\sf Perimeter}{W})},$
  • if $u$ and $v$ are distinct vertices, both of which are either corners of branch vertices of $W$, such that $\pi^{-1}(u),\pi^{-1}(v)\notin N(\delta)$, then $u,v$ belong to distinct cells of $\delta$,
  • no cell ${c\in C(\delta)}$ with ${c\subseteq\Delta}$ is a vortex, and
  • ${W-{\sf Perimeter}{W}}$ is a subgraph of ${\bigcup\{\Sigma(c)\mid c\subseteq\Delta\}}.$
[…]
Let $G$ be a graph, let ${r\geq 3}$ be an integer and let $W$ be an ${r}$-wall in $G.$ If ${(A,B)}$ is a separation of $G$ of order at most ${r-1},$ then exactly one of the sets ${A\setminus B}$ and ${B\setminus A}$ includes the vertex set of a column and a row of $W.$ If it is the set $A\setminus B,$ we say that $A$ is the $W$-majority side of the separation ${(A,B)}$; otherwise, we say that $B$ is the $W$-majority side.
Let $\Sigma$ be a surface and ${\delta=(\Gamma,\mathcal{D})}$ be a $\Sigma$-decomposition of $G.$ We say that $\delta$ is $W$-central if there is no cell ${c\in C(\delta)}$ such that ${V(\Sigma(c))}$ includes the $W$-majority side of a separation of $G$ of order at most ${r-1}.$
More generally, let ${Z\subseteq V(G)}$ with ${\lvert Z\rvert\leq r-1},$ let Σsuperscriptnormal-$\Sigma^{\prime}$ be a surface, and let $\delta^{\prime}$ be a Σsuperscriptnormal-$\Sigma^{\prime}$-decomposition of ${G-Z}.$ Then $\delta^{\prime}$ is a $W$-central decomposition of ${G-Z}$ if for all separations ${(A,B)}$ of ${G-Z}$ of order at most ${r-|Z|-1}$ such that ${B\cup Z}$ is the majority side of the separation ${(A\cup Z,Z\cup B)}$ of $G,$ there is no cell ${c\in C(\delta^{\prime})}$ such that ${V(\Sigma_{\delta^{\prime}}(c))}$ contains $B.$
[…]
Now let $k$ be some positive integer and let =$\mathfrak{H}=\{\mathscr{H}^{(i)}\mid i\in[k]\}$ be a parametric family of $k$ minor-monotone parameterized graphs. Moreover, let $\mathbf{a},$ $\mathbf{g},$ $\mathbf{v}_{\mathsf{b}},$ and $\mathbf{v}_{\mathsf{d}}$ be computable functions that map elements of $\mathbb{N}^{k}$ to $\mathbb{N},$ and let $\mathbf{R}\colon\mathbb{N}^{k+1}\rightarrow\mathbb{N}$ be another computable function.
A local structure theorem is a statement $\mathfrak{S}(\mathfrak{H},\mathbf{a},\mathbf{g}_{\mathsf{o}},\mathbf{g}_{\mathsf{n}},\mathbf{v}_{\mathsf{b}},\mathbf{v}_{\mathsf{d}},\mathbf{R})$ of the following form:
Let $r,t_{1},\dots,t_{k}\geq 0$ be integers. Let $G$ be a graph and let $W$ be an $\mathbf{R}(r,t_{1},\dots,t_{k})$-wall in $G$. Then either $G$ has an $\mathscr{H}^{(j)}_{t_{j}}$-minor for some $j\in[k],$ or there exists a set $A\subseteq V(G)$ of size at most $\mathbf{a}(t_{1},\dots,t_{k}),$ a surface $\Sigma$ which is of Euler-genus at most $\mathbf{g}_{\mathsf{o}}(t_{1},\dots,t_{k})$ if it is orientable and of Euler-genus at most $\mathbf{g}_{\mathsf{n}}(t_{1},\dots,t_{k})$ otherwise, a $W$-central $\Sigma$-decomposition $\delta$ of $G-A$ of depth at most $\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})$ and breadth at most $\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k}),$ and an $r$-subwall $W^{\prime}\subseteq W-A$ which is flat in $\delta$.
We say that $\mathfrak{S}(\mathfrak{H},\mathbf{a},\mathbf{g}_{\mathsf{o}},\mathbf{g}_{\mathsf{n}},\mathbf{v}_{\mathsf{b}},\mathbf{v}_{\mathsf{d}},\mathbf{R})$ is valid, if it is a true statement.
A valid local structure theorem is $\mathbf{n}$-orderly if there exists a function $\mathbf{n}\colon\mathbb{N}^{k}\rightarrow\mathbb{N}$ such that the $\Sigma$-decomposition $\delta$ guaranteed by $\mathfrak{S}$ also satisfies
  • for each vortex cell $c\in C(\delta)$ there exists a nest $\mathcal{C}_{c}$ of order $\mathbf{n}(t_{1},\dots,t_{k})$ in $\delta$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $c,$ and
  • for each vortex cell $c\in C(\delta)$ let $\Delta_{c}\subseteq\Sigma$ be the disk bounded by the trace of $\mathbf{n}(t_{1},\dots,t_{k})\in\mathcal{C}_{c},$ then for each pair of distinct vortex cells $c,c^{\prime}\in C(\delta)$ we have that $\Delta_{c}\cap\Delta_{c^{\prime}}=\emptyset.$
4 From local to global structure theorems
To fully capture the structure of graphs excluding a surface it does not suffice to only describe the structure with respect to a single large order tangle or well-linked set. In this section we describe how a local structure theorem can be used to obtain a sructural decomposition of the whole graph.
4.1 Local structure with respect to a tangle
Before we can show that a (valid) local structure theorem $\mathfrak{S}(\mathfrak{H},\mathbf{a},\mathbf{g}_{\mathsf{o}},\mathbf{g}_{\mathsf{n}},\mathbf{v}_{\mathsf{b}},\mathbf{v}_{\mathsf{d}},\mathbf{R})$ can be used to obtain a statement for the global structure of a graph $G$ we need to introduce some additional tools. Oir first step is to show that, by sacrificing a small set of vertices in each vortex one can obtain a tangle-based version of every local structure theorem. Moreover, a tangle-based version also immediately implies a well-linked-set-based version.
Theorem 4.1.
Let $\mathfrak{S}(\mathfrak{H}=\{\langle\mathscr{H}^{(j)}_{i}\rangle_{i\in\mathbb{N}}\mid j\in[k]\},\mathbf{a},\mathbf{g}_{\mathsf{o}},\mathbf{g}_{\mathsf{n}},\mathbf{v}_{\mathsf{b}},\mathbf{v}_{\mathsf{d}},\mathbf{R})$ be a valid local structure theorem.
Let $\mathsf{p}$ be a function such that for every positive integer $r,$ every graph with treewidth at least $\mathsf{p}(r)$ contains an $r$-wall as a subgraph.
Let $t_{1},\dots,t_{k},r$ be positive integers. Let $\mathcal{T}$ be a tangle in a graph $G$ whose order is $\geq 3\mathsf{p}(6\mathbf{R}{(r_{1},t_{1},\dots,t_{k})}^{2})+1$ with
$r_{1}\coloneqq\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k})\cdot(2\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})+1)+6+\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})+r.$
Then either there exists $i\in[k]$ such that $G$ has a minor isomorphic to $H^{i}_{t_{i}},$ or there exists $A\subseteq V(G),$ where
$|A|\leq\mathbf{v}_{\mathsf{b}}(T_{1},\dots,t_{k})(2\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})+1)+\mathbf{a}(t_{1},\dots,t_{k}),$
a surface $\Sigma$ which is of Euler-genus at most $\mathbf{g}_{\mathsf{o}}(t_{1},\dots,t_{k})$ if it is orientable and of Euler-genus at most $\mathbf{g}_{\mathsf{n}}(t_{1},\dots,t_{k})$ otherwise, and a $\Sigma$-decomposition $\delta$ of breadth at most $\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k})$ and depth at most $\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})$ for $G-A$ which is $(\mathcal{T}-A)$-central.
Moreover, let $\mathcal{V}$ be the collection of vortices of $\delta,$ then there exists a set $Z\subseteq A$ of size at least $|A|-\mathbf{a}(t_{1},\dots,t_{k})$ such that $Z$ there exists a near partition ${\{Z_{c}\}}_{c\in\mathcal{V}}$ of $Z$ into sets of size at most $12\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})+$ 1 each such that for each $c\in\mathcal{V},$ the vertices of $Z_{c}$ may be drawn into $c$ to turn $\delta$ into a $\Sigma$-decomosition $\delta^{\prime}$ of $G-(A\setminus Z)$ satisfying the same bounds on breadth and depth as $\delta.$
Proof.
Since $\mathcal{T}$ is a tangle of order at least $3\mathsf{p}(6\mathbf{R}{(r_{1},t_{1},\dots,t_{k})}^{2})+1,$ we may apply LABEL:prop_tangletowall to obtain an $\mathbf{R}(r_{1},t_{1},\dots,t_{k})$-wall $W$ in $G$ such that $\mathcal{T}_{W}$ is a restriction of $\mathcal{T}.$
Now assume that $G$ does not contain $\mathscr{H}^{(i)}_{t_{i}}$ as a minor for any $i\in[k],$ then, since $\mathfrak{S}(\mathscr{H}=\{{(\mathscr{H}^{(j)}_{i})}_{i\in\mathbb{N}}\mid j\in[k]\},\mathbf{a},\mathbf{g}_{\mathsf{o}},\mathbf{g}_{\mathsf{n}},\mathbf{v}_{\mathsf{b}},\mathbf{v}_{\mathsf{d}},\mathbf{R})$ is valid, there exists a set $A^{\prime}\subseteq V(G)$ of size at most $\mathbf{a}(t_{1},\dots,t_{k}),$ a surface $\Sigma$which is of Euler-genus at most $\mathbf{g}_{\mathsf{o}}(t_{1},\dots,t_{k})$ if it is orientable and of Euler-genus at most $\mathbf{g}_{\mathsf{n}}(t_{1},\dots,t_{k})$ otherwise, a $W$-central $\Sigma$-decomposition $\delta$ of $G-A$ of depth at most $\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})$ and breadth at most $\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k}),$ and an $r_{1}$-subwall $W_{1}\subseteq W-A$ which is flat in $\delta.$
By applying Subsection 3.1 to $G-A^{\prime},$ the tangle $\mathcal{T}-A^{\prime},$ the wall $W_{1},$ and the $\Sigma$-decomposition $\delta$ we obtain a set $A^{\prime\prime}\subseteq V(G)\setminus A^{\prime}$ with
$|A^{\prime\prime}|\leq\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k})(2\mathbf{v}_{\mathsf{d}}+1),$
and a $\Sigma$-decomposition $\delta^{\prime}$ of breadth at most $\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k})$ and depth at most $\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})$ which is $\mathcal{T}-A^{\prime}-A^{\prime\prime}$-central. With $A\coloneqq A^{\prime}\cup A^{\prime\prime}$ the first part of the proof is complete.
To conclude, observe that, since $A^{\prime\prime}$ stems from the vortices of $\delta$ by Subsection 3.1, the set $A^{\prime\prime}$ can be partitioned into sets $Z_{1},\dots,Z_{\ell},$ where $\ell\leq\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k})$ with $|Z_{i}|\leq 2\mathbf{v}_{\mathsf{d}}+1$ for each $i\in[\ell],$ such that $\delta$ can be obtained from $\delta^{\prime}$ by drawing, for each $i\in[\ell],$ the vertices of $Z_{i}$ into different vortices of $\delta.$
4.2 The global structure theorem
We are now ready to provide a transition from a local structure theorem to a global one.
Theorem 4.2.
Let $\mathfrak{S}(\mathfrak{H},\mathbf{a},\mathbf{g}_{\mathsf{o}},\mathbf{g}_{\mathsf{n}},\mathbf{v}_{\mathsf{b}},\mathbf{v}_{\mathsf{d}},\mathbf{R})$ be a valid local structure theorem.
Let $\mathsf{p}$ be the function from Theorem 3.1. Let $t_{1},\dots,t_{k},r$ positive integers and assume $r\geq 3$ and let
$\alpha$ $\coloneqq 2\mathsf{p}(6\mathbf{R}(r_{1},t_{1},\dots,t_{k})^{2})+1\text{, with}$
$r_{1}$ $\coloneqq\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k})\cdot(2\mathbf{v}_{\mathsf{d}}+1)+6+\mathbf{v}_{\mathsf{d}}+r$
Then for every graph $G$ and every integer $r\geq 3,$ either
  1. 1.
    $G$ has a minor isomorphic to $\mathscr{H}^{(j)}_{t_{j}}$ for some $j\in[k],$ or
  2. 2.
    $G$ has a tree decomposition $(T,\beta)$ of adhesion less than $\alpha$ such that for every $d\in V(T),$ either $|\beta(d)|\leq 4\alpha,$ or there exist a set $A_{d}\subseteq\beta(d)$ of size at most $4\alpha$ and a surface $\Sigma$ which is of Euler-genus at most $\mathbf{g}_{\mathsf{o}}(t_{1},\dots,t_{k})$ if it is orientable and of Euler-genus at most $\mathbf{g}_{\mathsf{n}}(t_{1},\dots,t_{k})$ otherwise, such that, for the torso $G_{d}$ of $G$ at $d,$ $G_{d}-A_{d}$ has a $\Sigma$-decomposition $\delta$ of width at most $\alpha$ and breadth at most $\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k})$ such that there exists an $r$-wall $W_{d}$ which is flat in $\delta,$ and there is no vertex of $G_{d}-A_{d}$ which is drawn in the interior of a non-vortex-cell of $\delta.$
    Moreover, let $X_{d}$ be the collection of all vertices of $G_{d}-A_{d}$ which are drawn in some vortex of $\delta$ together with the vertices of $A_{d}.$ Then for every neighbour $d^{\prime}$ of $d$ in $T$ we have that
    • i.
      $|(\beta(d)\cap\beta(d^{\prime}))\setminus X_{d}|\leq 3$ and
    • ii.
      $G_{d}-X_{d}$ is a minor of $G.$
Moreover, given an oracle for finding the $\Sigma$-decomposition corresponding to the local structure theorem when given a large enough wall, there exists an algorithm that either finds a minor isomorphic to $\mathscr{H}^{(j)}_{t_{j}}$ for some $j\in[k],$ or produces the second outcome of the statement above in time $\mathcal{O}_{t}(|V(G)|\cdot(Q+|V(G)|^{2}|E(G)|\log(|V(G)|))),$ where $Q$ is the running time of the oracle as above and $t\coloneqq\max_{i\in[k]}t_{i}$.
Proof.
We prove that the following, slightly stronger version of the first claim holds by induction on the number of vertices in $G.$
Strengthened claim: Let $G$ be a graph which does not have a minor isomorphic to $\mathscr{H}^{(j)}_{t_{j}}$ for any $j\in[k]$ and let $A\subseteq V(G)$ be a set of size at most $3\alpha.$ Then $G$ has a tree decomposition $(T,\beta)$ with root $f,$ where $A\subseteq\beta(f),$ and which is of adhesion less than $\alpha$ such that for every $d\in V(T)$ either
  • 1.
    $|\beta(d)|\leq 4\alpha,$ or
  • 2.
    there exist a set $A_{d}\subseteq\beta(d)$ of size at most $4\alpha,$ a and a surface $\Sigma$ which is of Euler-genus at most $\mathbf{g}_{\mathsf{o}}(t_{1},\dots,t_{k})$ if it is orientable and of Euler-genus at most $\mathbf{g}_{\mathsf{n}}(t_{1},\dots,t_{k})$ otherwise, such that, for the torso $G_{d}$ of $G$ at $d$, $G_{d}-A_{d}$ has a $\Sigma$-decomposition $\delta$ of width at most $\alpha$ and breadth at most $\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k}).$ Moreover, let $X_{d}$ be the collection of all vertices of $G_{d}-A_{d}$ which are drawn in some vortex of $\delta$ together with the vertices of $A_{d}.$ Then for every neighbour $d^{\prime}$ of $d$ in $T$ we have that
    • i.
      $|(\beta(d)\cap\beta(d^{\prime}))\setminus X_{d}|\leq 3$ and
    • ii.
      $G_{d}-X_{d}$ is a minor of $G.$
Assume the strengthened claim is false. We choose $G,$ $A,$ and $t_{1},\dots,t_{k},r$ to form a counterexample minimizing $|V(G)|+|V(G)\setminus A|.$ It follows that $|V(G)|>4\alpha$ since otherwise we could choose $(T,\beta)$ to be a trivial tree decomposition with a single node and a unique bag containing all of $V(G).$ Moreover, we have $|A|=3\alpha$ since otherwise we could add an arbitrary vertex of $V(G)\setminus A$ to $A$ and thereby contradict our choice for the counterexample.
Notice that there exists no separation $(X_{1},X_{2})$ of order less than $\alpha$ such that $|X_{i}\cap A|\leq|X_{1}\cap X_{2}|$ for both $i\in[2]$ as $|A|=3\alpha.$
Moreover, suppose that $|X_{i}\cap A|>|X_{1}\cap X_{2}|$ for both $i\in[2].$ Let $A_{i}\coloneqq(X_{i}\cap A)\cup(X_{1}\cap X_{2})$ and observe the following:
$|A\cap X_{i}|$ $=3\alpha-|A\cap X_{3-i}|+|A\cap X_{1}\cap X_{2}|$
$<3\alpha-|X_{1}\cap X_{2}|+|A\cap X_{1}\cap X_{2}|$
$=3\alpha-|(X_{1}\cap X_{2})\setminus A|.$
It now follows that $|A_{i}|\leq 3\alpha$ for both $i\in[2].$ Also notice that $|A\cup(X_{1}\cap X_{2})|\leq 4\alpha.$ Let us introduce a vertex $f$ and set $\beta(f)\coloneqq(X_{1}\cap X_{2})\cup A.$ Now let $G_{i}\coloneqq G[X_{i}]$ and observe that $G_{i},$ $A_{i},$ and $t_{1},\dots,t_{k},r$ do not form a counterexample by our minimality assumption. Hence, there exists a tree decomposition $(T_{i},\beta_{i})$ with root $f_{i}$ and $A_{i}\subseteq\beta(f_{i})$ which satisfies all requirements of the strengthened claim. We may form a tree decomposition $(T,\beta)$ by joining $f$ to the vertices $f_{1}$ and $f_{2},$ thereby creating the tree $T.$ It is straight forward to check that $(T,\beta)$ meets all requirements of the strengthened claim which is a direct contradiction to the choice of our counterexample. Observe that it is possible to determine if such a decomposition exists, at least with a slightly worse approximation factor, by using the techniques from [4].
Hence, in particular, we obtain that for each separation $(X_{1},X_{2})$ of order at most $\alpha-1$ there exists a unique $i\in[2]$ such that $|X_{i}\cap A|>|X_{1}\cap X_{2}|.$ This immediately implies that $A$ is a $(\alpha,\frac{2}{3})$-linked set. Hence $\mathcal{T}\coloneqq\mathcal{T}_{A}$ is a tangle of order $\alpha=2\mathsf{p}(6\mathbf{R}{(r_{1},t_{1},\dots,t_{k})}^{2})+1.$ We may now use Theorem 3.1 to find a large wall whose associated tangle is a truncation of $\mathcal{T}_{A}$ in time $2^{\mathcal{O}(\mathsf{poly}(t)^{c_{2}})}|V(G)|^{2}|E(G)|\log(|V(G)|)$. By Theorem 4.1 and the fact that $G$ does not contain $\mathscr{H}^{j}_{t_{j}}$ as a minor for any $j\in[k],$ we obtain a set $A_{f}\subseteq V(G)$ where
$|A_{f}|\leq\mathbf{v}_{\mathsf{b}}(T_{1},\dots,t_{k})(2\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})+1)+\mathbf{a}(t_{1},\dots,t_{k}),$
a surface $\Sigma$ which is of Euler-genus at most $\mathbf{g}_{\mathsf{o}}(t_{1},\dots,t_{k})$ if it is orientable and of Euler-genus at most $\mathbf{g}_{\mathsf{n}}(t_{1},\dots,t_{k})$ otherwise, and a $\Sigma$-decomposition $\delta$ of breadth at most $\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k})$ and depth at most $\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})$ for $G-A_{f}$ which is $(\mathcal{T}-A_{f})$-central.
What remains is to show how, for each cell $c$ of $\delta$ which has more than two vertices on its boundary and at least one vertex in its interior, we obtain a smaller graph with a bounded size set that will take the place of $A.$ For each such graph, as it does not form a counterexample, we then obtain a decomposition meeting the requirements of our strengthened claim. By combining all of these tree decompositions we will finally find a decomposition for $G$ and thereby contradict the existence of a counterexample.
Before we do this, however, let $d\coloneqq 2\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})$ and $a\coloneqq\mathbf{a}(t_{1},\dots,t_{k}).$ Now observe that $a+d+3<\mathbf{R}{(r_{1},t_{1},\dots,t_{k})}^{2}$ by the choice of $r_{1}$ and the fact that the deletion of $a$ vertices of an $\mathbf{R}(r_{1},t_{1},\dots,t_{k})$-wall still leaves an $r$-wall intact. It follows that $a+d+3<\frac{1}{2}\alpha$ and thus, for any separation $(X,Y)$ of $G$ of order at most $a+d+3$ we have $(X,Y)\in\mathcal{T}$ or $(Y,X)\in\mathcal{T}.$ In case $(X,Y)\in\mathcal{T},$ we also get, from the definition of $\mathcal{T},$ that $|(A\cap X)\cup(X\cap Y)|<\alpha.$ This observation will allow us to obtain tree decompositions as desired for all part of the graph we are about to separate from the graph drawn on the boundaries of the cells of $\delta.$
Let $X_{f}$ be the collection of all vertices of $G-A_{f}$ which are drawn in some vortex of $\delta$ together with the vertices of $A_{f}.$
Let $\mathcal{V}$ be the collection of all vortices of $\delta.$ For each $c\in\mathcal{V}$ let $(G_{c},\Omega_{c})$ be the corresponding vortex society, and let $n_{c}\coloneqq|V(\Omega_{c})|.$ Let us fix a linear order $v^{c}_{1},\dots,v^{c}_{n_{c}}$ of $V(\Omega_{c})$ such that there exists a linear decomposition $\langle L^{c}_{1},\dots,L^{c}_{n_{c}},v^{c}_{1},\dots,v^{c}_{n_{c}}\rangle$ of adhesion at most $\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})$ of $(G_{c},\Omega_{c}).$ We set $Y^{c}_{i}\coloneqq(L^{c}_{i}\cap L^{c}_{i-1})\cup(L^{c}_{i}\cap L^{c}_{i+1})\cup\{v^{c}_{i}\}$ where $L^{c}_{0}=L^{c}_{i+1}=\emptyset.$ Observe that every vertex drawn on the closure of $c$ must belong to some $L^{c}_{i},$ moreover, we have that $\bigcup_{i=1}^{n_{c}}L^{c}_{i}\subseteq X_{f}.$ Finally, let $A_{c}\coloneqq Y^{c}_{i}\cup A_{f}\cup(L^{c}_{i}\cap A).$
For every $c\in\mathcal{V}$ and $i\in[n_{c}]$ let $H^{c}_{i}\coloneqq G[L^{c}_{i}\cup A_{f}].$ Notice that $(V(G)\setminus(L^{c}_{i}\setminus Y^{c}_{i}),L^{c}_{i}\cup A_{f})\in\mathcal{T}$ is a separation of order at most $d+a+1$ in $G$ and $H^{c}_{i}-A_{f}=G_{c}[Y^{c}_{i}].$ Moreover, it follows from the discussion above that $|A_{c}|<\alpha,$ and thus $(V(G)\setminus(L^{c}_{i}\setminus(Y^{c}_{i}\cup A)),L^{c}_{i}\cup A_{f})$ is a separation of order less than $\alpha$ in $G.$ Notice that htis argument also yields the, bound $\alpha$ on the width of the linear decomposition $\langle L^{c}_{1},\dots,L^{c}_{n_{c}}v^{c}_{1},\dots,v^{c}_{n_{c}}\rangle.$
By the minimality of our counterexample, we obtain a tree decomposition $(T^{c}_{i},\beta^{c}_{i})$ with root $f^{c}_{i}$ satisfying the requirements of the strengthened claim for every $H^{c}_{i},$ $A_{c},$ and $t_{1},\dots,t_{k},r.$
Now let $\mathcal{C}$ be the collection of all non-vortex cells of $\delta.$ For every $c\in\mathcal{C}$ let $H^{c}$ be the subgraph of $G$ with vertex set $V(\Sigma(c))\cup A_{f}$ and edge set $E(\Sigma(c))$ together with all edges that have at least one endpoint in $A_{f}$ and the other in $V(H^{c}).$ Let $N^{c}$ be the set of vertices of $G$ which are drawn on the boundary of the closure of $c.$ Let $\Omega^{c}$ be the cyclic permutation of the vertices of $N\setminus A$ obtained by following along the boundary of $c$ clockwise. By deleting $A$ in addition to $A_{f}$ it is possible that we slightly changed the way $H^{c}$ attaches to the rest of $G.$ To fix this, we apply Theorem 3.2 to $(H^{c}-A,\Omega^{c})$ and enhance $\delta$ by possibly refining $c$ into smaller cells wherever possible. Note that the application of Theorem 3.2 guarantees that the formation of a torso by removing the interiors of non-vortex cells yields a minor of $G.$ From now on, we will assume that all cells have been refined in this way.
Observe that $(V(G)\setminus(V(\Sigma(c))\setminus N^{c}),V(H^{c}))$ is a separation of $G$ of order at most $a+3<\frac{1}{2}\alpha.$ By the discussion above and the fact that $\delta$ is $(\mathcal{T}-A_{f})$-central, it follows that $(V(G)\setminus(V(\Sigma(c))\setminus N^{c}),V(H^{c}))\in\mathcal{T}$ and $|A\cap V(H^{c})|\leq a+3<\frac{1}{2}\alpha.$ Hence, $(V(G)\setminus(V(\Sigma(c))\setminus(N^{c}\cup A)),V(H^{c}))$ is a separation of order less than $\alpha$ in $G.$
Let $H^{c}_{1},\dots H^{c}_{n_{c}}$ be the components of $H^{c}-A-A_{f}.$ Let $A^{c}_{i}\coloneqq A_{f}\cup(V(H^{c})\cap A)\cup(V(H^{c}_{i})\cap(V(G)\setminus(V(\Sigma(c))\setminus N^{c}))).$ For each $c\in\mathcal{C}$ and $i\in[n_{c}]$ we obtain, from the minimality of our counterexample, a tree decomposition $(T^{c}_{i},\beta^{c}_{i})$ with root $f^{c}_{i}$ for $H^{c}_{i},$ $A^{c}_{i},$ and $t_{1},\dots,t_{k},$ which satisfies the requirements of the strengthened claim.
We may now construct a tree decomposition for $G$ as follows. Let $f$ be a vertex, let $X_{f}\coloneqq A\cup A_{f}\cup\bigcup_{c\in\mathcal{V}}\bigcup_{i=1}^{n_{c}}Y^{c}_{i}$ and set
$\beta(f)\coloneqq V(G)\setminus((\bigcup_{c\in\mathcal{C}\cup\mathcal{V}}\bigcup_{i\in[n_{c}]}V(H^{c}_{i}))\cup\bigcup_{c\in\mathcal{C}\cup\mathcal{V}}\bigcup_{i\in[n_{c}]}A^{c}_{i}).$
It follows from the construction of the $A^{c}_{i}$ and the definition of $\delta$ that, for the torso $G_{f}$ of $G$ at $f$,$G_{f}-A-A_{f}$ has a $\Sigma$-decomposition $\delta^{\prime}$ of depth at most $\mathbf{v}_{\mathsf{d}}(t_{1},\dots,t_{k})$ and breadth at most $\mathbf{v}_{\mathsf{b}}(t_{1},\dots,t_{k}).$ Moreover, $G_{f}-X_{f}$ is a minor of $G$ since $X_{f}$ contains all vertices which are drawn in the vortices of $\delta^{\prime}$ together with the apex set $A\cup A_{f}$ and anything that attaches to this graph does so in separations of order at most three which are obtained from applications of Theorem 3.2.
For every $c\in\mathcal{V}\cup\mathcal{C}$ and every $i\in[n_{c}]$ introduce the edge $f^{c}_{i}f.$ And enhance $\beta$ by $\beta^{c}_{i}.$ Let $(T,\beta)$ be the resulting tree decomposition with root $f.$ We conclude by observing that $(T,\beta)$ satisfies the requirements of the strengthened claim and thus contradicts the choice of our counterexample. ∎
4.3 The local structure of a surface
The final piece of technology we will need is to describe how Kawarabayashi, Thomas, and Wollan capture a surface in their version of the local structure theorem [32].
The purpose of the following definitions is to introduce the concept of $\Sigma$-configurations which are meant as a strengthening of $\Sigma$-decompositions in a way that makes it possible to describe how a clique minor might appear during the inductive process of constructing the surface in the proof of Subsection 4.3. Please note that we will not make full use of the infrastructure provided by a $\Sigma$-configuration. This is, because a part of it is designed to be used within an induction where the surface is built by iteratively finding large transactions.
[…]
Let $\rho=(\Gamma,\mathcal{D})$ be a rendition of a society $(G,\Omega)$ in a surface $\Sigma$ and let $\Delta\subseteq\Sigma$ be an arcwise connected set. A nest in $\rho$ around $\Delta$ of order $s$ is a sequence $\mathcal{C}=(C_{1},C_{2},\dots,C_{s})$ of disjoint cycles in $G$ such that each of them is grounded in $\rho,$ and the trace of $C_{i}$ bounds a closed disk $\Delta_{i}$ in such a way that $\Delta\subseteq\Delta_{1}\subsetneq\Delta_{2}\subsetneq\dots\subsetneq\Delta_{s}\subseteq\Sigma.$
Let $(G,\Omega)$ be a society and $\rho=(\Gamma,\mathcal{D})$ be a rendition of $(G,\Omega)$ in a surface $\Sigma.$ Let $C$ be a cycle in $G$ that is grounded in $\rho$ such that the trace $T$ of $C$ bounds a closed disk $D$ in $\Sigma.$ Let $L$ be the subgraph of $G$ with vertex set $\pi_{\rho}(N(\rho)\cap T)$ and no edges. We define the outer graph of $C$ in $\rho$ as the graph
$L\cup\bigcup_{c\in C(\rho)\text{ and }c\not\subseteq D}\Sigma(c)$
We define the inner graph of $C$ in $\rho$ as the graph
$L\cup\bigcup_{c\in C(\rho)\text{ and }c\subseteq D}\Sigma(c)$
Finally, we define the inner society of $C$ in $\rho$ as follows. Let $\Omega^{\prime}$ be the cyclic permutation of $V(L)$ obtained by traversing $T$ clockwise. If $G^{\prime}$ is the inner graph of $C$ in $\rho,$ then the inner society of $C$ in $\rho$ is the society $(G^{\prime},\Omega^{\prime}).$
Let $\mathcal{C}=\{C_{1},\dots,C_{s}\}$ be a collection of pairwise vertex-disjoint cycles and $\mathcal{P}=\{P_{1},\dots,P_{k}\}$ be a linkage. We say that $\mathcal{P}$ is orthogonal to $\mathcal{C}$ if every $P_{i}$ has a non-empty intersection with each cycle of $\mathcal{C}$ such that there exists an endpoint $p_{i}$ of $P_{i}$ for which the cycles $C_{1},\dots,C_{s}$ appear in the order listed when traversing along $P_{i}$ starting from $p_{i}$, and $P_{i}\cap C_{j}$ is a (possibly trivial) subpath of $P_{i}$ for all $i\in[k]$, $j\in[s]$.
Figure 11: An illustration of a pair $(\mathcal{C},\mathcal{P})$ overlaying a wall $W$.
Let $G$ be a graph, $W$ a wall in $G,$ $\mathcal{C}$ a nest in $G,$ and $\mathcal{P}$ a linkage which is orthogonal to $\mathcal{C}.$
We say that $(\mathcal{P},\mathcal{C})$ overlays $W$ if we can fix the horizontal paths of $W$ to be $H_{1},\dots,H_{k}$ and the vertical paths $V_{1},\dots,V_{\ell}$ such that there exist injective maps $\phi_{\text{hor}}\colon\mathcal{C}\to\{H_{1},\dots,H_{k}\}$ and $\phi_{\text{vert}}\colon\mathcal{P}\to\{V_{1},\dots,V_{\ell}\}$ such that for all $P\in\mathcal{P}$ and $C\in\mathcal{C},$ $\phi_{\text{hor}}(C)\cap\phi_{\text{vert}}(P)\subseteq P\cap C.$ See Figure 11 for an illustration.
Let $C$ be some cycle grounded in a $\Sigma$-decomposition $\delta$ of a graph $G$. We say that a linkage $\mathcal{L}$ is coterminal with a linkage $\mathcal{P}$ up to level $C$ if there exists $\mathcal{P}^{\prime}\subseteq\mathcal{P}$ such that $H\cap\mathcal{L}=H\cap\mathcal{P^{\prime}}$ where $H$ denotes the outer graph of $C$. When $\mathcal{C}=\{C_{1},\dots,C_{\ell}\}$ is a nest in $\delta$, and it is clear from the context to which nest we are referring, we will abbreviate “coterminal with $\mathcal{P}$ up to level $C_{i}$” by “coterminal with $\mathcal{P}$ up to level $i$”.
[…]
Let $G$ be a graph, let $\Sigma=\Sigma^{(h,c)}$ be a surface, and let $\delta=(\Gamma,\mathcal{D})$ be a $\Sigma$-decomposition of $G.$ Let $W$ be a wall in $G,$ $W_{1}$ a subwall of $W,$ and $W_{0}$ an $r$-subwall of $W_{1}.$ Assume that there exist cells $c_{0},c_{1}\in C(\delta)$ and a closed disk $\Delta_{0}\subseteq\Sigma$ such that:
  • No cell in $C(\delta)\setminus\{c_{0},c_{1}\}$ is a vortex.
  • The vortex society $(G_{0},\Omega_{0})$ of $c_{0}$ in $\delta$ has a vortex-free rendition, $W_{0}\subseteq G_{0},$ and there exists a vortex free rendition of $(G_{0},\Omega_{0})$ for which $W_{0}$ is flat.
  • The cell $c_{0}\subseteq\Delta_{0}\subseteq\Sigma\setminus c_{1}$ and the boundary of $\Delta_{0}$ intersect no cell of $\delta.$
  • The complementary society $(G_{2},\Omega_{2})$ to the vortex society of $c_{0}$ has a cylindrical rendition $\rho_{2}=(\Gamma_{2},\mathcal{D}_{2},c_{2})$ with a nest $\mathcal{C}_{2},$ where $\Gamma_{2}=\Gamma,$ $C(\rho_{2})\setminus c_{2}=\{c\in C(\delta)\mid c\subseteq\Delta_{0}\}\setminus\{c_{0}\},$ $N(\rho_{2})=N(\delta)\cap\Delta_{0},$ $\Sigma_{\rho_{2}}(c_{2})=\bigcup_{c\not\subseteq\Delta_{0}}\Sigma_{\delta}(c),$ and $\Sigma_{\delta}(c)=\Sigma_{\rho_{2}}(c)$ for every $c\in C(\rho_{2})\setminus\{c_{2}\}.$
  • The vortex society of $c_{1}$ in $\delta$ has a cylindrical rendition $\rho_{1}=(\Gamma_{1},\mathcal{D}_{1},c_{1}^{\prime})$ with a nest $\mathcal{C}_{1}.$
Let $\widetilde{\Delta_{0}}$ denote the set of vertices of $N(\rho)$ on the boundary of $\Delta_{0}.$ Let $X_{1},X_{2}$ be two complementary segments of $\Omega_{2}.$ Assume that there exists a set $\gamma=\{(\mathcal{L}_{i},X_{i}^{2})\mid i\in[c+h]\}$ and disjoint linkages $\mathcal{P}_{1},\mathcal{P}_{2}$ in $(G_{2},\Omega_{2})$ such that:
  • $\mathcal{P}_{1}$ is a linkage from $X_{1}$ to $\Sigma_{\rho_{1}}(c_{1}^{\prime})$ which is orthogonal to both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ and $(\mathcal{P}_{1},\mathcal{C}_{1})$ overlays $W_{1}.$
  • The linkage $\mathcal{P}_{2}$ is a linkage from $X_{2}$ to $\widetilde{\Delta_{0}}$ which is orthogonal to $\mathcal{C}_{2}$ and $(\mathcal{P}_{1}\cup\mathcal{P}_{2},\mathcal{C}_{2})$ overlays $W_{1}.$
  • For $i\in[c+h],$ $\mathcal{L}_{i}$ is a transaction on $(G_{2},\Omega_{2})$ and $X^{2}_{i}\subseteq X_{2}$ is a segment of $\Omega_{2}.$ The transactions $\mathcal{L}_{1},\dots,\mathcal{L}_{c+h}$ are pairwise disjoint with endpoints in $X_{i}^{2}$ and, moreover, $X_{1}^{2},\dots,X^{2}_{c+h}$ are also pairwise disjoint. For all $i,$ $\mathcal{L}_{i}$ is coterminal with $\mathcal{P}_{2}$ up to level $1$ of $\mathcal{C}_{2}$ in $\rho_{2}$ and $\mathcal{L}_{i}$ is disjoint from $\Sigma_{\delta}(c_{1})$ and disjoint from $\mathcal{P}_{1}.$
  • The transactions $\mathcal{L}_{1},\dots,\mathcal{L}_{h}$ are handle transactions, each of thickness $θ,$ and the transactions $\mathcal{L}_{h+1},\dots,\mathcal{L}_{c+h}$ are crosscap transactions, each of thickness $θ.$
  • Let $\mathcal{P}_{1}=\{P_{1},\dots,P_{m}\},$ for $i\in[m]$ let $x_{i}$ be the endpoint of $P_{i}$ in $X_{1}$ and let $y_{i}$ be the last vertex of $P_{i}$ in $\widetilde{c_{1}}.$ If $x_{1},\dots,x_{m}$ appear in $\Omega_{2}$ in the order listed, then $y_{1},\dots,y_{m}$ appear on $\widetilde{c_{1}}$ in the order listed.
In those circumstances (neglecting the wall $W$) we say that $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ is a $\Sigma$-configuration with parameters $(r,|\mathcal{C}_{1}|,|\mathcal{C}_{2}|,|\mathcal{P}_{1}|,θ,c,h).$ The cell $c_{1}$ is called the exceptional cell of $\delta$.
For an illustration of a $\Sigma$-configuration consider (a), (b), and (c) in Figure 12.
Figure 12: A $\Sigma$-configuration of a graph (a), (b), and (c)) together with a rendition of bounded breadth and depth of its exceptional cell (d)). (a) The vortex-free rendition of the vortex society of $c_{0}$ together with the wall $W_{0}$ which is grounded in this rendition. On the boundary of $c_{0}$ one can find one endpoint of every path in $\mathcal{P}_{1}$. Notice that each such endpoint coincides with a branch vertex of $W_{0}$. (b) A rendition of the complementary society $(G_{2},\Omega_{2})$ to the vortex society of $c_{0}$. This rendition contains the nest $\mathcal{C}_{2}$ of the special cell $c_{2}$, a part of the linkage $\mathcal{P}_{1}$ which connects $W_{0}$ to the boundary of $c_{1}^{\prime}$, the transactions $\mathcal{L}_{i}$ which are either handle or crosscap transactions on $(G_{2},\Omega_{2})$ and reach into the interior of $c_{2}$. The linkage $\mathcal{P}_{2}$ (not highlighted) consists of the subpaths of the members of the $\mathcal{L}_{i}$ connecting the boundary of $c_{1}$ to the boundary of $c_{2}$. Finally, the figure depicts the exceptional cell $c_{1}$. (c) The cylindrical rendition of the vortex society of $c_{1}$, its vortex $c_{1}^{\prime}$, its nest $\mathcal{C}_{1}$, and the remaining part of $\mathcal{P}_{1}$. (d) An alternative rendition of the vortex society of $c_{1}$, this time a rendition of bounded breadth and depth where each vortex is surrounded by a private nest and each such nest together with its vortex lives in a private disk. This rendition is guaranteed to exist due to Subsection 4.3.
[…]
Let $r,s_{2},M,t\geq 1$ be integers with $M\geq t,$ $s_{2}\geq 2t.$ Let $\mu=\left\lfloor\frac{1}{10^{5}t^{26}}t^{10^{7}t^{26}}\cdot M\right\rfloor.$ Let
$R=49152t^{24}(r+2s_{2})+\mu.$
Let $G$ be a graph, and let $W$ be an $R$-wall in $G.$ Then either $G$ contains $K_{t}$ as a minor, or there exist integers $c,h\geq 0$ with $c+h\leq{t+1\choose 2},$ a set $A\subseteq V(G)$ of size at most $\mu,$ a surface $\Sigma=\Sigma^{(h,c)}$, and a $\Sigma$-configuration $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,c,h)$.
Moreover, the vortex society of $c_{1}$ has a rendition $\rho_{1}=(\Gamma_{1},\mathcal{D}_{1})$ of breadth at most $2t^{2}$ and depth at most $\mu$ and such that
  • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $c,$ and
  • for each vortex cell $c\in C(\rho_{1})$ let $\Delta_{c}\subseteq\Sigma$ be the disk bounded by the trace of $C_{M}\in\mathcal{C}_{c},$ then for each pair of distinct vortex cells $c,c^{\prime}\in C(\rho_{1})$ we have that $\Delta_{c}\cap\Delta_{c^{\prime}}=\emptyset.$
For an illustration of the outcome of Subsection 4.3 see (a), (b), and (d) in Figure 12. Please notice that the structure of the nests for the vortices of $\rho_{1}$ is not made explicit in [32]. However, the existence of these nests follows from the proofs as explained, for example, in [47].
5 The upper bound
We are finally ready to prove the upper bound portion of Theorem 1.2. To do this we will first show how to extract a Dyck-grid of chosen order from a $\Sigma$-configuration. Then, we use this technique to prove a local structure theorem for the case where $\Sigma$ belongs to a fixed closed set of surfaces.
5.1 A local structure theorem for excluding Dyck-grids
[…]
Let $c,h,t,r,s_{2}$ be non-negative integers. In case $c=0$ let $c_{0}=h_{0}=0,$ if $c=2p+1$ let $c_{0}=1$ and $h_{0}=p,$ and if $c=2p\neq 0$ let $c_{0}=2$ and $h_{0}=p-1.$ Finally, let $g=2h+c.$ Assume that
$r\geq 16t\text{, }M\geq(162^{2g}+4)t,\text{ and }s_{2}\geq(162^{2g}+4)t.$
If $G$ is a graph with a $\Sigma$-configuration $(W_{1},W_{0},\delta,c_{0},c_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ with parameters $(r,M,s_{2},M,M,c,h)$ where $\Sigma=\Sigma^{(h,c)}$ then $G$ contains an $\mathscr{D}^{h+h_{0},c_{0}}_{t}$-expansion which is grounded in $\delta$.
Proof.
The claim follows quickly from the assumptions on the numbers and the definition of $\Sigma$-configurations.
Let us first describe how to find a mixed surface grid in the $\Sigma$-configuration. Consider the $\Sigma$-configuration $(W_{1},W_{0},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma).$
In case $c+h=0$ we may conclude immediately as $W_{0}$ is an $r$-wall where $r\geq 16t.$ Hence, $W_{0}$ contains the $(t,4t)$-cylindrical grid, and therefore $\mathscr{D}_{t}^{(0,0)},$ as a minor. Moreover, $W_{0}$ is grounded in $\delta$.
So from now on we may assume that $c+h\geq 1.$ We first collect the cycles necessary for the cylindrical grid which will host the crosscap and handle transactions. Let us number the cycles of $\mathcal{C}_{2}$ as $C_{1},C_{2},\dots,C_{s_{2}}$ as in the definition of a nest. Let $H_{0}$ be the subgraph of $G$ consisting of the union of the cycles $C_{1},\dots,C_{162^{2g}t}.$ For each $i\in[c+h]$ let $Q_{i}$ be the subgraph of $G$ obtained as follows. For each $L\in\mathcal{L}_{i}$ there exist paths $P_{1},P_{2}\in\mathcal{P}_{1}$ such that $L$ is coterminal with $\{P_{1},P_{2}\}$ up to level $162^{2g}t.$ For each $j\in[2],$ let $x_{j}$ be the last vertex of $P_{j}$ on $C_{162^{2g}t}$ when traversing along $P_{j}$ starting on $V(\Omega_{2}).$ Moreover, let $y_{j}$ be the last vertex of $P_{j}$ in $C_{1}\cap L$ when traversing along $P_{j}$ starting on $x_{j}$ and moving towards $\Sigma_{\rho_{1}}(c_{1}^{\prime}).$ Let $L^{\prime}$ be the subpath of $L$ with endpoints $y_{1}$ and $y_{2}.$ Finally, we define $\hat{L}$ to be the path $x_{1}P_{1}y_{1}L^{\prime}y_{2}P_{2}x_{2}.$ Observe, that both $P_{1}$ and $P_{2}$ are orthogonal with $\{C_{1},\dots,C_{162^{2g}t}\},$ and $L^{\prime}$ is internally disjoint from $H_{0}.$ Then $Q_{i}$ is defined as the union of the paths $\hat{L}$ for all $L\in\mathcal{L}_{i}.$ Now, let $H_{1}$ be the union of $H_{0}$ with the graphs $Q_{i}$ for all $i\in[c+h].$
Observe that $H_{1}$ contains a mixed surface grid of order $162^{2g}t$ with $c$ crosscaps and $h$ handles as a minor since each of the linkages $\mathcal{L}_{i}$ has order at least $(162^{2g}+4)t$ which allows forgetting $4t$ of the consecutive $L^{\prime}$-paths.
Finally, notice that every cycle in $\mathcal{C}_{2}$ is grounded in $\delta$. Moreover, any path $\hat{L}$ as above closes two cycles with $C_{162^{2g}t}$, each of which is grounded in $\delta$. It follows that $H_{1}$ is grounded in $\delta$ and thus, the claim now follows from Theorem 2.1. ∎
We now have everything in place to prove the local structure theorem for excluding Dyck-grids. The statement we prove here is more general than the one required Theorem 1.1 as we want to allow for the exclusion of any pair of incomparable surfaces. That is, in the following we fix an Euler-genus $g_{1}=2h_{1}$ and an Euler-genus $g_{2}=c_{2}+2h_{2}$ where $c_{2}\in[2]$ and $g_{1}$ represents a surface of orientable genus $h_{1}$ while $g_{2}$ represents a non-orientable surface. To ensure that the two surfaces are incomparable, we also require $0\leq h_{2}<h_{1}$ since otherwise we could find a Dyck-grid with $h_{1}$ handles as a minor in any Dyck-grid with $h_{2}$ handles and $c_{2}$ crosscaps.
Theorem 5.1.
Let $k,h_{1},h_{2},c_{2}$ be non-negative integers where $c_{2}\in[0,2]$ and $0\leq h_{2}<h_{1}.$ Let $g_{1}\coloneqq 2h_{1},$ $g_{2}\coloneqq c_{2}+2h_{2}$ if $c_{2}\neq 0$ and $g_{2}\coloneqq 0$ otherwise, and $g\coloneqq\max\{g_{1},g_{2}\}$. Moreover, fix $t\coloneqq 12(162^{2g}+4)k.$
Define
$\mathbf{a}_{g_{1},g_{2}}$ $\coloneqq t^{10^{7}t^{26}}$
$\mathbf{g}_{\mathsf{o},g_{1},g_{2}}$ $\coloneqq g_{1}-1$
$\mathbf{g}_{\mathsf{n},g_{1},g_{2}}$ $\coloneqq g_{2}-1\text{ if }c_{0}\neq 0\text{, and }g_{1}\text{ otherwise}$
$\mathbf{v}_{\mathsf{b},{g_{1},g_{2}}}$ $\coloneqq 2t^{2}$
$\mathbf{v}_{\mathsf{d},{g_{1},g_{2}}}$ $\coloneqq t^{10^{7}t^{26}}$
$\mathbf{R}_{g_{1},g_{2}}$ $\coloneqq 49152t^{24}(r+16t)+\left\lceil\frac{5}{2}t^{10^{7}t^{26}}\right\rceil.$
Moreover, we set
$\mathfrak{D}_{g_{1},g_{2}}\coloneqq\{\mathscr{D}^{(h_{1},0)}=\langle\mathscr{D}^{(h_{1},0)}_{i}\rangle_{i\in\mathbb{N}},\mathscr{D}^{(h_{2},c_{2})}=\langle\mathscr{D}^{(h_{2},c_{2})}_{i}\rangle_{i\in\mathbb{N}}\}$
in case $c_{2}\neq 0$ and
$\mathfrak{D}_{g_{1},g_{2}}\coloneqq\{\mathscr{D}^{(h_{1},0)}=\langle\mathscr{D}^{(h_{1},0)}_{i}\rangle_{i\in\mathbb{N}}\}$
otherwise. Then the local structure theorem $\mathfrak{S}(\mathfrak{D}_{g_{1},g_{2}},\mathbf{a}_{g_{1},g_{2}},\mathbf{g}_{\mathsf{o},g_{1},g_{2}},\mathbf{g}_{\mathsf{n},g_{1},g_{2}},\mathbf{v}_{\mathsf{b},{g_{1},g_{2}}},\mathbf{v}_{\mathsf{d},{g_{1},g_{2}}},\mathbf{R}_{g_{1},g_{2}})$ is valid.
To prove Theorem 5.1 we start by showing that we can always either find a Dyck-grid of order $k$ which corresponds to one of the two fixed Euler-genus $g_{1}$ and $g_{2}$, or we find a $\Sigma$-configuration whose parameters depend only on $k,$ $g_{1},$ and $g_{2}$ where the Euler-genus of $\Sigma$ is strictly less than $g_{1}$ if $\Sigma$ is orientable, and strictly less than $g_{2}$ if $\Sigma$ is non-orientable and $g_{2}\neq 0$ or at most $g_{1}$ if $\Sigma$ is non-orientable but $g_{2}=0,$
[…]
Let $k,h_{1},h_{2},c_{2}$ be non-negative integers where $c_{2}\in[0,2]$ and $0\leq h_{2}<h_{1}.$ Let $g_{1}\coloneqq 2h_{1},$ $g_{2}\coloneqq 2h_{2}+c_{2}$ if $c_{2}\neq 0$ and $g_{2}\coloneqq 0$ otherwise, and $g\coloneqq\max\{g_{1},g_{2}\}$. Moreover, let $t\geq 12(162^{2g}+4)k$ be an integer.
Finally, let $r,s_{2},M\geq 1$ be integers with $r\geq 16t,$ $s_{2}\geq 2M\geq 2t,$ and set $\mu\coloneqq\left\lfloor\frac{1}{10^{5}t^{26}}t^{10^{7}t^{26}}\cdot M\right\rfloor.$ Fix
$R\coloneqq 49152t^{24}(r+2s_{2})+\mu.$
Let $G$ be a graph, and let $W$ be an $R$-wall in $G.$
Then either $G$ contains $\mathscr{D}^{(h_{1},0)}_{k}$ or, in case $c_{2}\neq 0,$ $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor, or there exist integers $c,h\geq 0$ with $2h<g_{1},$ $2h+c<g_{2}$ if $c\neq 0$ and $c_{2}\neq 0,$ and $2h+c\leq g_{1}$ if $c\neq 0$ and $c_{2}=0,$ a set $A\subseteq V(G)$ of size at most $\mu,$ a surface $\Sigma=\Sigma^{(h,c)}$, and a $\Sigma$-configuration $(W_{1},W_{0},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,c,h)$.
Moreover, the vortex society of $o_{1}$ has a rendition $\rho_{1}$ of breadth at most $2t^{2}$ and depth at most $\mu$ such that
  • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $c,$ and
  • for each vortex cell $c\in C(\rho_{1})$ let $\Delta_{c}\subseteq\Sigma$ be the disk bounded by the trace of $C_{M}\in\mathcal{C}_{c},$ then for each pair of distinct vortex cells $c,c^{\prime}\in C(\rho_{1})$ we have that $\Delta_{c}\cap\Delta_{c^{\prime}}=\emptyset.$
Proof.
We use Subsection 4.3 to either obtain a $K_{t}$-minor or a set $A\subseteq V(G)$ of size at most $\mu$ together with integers $c,h\leq{t+1\choose 2},$ a surface $\Sigma$ of Euler-genus $2h+c,$ and a $\Sigma$-configuration $(W_{1},W_{2},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\gamma)$ of $G-A$ with parameters $(r,M,s_{2},M,M,c,h)$ such that the vortex society of $c_{1}$ in $\delta$ has depth at most $\mu$ and breadth at most $2t^{2}$ together with the desired properties for the nests.
In case we find the $K_{t}$-minor, we contain every graph on $12(162^{2g}+4)k$ vertices as a minor and thus, in particular, both $\mathscr{D}^{(h_{1},0)}_{k}$ and $\mathscr{D}^{(h_{2},c_{2})}_{k}.$ So in this case we are done.
In case $c=2n\neq 0,$ let $h^{\prime}\coloneqq h+n-1$ and $c^{\prime}\coloneqq 2$, similarly, if $c=0$ we set $h^{\prime}\coloneqq h$ and $c^{\prime}\coloneqq c=0$. Moreover, if $c=2n+1,$ let $h^{\prime}\coloneqq h+n$ and $c^{\prime}\coloneqq 1.$ By Subsection 5.1 we find $D\coloneqq\mathscr{D}^{(h^{\prime},c^{\prime})}_{12k}$ as a minor of $G.$
What follows is an analysis on the numbers $h^{\prime},c^{\prime},c,$ and $h$ in relation to $g_{1}$ and $g_{2}.$
Suppose $h^{\prime}\geq h_{1},$ then $D$ contains $\mathscr{D}_{k}^{(h_{1},0)}$ as a minor, and we are done. So we may assume $h^{\prime}<h_{1}$ in the following. This settles the cases where $c=0$ or $c_{2}=0$ since both situations imply that $2h^{\prime}+c^{\prime}\leq 2h<g_{1}.$
From now on we may assume $c,c_{2}\neq 0$ which implies that $c^{\prime},c_{2}\in[2].$
Let us first assume that $c^{\prime}=1$ and $2h^{\prime}+c^{\prime}\geq 2h_{2}+c_{2}.$ If $c_{2}=1$ we immediately obtain $h^{\prime}\geq h_{2}.$ Hence, we may find $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor in $D$ as desired. So we may assume $c_{2}=2.$ In this case, however, since the two sums have different parity we obtain $h^{\prime}>h_{2}.$ We now use Subsection 2.3 to obtain the surface grid $D^{\prime}$ of order $k$ with $h^{\prime}-1$ handles and $c^{\prime}+2$ crosscaps as a minor within $D.$ As $h^{\prime}-1\geq h_{2}$ and certainly $c^{\prime}+2>c_{2}$ we find $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor in $D^{\prime}.$
For the final case we assume $c^{\prime}=2$ and again $2h^{\prime}+c^{\prime}\geq 2h_{2}+c_{2}.$ Since this means $c^{\prime}\geq 2\geq c_{2}\geq 1$ we also obtain $h^{\prime}\geq h_{2}$ which immediately implies the presence of a $\mathscr{D}^{(h_{2},c_{2})}_{k}$-minor in $D$ as desired and our proof is complete. ∎
5.2 Proof of Theorem 5.1
With Subsection 5.1 at hand our proof of Theorem 5.1 becomes an analogue of the proof for Theorem 2.11 from [32]. However, we split the proof in half to make an intermediate structure accessible for future applications. The reason for this is the requirement of a local structure theorem to produce a $\Sigma$-decomposition which is $W$-central with respect to a large wall. The way $W$-centrality is achieved in [32] is by removing additional vertex sets from the vortices which potentially damages the surrounding nests. Some applications might need access to the nests around vortices after the fact, and thus we include an intermediate step here.
The main reason most of the structural theory talks about walls rather than grids is that walls are subcubic graphs and as such, if they are contained as minors in a graph, one can find a subdivision as a subgraph. For a similar reason we define Dyck-walls. Let $c\in[0,2]$, $h$, and $t$ be non-negative integers. The elementary $(h,c;t)$-Dyck-wall is obtained from $\mathcal{D}^{h,c}_{2t}$ by deleting the cycles $C_{t+1},\dots,C_{2t}$ and by deleting the edge $v^{i}_{j}v^{i+1}_{j}$ for every odd $i\in[t-1]$ and every odd $j\in[8t]$ and for every even $i\in[t-1]$ and even $j\in[8t]$. Moreover, for each handle and crosscap that was added to create $\mathcal{D}^{h,c}_{2t}$ delete every edge starting on $v^{1}_{\ell}$ where $\ell$ is even. That is, we delete every second edge. An $(h,c;t)$-Dyck-wall is a subdivision of the elementary $(h,c;t)$-Dyck-wall.
Consider an embedding of a $(h,c;t)$-Dyck-wall $D$ in a surface $\Sigma$ of Euler-genus $2h+c$. Then $C_{t}$ bounds a disk in which no vertex of $D$ is drawn. We call this disk the simple face of $D$. Similarly, there exists a unique face which is incident to $C_{1}$ and every handle and every crosscap of $D$. We call this face the exceptional face of $D$.
Moreover, we will need the following seminal result of Grohe. We say that a graph $G$ is quasi-$4$-connected if for every separation $(A_{1},A_{2})$ of order at most three in $G$ there exists a unique $i\in[2]$ such that $|A_{i}\setminus A_{3-i}|\leq 1$.
[…]
Every graph $G$ has a tree-decomposition $(T,\beta)$ of adhesion at most three such that for all $t\in V(T)$ the torso $G_{t}$ of $G$ at $t$ is a minor of $G$ that is either quasi-$4$-connected or a complete graph of order at most four.
We call a graph $H$ that appears as the torso of a node in a tree-decomposition of $G,$ as in Subsection 5.2, a quasi-$4$-connencted component of $G$.
We now provide the following consequence of Subsection 5.1 that reveals several features of the local structure of a graph where a Dyck-grid is excluded. Subsection 5.2 is of independent interest and may serve as the departure point of future developments.
[…]
Let $h_{1},h_{2},c_{2}$ be non-negative integers where $c_{2}\in[0,2]$ and $0\leq h_{2}<h_{1}.$ We set
$g_{1}$$\coloneqq$ $2h_{1},$
[…]
$g$$\coloneqq$ $\max\{g_{1},g_{2}\}.$
Let also $k$, $r$, q be non-negative integers, where, $r\geq 16\cdot 12(162^{2g}+4)k$ and $q\geq 6k$, and set
$t$$\coloneqq$ $12(162^{2g}+4)k,$
$M$$\coloneqq$ $2(162^{2g}+4)q,$
$\mu$$\coloneqq$ $\left\lfloor\frac{1}{10^{5}t^{26}}t^{10^{7}t^{26}}\cdot M\right\rfloor,\mbox{\leavevmode\nobreak\ and\leavevmode\nobreak\ }$
$R$$\coloneqq$ $49152\cdot t^{24}(r+4M)+\mu.$
If a graph $G$ contains an $R$-wall $W$, then either,
  1. 1.
    $G$ contains $\mathscr{D}^{(h_{1},0)}_{k}$ or, in case $c_{2}\neq 0,$ $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor, or
  2. 2.
    there exist
    • a surface $\Sigma\coloneqq\Sigma^{(h,c)}$ where
      • $2h<g_{1}$,
      • if $c\neq 0$ and $c_{2}\neq 0$ then $2h+c<g_{2}$, and
      • if $c\neq 0$ and $c_{2}=0$ then $2h+c\leq g_{1}$,
    • a set $A\subseteq V(G)$ of size at most $\mu,$ and
    • a $\Sigma$-decomposition $\delta=(\Gamma,\mathcal{D})$ of $G-A$ with the following properties:
      1. (a)
        $\delta$ is of breadth at most $2t^{2}$ and depth at most $\mu$,
      2. (b)
        if we define $c_{0}$ and $h_{0}$ as follows:
        • if $c=0$, let $c_{0}\coloneqq 0$ and $h_{0}\coloneqq h$,
        • if $c>0$ and $c+2h$ is even, let $c_{0}\coloneqq 2$ and $h_{0}\coloneqq\frac{c+2h}{2}-1$, and
        • otherwise, let $c_{0}\coloneqq 1$ and $h_{0}\coloneqq\frac{c+2h-1}{2}$,
        then there exists an $(h_{0},c_{0};q)$-Dyck-wall $D$ in $G-A$ such that $D$ is grounded in $\delta$
      3. (c)
        there exists an $r$-subwall $W^{\prime}$ of $W$ such that $W^{\prime}$ is grounded in $\delta,$ and $W^{\prime}$ is drawn in the interior of the disk defined by the trace of the boundary of the simple face of $D$,
      4. (d)
        for each vortex cell $v\in C(\delta)$ there exists a nest $\mathcal{C}_{v}$ of order $M$ in $\delta$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $v,$
      5. (e)
        for each vortex cell $v\in C(\delta)$ let $\Delta_{v}\subseteq\Sigma$ be the disk bounded by the trace of $C_{M}\in\mathcal{C}_{v},$ then for each pair of distinct vortex cells $v,v^{\prime}\in C(\delta)$ we have that $\Delta_{v}\cap\Delta_{v^{\prime}}=\emptyset,$
      6. (f)
        let $\widetilde{\Delta}$ be the disk in $\Sigma$ which is bounded by the trace of the exceptional face of $D$ and does not contain any vertex of $W^{\prime}$, for every every vortex $v\in C(\delta)$ we have $\Delta_{v}\subseteq\widetilde{\Delta}$,
      7. (g)
        let $X$ be the set of all vertices of $G$ drawn in the interior of cells of $\delta$ and let $G_{\delta}$ be the torso of the set $V(G)\setminus(X\cup A)$ in $G-A$, then $G_{\delta}$ is a quasi-$4$-connected component of $G-A$, and
      8. (h)
        finally if holds that $\mathcal{T}_{W^{\prime}}\subseteq\mathcal{T}_{W}$ and $\mathcal{T}_{D}\subseteq\mathcal{T}_{W}$.
Proof.
By our choice of parameters we may immediatelly call upon Subsection 5.1. The first possible outcome of Subsection 5.1 yields a minor model of $\mathscr{D}^{(h_{1},0)}_{k}$ or, in case $c_{2}\neq 0,$ of $\mathscr{D}^{(h_{2},c_{2})}_{k}.$
The second outcome provides us with integers $c,h\geq 0$ with $2h<g_{1},$ $2h+c<g_{2}$ if $c\neq 0$ and $c_{2}\neq 0,$ and $2h+c\leq g_{1}$ if $c\neq 0$ and $c_{2}=0,$ a set $A\subseteq V(G)$ of size at most $\mu,$ a surface $\Sigma$ of Euler-genus $c+2h,$ which can be obtained from the sphere by adding $h$ handles and $c$ crosscaps, and a $\Sigma$-configuration $(W_{1},W_{0},\delta,o_{0},o_{1},\Delta_{0},\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{P}_{1},\mathcal{P}_{2},\gamma)$ of $G-A$ with parameters $(r,M,2M,M,M,c,h)$.
Additionally, we have that the vortex society of $o_{1}$ has a rendition $\rho_{1}=(\Gamma_{1},\mathcal{D}_{1})$ of breadth at most $2t^{2}$ and depth at most $\mu$ such that
  • for each vortex cell $c\in C(\rho_{1})$ there exists a nest $\mathcal{C}_{c}$ of order $M$ in $\rho_{1}$ around the unique disk $\Delta\in\mathcal{D}$ corresponding to $c,$ and
  • for each vortex cell $c\in C(\rho_{1})$ let $\Delta_{c}\subseteq\Sigma$ be the disk bounded by the trace of $C_{M}\in\mathcal{C}_{c},$ then for each pair of distinct vortex cells $c,c^{\prime}\in C(\rho_{1})$ we have that $\Delta_{c}\cap\Delta_{c^{\prime}}=\emptyset.$
We extend $\delta$ to a $\Sigma$-decomposition $\delta^{\prime}=(\Gamma^{\prime},\mathcal{D})$ of $G-A$ by fixing a vortex free rendition of the vortex society of $o_{0}$ such that the wall $W_{0}$ is flat in the rendition and fixing a rendition of breath at most $2t^{2}$ and depth at most $\mu$ of the vortex society of $o_{1}.$ Thus, $\delta^{\prime}$ has breath at most $2t^{2}$ and depth at most $\mu.$
Next, Subsection 5.1 provides us with a minor-model $D^{\prime}$ of $\mathscr{D}^{h_{0},c_{0}}_{2q}$ which is grounded in $\delta^{\prime}$. Notice that $D^{\prime}$ contains an $(h_{0},c_{0};q)$-Dyck-wall $D$ as a subgraph which still is grounded in $\delta^{\prime}$. In particular, $D^{\prime}$ is also grounded in $\delta$, implying that $D$ is as well. From the construction of $D^{\prime}$ from the mixed surface grid in Subsection 5.1 it follows that $o_{0}$ sits in the disk bounded by the trace of the simple face of $D$ while $o_{1}$ belongs to the disk bounded by the exceptional face of $D$.
We now start to further modify $\delta^{\prime}$ by iterating the following operation exhaustively. Let $(Y_{1},Y_{2})$ be a separation of order at most three in $G-A$ such that $Y_{1}\cap Y_{2}$ consists of nodes of $\delta^{\prime}$ and assume that $(Y_{1},Y_{2})$ does not cross any separation induced by a non-vortex cell of $\delta^{\prime}$. That is, we assume that there does not exist a non-vertex cell $x\in C(\delta^{\prime})$ such that $V(\sigma^{\prime}(x))\cap(Y_{1}\setminus Y_{2})\neq\emptyset$ and $V(\sigma^{\prime}(x))\cap(Y_{2}\setminus Y_{1})\neq\emptyset$. By the submodularity of separations we may always find such a separation $(Y_{1},Y_{2})$ if some separation of order at most three with only nodes in its separator exists. We start with separations of order one, then consider separations of order two, and finally separations of order three. Without loss of generality we may assume that $Y_{1}$ contains a row and a column of $W_{0}$. Moreover, we assume that $Y_{2}$ is maximal with respect to set containment. This implies that $Y_{1}\setminus Y_{2}$ is connected. Indeed, it follows from the definition of $\Sigma$-configurations and the $3$-connectivity of Dyck-grids that also at most one branch vertex of $D$ may be contained in $Y_{2}\setminus Y_{1}$. Now suppose $Y_{2}\setminus Y_{1}$ contains a node of $\delta^{\prime}$. Let $C_{(Y_{1},Y_{2})}$ be the collection of all cells $x$ of $\delta^{\prime}$ such that $V(\sigma^{\prime}(x))\cap(Y_{2}\setminus Y_{1})\neq\emptyset$. It follows that $V(\sigma^{\prime}(x))\subseteq Y_{2}$ for all $x\in C_{(Y_{1},Y_{2})}$ by our choice of $(Y_{1},Y_{2})$. We may now define a new cell $c_{(Y_{1},Y_{2})}$ replacing all cells in $C_{(Y_{1},Y_{2})}$ as follows. Let $\Delta_{(Y_{1},Y_{2})}$ be the disk which is the closure of $\bigcup_{c\in C_{(Y_{1},Y_{2})}}c$. Now let $c_{(Y_{1},Y_{2})}$ be the cell obtained by removing the points $\pi^{\prime-1}(y)$, $y\in Y_{1}\cap Y_{2}$, from $\Delta_{(Y_{1},Y_{2})}$, and notice that all vertices of $Y_{2}\setminus Y_{1}$ are drawn into the interior of $c_{(Y_{1},Y_{2})}$. Then forget all cells in $C_{(Y_{1},Y_{2})}$. Let $\delta^{\prime\prime}=(\Gamma^{\prime\prime},\mathcal{D}^{\prime\prime})$ be the $\Sigma$-decomposition obtained once the above operation cannot be applied any more. Notice that $D$ is still grounded in $\delta^{\prime\prime}$ and so is $W_{0}$. The same is true for all cycles from the nests of the vortices of $\delta^{\prime\prime}$. This implies that the torso $G_{\delta^{\prime\prime}}$ contains a Dyck-wall $D^{\prime\prime}$ of the same order as $D$ which can be obtained by contracting some edges of $D$ and is quasi-$4$-connected.
If follows that the set $S$ of vertices of $G-A$ corresponding to the nodes of $\delta^{\prime\prime}$ cannot be split further by separations of order at most three and $S$ is minimal with this property. Hence, by applying Subsection 5.2 there must exist a bag of the resulting tree-decomposition whose vertex set is exactly $S$.
So far we can observe that $\delta^{\prime\prime}$ satisfies condition a) since it is of breadth at most $2t^{2}$ and depth at most $\mu$. It also holds that $D$ is grounded in $\delta^{\prime\prime}$ and none of its vertices are drawn in $\Delta_{c}$ for any vortex of $\delta^{\prime\prime}$, hence property b) is satisfied. The wall $W_{0}$ is an $r$-subwall of $W$ which is drawn in the disk bounded by the trace of the simple face of $D$ and rooted in $\delta^{\prime\prime}$ which implies property c). Moreover, every vortex of $\delta^{\prime\prime}$ is contained in the disk bounded by the exceptional face of $D$ as required by d) and all vortices are equipped with nests of order $M$ which is property e).
These nests are pairwise disjoint and belong to their private disks, hence f) is satisfied. Finally, the construction above ensures that $G_{\delta^{\prime\prime}}$ is quasi-$4$-connected and so we also get property g).
So we are left with discussing the tangles. For this observe that, by the definition of $\Sigma$-configurations, $W_{0}$ is a subwall of $W_{1}$, $W_{1}$ is a subwall of $W$, and $(\mathcal{P}_{1}\cup\mathcal{P}_{2},\mathcal{C}_{2})$ overlays $W_{1}$. Recall that $D$ consists entirely of cycles from $\mathcal{C}_{2}$ and the paths from $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$. Since $r,q<M$ it follows that the $W$-majority side of any separation of order less than $r$ must also be the $W_{0}$-majority side and the $W$-majority side of any separation of order less than $q$ must also be the $D$-majority side. Hence, also property h) holds as desired and our proof is complete. ∎
An advantage of Subsection 5.2 over Theorem 5.1 is that here the parameter $M$ is freely choosable, as long as it is at least $t$. This means that one may increase the order of the Dyck-wall and the nests around the vortices arbitrarily and independent of the excluded minors. We continue with the final step in the proof of Theorem 5.1.
The proof of Subsection 5.2 can be turned into an algorithm that given a graph $G$, $k,h_{1},h_{2},c_{2},r,M$ and an $R$-wall as above, computes the structures implied by Subsection 5.2 in time
$\mathcal{O}(M^{2}\cdot 2^{\mathsf{poly}(k\cdot 2^{\max\{2h_{1},2h_{2}+c_{2}\}})}|V(G)|^{2}+|V(G)|^{3}).$
This running time comes from two different sources. The quadratic, in $|V(G)|$, part is due to us calling Subsection 5.1 which in turn calls Subsection 5.1 from [32]. Here the only additional work that needs to be done is the extraction of the Dyck-wall as describe din the proof of Subsection 5.1. Since this is purely handling a bounded size object, the main contribution stems indeed from Subsection 5.1. The second term, cubic in $|V(H)|$, comes from the refinement of our $\Sigma$-decomposition. In particular from the requirement that $G_{\delta^{\prime\prime}}$ is quasi-$4$-connected in the end which makes it necessary to consider all possible separations of order at most three.
[…]
Let $d\geq 0$ be an integer. Every society of depth at most $d$ has a linear decomposition of adhesion at most $d.$
Proof of Theorem 5.1.
Let us start by discussing the numbers. We set
$M$ $\coloneqq t,$
$s_{2}$ $\coloneqq 2t,$
$r^{\prime}$ $\coloneqq r+16t+\left\lfloor\frac{1}{10^{5}t^{24}}t^{10^{7}t^{26}}\right\rfloor\text{, and}$
$\mu$ $\coloneqq\left\lfloor\frac{1}{10^{5}t^{25}}t^{10^{7}t^{26}}\right\rfloor.$
Let $R\coloneqq 49152t^{24}(r^{\prime}+2s_{2})+\mu$ be the constant from Subsection 5.1 and notice that
$R=49152t^{24}(r^{\prime}+2s_{2})+\mu$ $=49152t^{24}(r^{\prime}+4t)+\mu$
$\leq 49152t^{24}\left(r+16t+\left\lfloor\frac{1}{10^{5}t^{24}}t^{10^{7}t^{26}}\right\rfloor+4t\right)+\left\lfloor\frac{1}{10^{5}t^{25}}t^{10^{7}t^{26}}\right\rfloor$
$\leq 49152t^{24}(r+16t)+\frac{49152t^{24}}{10^{5}t^{24}}t^{10^{7}t^{26}}+2\cdot 10^{5}t^{25}+\left\lfloor\frac{1}{10^{5}t^{25}}t^{10^{7}t^{26}}\right\rfloor$
$\leq 49152t^{24}(r+16t)+\frac{1}{2}t^{10^{7}t^{26}}+2\cdot\frac{10^{5}t^{25}}{10^{5}t^{25}}t^{10^{7}t^{26}}$
$\leq 49152t^{24}(r+16t)+\left\lceil\frac{5}{2}t^{10^{7}t^{26}}\right\rceil=\mathbf{R}_{g_{1},g_{2}}.$
Moreover, we have ensured that $r^{\prime}\geq 16t.$ Let $W$ be an $\mathbf{R}_{g_{1},g_{2}}$-wall in a graph $G.$
We may call upon Subsection 5.2 with the numbers $k,h_{1},h_{2},c_{2},$ $r^{\prime},s_{2},M$ and $\mu$ and the wall $W.$ In case we find $\mathscr{D}^{(h_{1},0)}_{k}$ or $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as a minor we are done, so we may assume the final outcome of Subsection 5.2 to hold.
This means we find a set $A\subseteq V(G)$ of size at most $\mu$ and $\Sigma$-decomposition $\delta=(\Gamma,\mathcal{D})$ of $G-A$ satisfying properties a) to h) from Subsection 5.2. What remains is to ensure that $\delta$ is $W$-central.
Let $y_{1},\dots,y_{n}\in C(\delta)$ be the vortices of the $\Sigma$-decomposition $\delta.$ Let $P_{1},\dots,P_{\mathbf{R}_{g_{1},g_{2}}}$ be the horizontal paths and $Q_{1},\dots,Q_{\mathbf{R}_{g_{1},g_{2}}}$ be the vertical paths of $W.$ Moreover, let
$\mathcal{W}\coloneqq\{P_{i}\cup Q_{j}\mid i,j\in[\mathbf{R}_{g_{1},g_{2}}]\text{ where }(V(P_{i})\cup V(Q_{j}))\cap A=\emptyset\}.$
Observe that for every $T\in\mathcal{W}$ there exist at least $r^{\prime}-\mu$ vertex disjoint paths in $W-A$ that link $T$ and $W_{0}.$ Thus, the only cells $y\in C(\delta^{\prime})$ such that an element of $\mathcal{W}$ is a subgraph of $\Sigma_{\delta^{\prime}}(y)$ are the vortices $y_{i},$ $i\in[n].$
Claim: Let $i\in[n]$ and let $(G_{i},\Omega_{i})$ be the vortex society of $y_{i}.$ There exists a set $A_{i}\subseteq V(G_{i})$ with $|A_{i}|\leq 2\mu+1$ such that $G_{i}-A_{i}$ does not contain any element of $\mathcal{W}$ as a subgraph.
By Subsection 5.2 there exists a linear decomposition $\langle X_{1},\dots,X_{m},v_{1},\dots,v_{m}\rangle$ of $(G_{i},\Omega_{i})$ which has adhesion at most $\mu.$ For every $T\in\mathcal{W}$ with $T\subseteq G_{i}$ let $a_{T}\in[m]$ be the smallest integer such that $V(T)\cap X_{a_{T}}\neq\emptyset$ and let $b_{T}\in[m]$ be the largest integer with $V(T)\cap X_{b_{T}}\neq\emptyset$ Notice that the fact that each $T\in\mathcal{W}$ with $T\subseteq G_{i}$ is connected in $G_{i}$ implies that $V(T)\cap X_{j}\neq\emptyset$ for all $j\in[a_{T},b_{T}].$ Let $b$ be the smallest value in the set $\{b_{T}\mid T\in\mathcal{W}\text{ and }T\subseteq G_{i}\}.$ It follows from the fact that $T\cup T^{\prime}$ is connected for all $T^{\prime},T\in\mathcal{W}$ that $V(T)\cap X_{b}\neq\emptyset$ for every $T\in\mathcal{W}$ with $T\subseteq G_{i}.$
Let $A_{i}\coloneqq(X_{b}\cap X_{b-1})\cup(X_{b}\cap X_{b+1})\cup\{v_{b}\}.$ Since the linear decomposition is of adhesion at most $\mu$ we obtain that $|A_{i}|\leq 2\mu+1.$
Moreover, $G_{i}-A_{i}$ does not contain any member of $\mathcal{W}$ as a subgraph as desired. To see this let us assume, towards a contradiction, that there exists some $T\in\mathcal{W}$ with $T\subseteq G_{i}$ such that $T\subseteq G_{i}-A_{i}.$ Since $A_{i}$ separates the vertices of $X_{b}\setminus A_{i}$ from the rest of $G$ and $T$ must intersect $X_{b},$ the only way how $T$ could avoid $A_{i}$ is if $V(T)\subseteq X_{b}\setminus A_{i}.$ Recall that there exist $r^{\prime}-\mu$ pairwise vertex disjoint paths in $W-A$ that that link $T$ to $W_{0}$ where $V(W_{0})\cap V(G_{i})=\emptyset.$ Furthermore, we have
$r^{\prime}-\mu$ $=r+16t+\left\lfloor\frac{1}{10^{5}t^{24}}t^{10^{7}t^{26}}\right\rfloor-\left\lfloor\frac{1}{10^{5}t^{25}}t^{10^{7}t^{26}}\right\rfloor$
$\geq\frac{t}{10^{5}t^{25}}t^{10^{7}t^{26}}-\frac{1}{10^{5}t^{25}}t^{10^{7}t^{26}}-1$
$=\frac{t-1}{10^{5}t^{25}}t^{10^{7}t^{26}}-1$
$>3\mu>|A_{i}|.$
So $A_{i}$ cannot separate $T$ from $W_{0}$, and thus we have reached a contradiction and the proof of the claim is complete.
Set $\hat{A}\coloneqq A\cup\bigcup_{i\in[n]}A_{i}$ and observe that
$|\hat{A}|$ $\leq 2t^{2}(2\mu+1)+\mu$
$\leq\frac{5t^{2}}{10^{7}t^{25}}t^{10^{7}t^{26}}$
$\leq t^{10^{7}t^{26}}.$
Moreover, from the construction of $\hat{A}$ we obtain that $\hat{A}\setminus A\subseteq\bigcup_{i\in n}V(\Sigma_{\delta^{\prime}}(y_{i})),$ that is, every vertex of $\hat{A}$ that does not belong to $A$ belongs to some vortex of $\delta.$ In particular, this means that $\hat{A}$ is disjoint from $W_{0}.$ Finally, let $\delta^{\prime}$ be the restriction of $\delta$ to the graph $G-\hat{A}.$ Now, $\delta^{\prime}$ is a $\Sigma$-decomposition of $G-\hat{A}$ which is $W$-central, has breadth at most $2t^{2}$ and depth at most $\mu\leq t^{10^{7}t^{26}}.$ Moreover, $W_{0}\subseteq W-\hat{A}$ contains an $r$-wall which is flat in $\delta^{\prime}$ and thus, our proof is complete. ∎
Theorem 5.2.
Let $\mathbb{S}$ be a non-empty finite and closed set of surfaces. Let $g_{1}$ be the Euler-genus of the orientable surface in $\mathsf{sobs}(\mathbb{S})$ and $g_{2}$ be the Euler-genus of the non-orientable surface in $\mathsf{sobs}(\mathbb{S})$ if such a surface exists, otherwise set $g_{2}\coloneqq 0.$
There exists a function $f\colon\mathbb{N}\to\mathbb{N}$ such that for every $k\in\mathbb{N},$ every graph $G$ which excludes the orientable Dyck-grid of order $k$ and Euler-genus $g_{1}$ and, if $g_{2}\neq 0,$ the non-orientable Dyck-grid of order $k$ and Euler-genus $g_{2}$ satisfies ${\mathbb{S}}\mbox{-}\mathsf{hw}(G)\leq f(k).$
In particular, $G$ has a tree-decomposition $(T,\beta)$ of adhesion less than $f(k)$ such that for every $d\in V(T),$ either $|\beta(d)|\leq f(k),$ or there exist a set $A_{d}\subseteq\beta(d)$ of size at most $f(k)$ and a surface $\Sigma\in\mathbb{S}$ such that $\mathsf{Torso}_{G}(d)-A_{d}$ has a $\Sigma$-decomposition $\delta$ of width at most $f(k)$ and breadth at most $f(k).$
[…]
  • i.
    for every neighbor $d^{\prime}$ of $d$ in $T$ we have that $|(\beta(d)\cap\beta(d^{\prime}))\setminus X_{d}|\leq 3$ and
  • ii.
    $\mathsf{Torso}_{G}(d)-X_{d}$ is a minor of $G.$
Proof.
Let us first assume that $\mathbb{S}=\{\Sigma^{\varnothing}\}.$ In this case $\mathsf{sobs}(\mathbb{S})$ contains only surfaces homeomorphic to the sphere. Thus, the orientable Dyck-grid and the non-orientable one are both of Euler-genus zero. Therefore, this case follows immediately from Subsection 3.1.
From now on we may assume $\mathbb{S}$ to contain a non-void surface.
Suppose $g_{2}=0.$ This means that $\mathsf{sobs}(\mathbb{S})$ contains only orientable surfaces of Euler-genus $g_{1}.$ Notice that this implies that $\mathbb{S}$ contains a non-orientable surface of Euler-genus $g_{1}$ and $g_{1}+1$.
Let $g_{1}=2h_{1}.$ Now let $k\in\mathbb{N}$ be some integer and $G$ be some graph that excludes $\mathscr{D}^{(h_{1},0)}_{k}$ as a minor. In case $k\in[0,1]$ this graph is planar, and thus our claim follows from Subsection 3.1. Thus, we may assume $k\geq 2.$ Finally, let $t\coloneqq 12(162^{2g_{1}}+4)k.$ Let
$\mathbf{a}_{g_{1},0}$ $\coloneqq t^{10^{7}t^{26}}$
$\mathbf{g}_{\mathsf{o},g_{1},0}$ $\coloneqq g_{1}-1$
$\mathbf{g}_{\mathsf{n},g_{1},0}$ $\coloneqq{g_{1}}$
$\mathbf{v}_{\mathsf{b},{g_{1},0}}$ $\coloneqq 2t^{2}$
$\mathbf{v}_{\mathsf{d},{g_{1},0}}$ $\coloneqq t^{10^{7}t^{26}}$
$\mathbf{R}_{g_{1},0}$ $\coloneqq 49152t^{24}(r+16t)+\left\lceil\frac{5}{2}t^{10^{7}t^{26}}\right\rceil.$
Moreover, let $\mathfrak{D}\coloneqq\{\mathscr{D}^{(h_{1},0)}=\langle\mathscr{D}^{(h_{1},0)}_{i}\rangle_{i\in\mathbb{N}},\}.$ Then $\mathfrak{S}(\mathfrak{D}_{g_{1},0},\mathbf{a}_{g_{1},0},\mathbf{g}_{\mathsf{o},g_{1},0},\mathbf{g}_{\mathsf{n},g_{1},0},\mathbf{v}_{\mathsf{b},{g_{1},0}},\mathbf{v}_{\mathsf{d},{g_{1},0}},\mathbf{R}_{g_{1},0})$ is valid by Theorem 5.1.
Our claim now follows immediately from Theorem 4.2.
Hence, we may now assume $\mathsf{sobs}(\mathbb{S})$ to contain a non-void surface which is non-orientable and one that is orientable. Thus, there exist integers $g_{1}$ and $g_{2}$ where $g_{2}\neq 0$ such that every orientable surface in $\mathsf{sobs}(\mathbb{S})$ has Euler-genus $g_{1}$ and every non-orientable surface in $\mathsf{sobs}(\mathbb{S})$ has Euler-genus $g_{2}.$ In particular, we may assume for every orientable $\Sigma\in\mathsf{sobs}(\mathbb{S})$ and every non-orientable $\Sigma^{\prime}\in\mathsf{sobs}(\mathbb{S})$ that $\Sigma\not\preceq\Sigma^{\prime}$ and $\Sigma^{\prime}\not\preceq\Sigma.$
Let $g_{1}=2h_{1}$ and $g_{2}=2h_{2}+c_{2}$ where $c_{2}\in[2].$ From the discussion above it follows that $h_{2}<h_{1}.$ Now let $k\in\mathbb{N}$ be some integer and $G$ be some graph that excludes both $\mathscr{D}^{(h_{1},0)}_{k}$ and $\mathscr{D}^{(h_{2},c_{2})}_{k}$ as minors. In case $k\in[0,1]$ both of these graphs are planar, and thus our claim follows from Subsection 3.1. Thus, we may assume $k\geq 2.$ Finally, let $g\coloneqq\max\{g_{1},g_{2}\}$ and $t\coloneqq 12(162^{2g}+4)k.$ Let $\mathbf{a}_{g_{1},g_{2}}\coloneqq\mathbf{a}_{g_{1},0}$, $\mathbf{g}_{\mathsf{o},g_{1},g_{2}}\coloneqq\mathbf{g}_{\mathsf{o},g_{1},0}$, $\mathbf{v}_{\mathsf{b},{g_{1},g_{2}}}\coloneqq\mathbf{v}_{\mathsf{b},{g_{1},0}}$, $\mathbf{v}_{\mathsf{d},{g_{1},g_{2}}}\coloneqq\mathbf{v}_{\mathsf{d},{g_{1},0}}$, $\mathbf{R}_{g_{1},g_{2}}\coloneqq\mathbf{R}_{g_{1},0}$, and
$\mathbf{g}_{\mathsf{n},g_{1},g_{2}}\coloneqq{g_{2}-1}$
Moreover, let $\mathfrak{D}\coloneqq\{\mathscr{D}^{(h_{1},0)}=\langle\mathscr{D}^{(h_{1},0)}_{i}\rangle_{i\in\mathbb{N}},\mathscr{D}^{(h_{2},c_{2})}=\langle\mathscr{D}^{(h_{2},c_{2})}_{i}\rangle_{i\in\mathbb{N}}\}.$ Then, by Theorem 5.1, the local structure theorem $\mathfrak{S}_{(g_{1},g_{2})}(\mathfrak{D}_{g_{1},g_{2}},\mathbf{a}_{g_{1},g_{2}},\mathbf{g}_{\mathsf{o},g_{1},g_{2}},\mathbf{g}_{\mathsf{n},g_{1},g_{2}},\mathbf{v}_{\mathsf{b},{g_{1},g_{2}}},\mathbf{v}_{\mathsf{d},{g_{1},g_{2}}},\mathbf{R}_{g_{1},g_{2}})$ is valid.
Our claim now follows immediately from Theorem 4.2. ∎
6 Bounding annotated treewidth
Recall that Theorem 5.2 provides a tree-decomposition $(T,\beta)$ where for each $t\in V(T)$ there exists a set $A\subseteq\beta(t)$ such that the torso $G_{t}$ of $G$ at $t$ has a $\Sigma$-decomposition $d$ of width at most $f(k)$ and breadth at most $f(k)$ after deleting $A$. In the statement of the same lemma we also considered the set $X_{t}$ containing all vertices of $G_{t}-A_{t}$ which are drawn in some vortex of $\delta$ together with the vertices of $A_{t}.$ The purpose of this section is to prove that $X_{t}$ is a low-treewidth modulator of the torso $G_{d}$ whose annotated treewidth is bounded by a function the width and the breath of $\delta$ as well as the Euler-genus of $\Sigma.$ This will permit us to substitute the use of the apex vertices $A_{d}$ and the vortices of $\delta$ by a single low-treewidth modulator.
In this section we present several results on the biggest grid parameter ${\sf bg}$ and its annotated version. Some of these results will also be useful for the proofs of the lower bounds in Section 7. We start with the formal definitions of annotated extension and ${\sf bg}.$
Annotated extensions of parameters.
We use $\mathcal{A}_{\text{all}}$ in order to denote the set of all annotated graphs. Given a minor-monotone graph parameter ${\sf p}:\mathcal{G}_{\rm all}\to\mathbb{N},$ we define its annotated extension ${\sf p}:\mathcal{A}_{\rm all}\to\mathbb{N}$ so that
${\sf p}(G,X)\coloneqq\max\{{\sf p}(H)\mid\mbox{$H$ is an $X$-minor of $G$}\}.$
Notice that $0={\sf p}(G,\emptyset)\leq{\sf p}(G,X)\leq{\sf p}(G,V(G))={\sf p}(G).$
In this paper we apply the above definition for ${\sf tw}$ and ${\sf bg}$ as defined below.
The ${\sf bg}$ parameter.
The biggest grid parameter ${\sf bg}$ outputs the maximum $k$ for which $G$ contains a $(k\times k)$-grid as a minor. Recall that $\{\Sigma^{0,0}\}\mbox{-}\mathsf{bg}$ is the maximum $k$ for which $G$ contains the annulus grid $\mathscr{A}_{k}$ as a minor, therefore ${\sf bg}\sim_{\mathsf{L}}\{\Sigma^{0,0}\}\mbox{-}\mathsf{bg}.$
Notice that the treewidth and the biggest grid parameters are asymptotically equivalent, in fact it holds that ${\sf tw}\sim_{{\mathsf{P}}}{\sf bg}$ (by Subsection 3.1 and the asymptotic equivalence between grids and walls). In this section we study the annotated extension of ${\sf bg}$: for every annotated graph $(G,X),$ ${\sf bg}(G,X)$ is the biggest $k$ such that $G$ contains a $(k\times k)$-grid as an $X$-minor.
Notice that ${\sf bg}(G,X)\geq k$ if and only if there is a collection $\mathcal{S}=\{S_{1},\ldots,S_{k^{2}}\}$ of connected and pairwise-disjoint vertex sets of $G,$ each containing at least one vertex in $X$ and such that the contraction, in $G[\boldsymbol{\boldsymbol{\bigcup}}\mathcal{S}],$ of all edges with both endpoints in the same $S_{i}$ results in a graph containing a $(k\times k)$-grid as a spanning subgraph.
6.1 Walls in surfaces and their radii
An annotated $r$-wall is a pair $(W,S)$ where
  • $W$ is an $r$-wall,
  • every path of $W$ between two vertices of degree three in $W$ contains exactly one vertex from $S.$
  • all vertices of $S$ have degree two in $W.$
A hairy annotated $r$-wall is a pair $(\hat{W},X)$ such that $\hat{W}=W\cup\boldsymbol{\boldsymbol{\bigcup}}\mathcal{P}$ and $X\subseteq V(\hat{W})$ where
  • $W$ is an $r$-wall that contains a vertex set $S\subseteq V(W)$ where $|S|=|X|$ and such that $(W,S)$ is an annotated $r$-wall,
  • $\mathcal{P}$ is a collection of $|X|$ vertex-disjoint paths each with one endpoint in $S$ and with the other in $X$ and such that $V(W)\cap\bigcup\mathcal{P}=S.$
The branch vertices of $\hat{W}$ are the branch vertices of the $r$-wall $W.$
The following observation follows easily by the fact that every $r$-wall is a subdivision of a subgraph of a $(2k\times 2k)$-grid.
[…]
There is a universal constant $c$ such that for every graph $G$ and every set $X\subseteq V(G),$ if ${\sf bg}(G,X)\geq c{k},$ then $G$ contains a hairy annotated $k$-wall $(W,X)$ as a subgraph.
Figure 13: A hairy annotated $5$-wall $(\hat{W},X).$ The edges and the internal nodes of the paths in $\mathcal{P}$ are orangle. The set $S$ is depicted by the magenta vertices and the set $X$ is depicted by the blue circles. The branch vertices are the square vertices.
We say that a vertex of a graph is a branch vertex if it has degree at least 3. Given two vertices $x,y,$ we define their branch distance, denoted by $\mathsf{bdist}_{G}(x,y)$ as the minimum number of branch vertices of a path between $x$ and $y.$ The branch radius of $G$ is at most $r$ if there is an $x\in V(G)$ such that for every $y\in V(G),$ $\mathsf{bdist}_{G}(x,y)\leq r.$
[…]
Let $(\hat{W},X)$ be a hairy annotated $r$-wall and let $Q$ be the graph obtained by adding a $v_{\text{new}}$ and making it adjacent to all vertices of $X.$ Then $Q$ has branch radius at most two.
Let $\Sigma$ be a surface without boundary and let $G$ be a graph drawn in $\Sigma$ without crossings. Towards simplifying notation, we do not distinguish the graph and its particular embedding, and we use $F(G)$ to denote the set of faces of $G,$ i.e., the set of connected components of the set $\Sigma\setminus G.$ Given $x,y\in V(G)\cup F(G),$ a radial $(x,y)$-path in $G$ of length at most $r$ is a sequence of vertices and faces of length $r+1,$ where faces and vertices alternate, which starts from $x$ and finishes in $y,$ and such that if $f\in F(G)$ and $v\in V(G)$ are consecutive in the sequence, then $v$ is a point of the boundary of $f.$ The radial distance between $x$ and $y,$ denoted by $\mathsf{rdist}_{G,\Sigma}(x,y)$ is the minimum length of a radial $(x,y)$-path in the embedding of $G$ in $\Sigma.$ The radial radius of $G$ is at most $r$ if there is an $x\in V(G)\cup F(G)$ such that for every $y\in V(G)\cup F(G),$ $\mathsf{rdist}_{G,\Sigma}(x,y)\leq r.$
The following has been proven in different forms, see eg, [10, 18, 24].
[…]
Let $G$ be a graph drawn in a surface $\Sigma$ of Euler-genus $g.$ There is a universal constant $c$ such that if the treewidth of G is more than $cr(g+1),$ then $G$ contains an $r$-wall $W$ as a subgraph which is a subset of a contractible closed disk of $\Sigma$ whose boundary is the perimeter of $W.$
[…]
Let $G$ be a graph drawn in a surface $\Sigma$ of Euler-genus at most $k$ with radial radius at most $r.$ Then ${\sf tw}(G)\in\mathcal{O}(g\cdot r)$
Proof.
Let $q$ be the maximum integer such that $G$ contains a $2(q+1)$-wall $W$ as a subgraph that is a subset of a closed disk of $\Sigma$ whose boundary is the perimeter of $W.$ Notice that the radial radius of $W$ is more than $q,$ which in turn implies that also the radial radius of $G$ is at least $q.$ As $G$ has radial radius at most $r,$ we have that $q\leq r.$ This, combined with Subsection 6.1 implies that ${\sf tw}(G)\in\mathcal{O}(g\cdot r).$
6.2 Bounding the biggest grid
Let us denote the $(k\times k)$-grid by $\Gamma_{k}.$ The perimeter of $\Gamma_{k},k\geq 3,$ is defined in the same way as it is defined for walls. We say that the $1$st layer of $\Gamma_{k}$ is its perimeter. Given some $s\leq\lfloor k/2\rfloor$ we define the $s$-th layer of $\Gamma_{k}$ as the perimeter of the $(k-2s+2\times k-2s+2)$-grid remaining if we remove the first $s-1$ layers of $\Gamma_{k}.$ We refer to the first $s$ layers of $\Gamma_{k}$ as its $s$ outermost layers.
Given an $(m\times m)$-grid $\Gamma_{m},m\geq 3,$ and some $k$ where $3\leq k\leq m$ and such that $m-k$ is even, we define the central $(k,k)$-subgrid of $\Gamma_{m}$ as the unique subgraph of $\Gamma_{m}$ obtained by removing its ${\frac{m-k}{2}}$ outermost layers.
Let $G$ and $H$ be graphs, $X\subseteq V(G)$, and assume that $H$ is an $X$-minor of $G,$ certified by the collection $\mathcal{S}=\{S_{v}\mid v\in V(H)\}$ of pairwise-disjoint connected subsets of $V(G).$ If $x\in S_{v},$ then we say that $x$ is a precursor of $v$ in $G$ and that $v$ is the heir of $x$ in $H.$
[…]
Let $(H,Y)$ and $(G,X)$ be annotated graphs where $H$ is an $X$-minor of $G$. Then ${\sf bg}(G,X)\geq{\sf bg}(H,Y).$
An important tool for the study of ${\sf bg}$ is the following.
[…]
Let $\Gamma_{m}$ be the $(m\times m)$-grid and let $S$ be a subset of the vertices in the central $(m-2\ell)\times(m-2\ell)$-subgrid of $\Gamma_{m},$ where =|$\ell=\lfloor\sqrt[4]{|S|}\rfloor.$ $G$ contains the $(\ell\times\ell)$-grid as an $S$-minor.
[…]
Let $G$ be a graph and let $X_{1},\ldots,X_{r}\subseteq V(G)$ such that, for $i\in[r],$ ${\sf bg}(G,X_{i})\leq\ell.$ Then ${\sf bg}(G,X_{1}\cup\cdots\cup X_{r})\in\mathcal{O}(r^{4}\ell^{2}).$
Proof.
Let $X=X_{1}\cup\cdots\cup X_{r}$ and let $q={\sf bg}(G,X).$ Let also $\Gamma_{q}$ be a $(q\times q)$-grid $X$-minor in $G,$ certified by a collection $\mathcal{S}=\{S_{v}\mid v\in V(\Gamma_{q})\}$ of pairwise-disjoint connected subsets of $V(G).$ We choose $i$ so that the cardinality of the set $Z_{i}\coloneqq\{u\in V(\Gamma_{q})\mid X_{i}\cap S_{u}\neq\emptyset\}$ is maximized, i.e., $X_{i}$ has the majority of intersections with the sets of $\mathcal{S}.$ As we fix $i,$ for simplicity we use $Z=Z_{i}.$
Clearly $Z$ is a subset of $V(\Gamma_{q})$ where $|Z|\geq q^{2}/r.$ For each (positive) margin $b,$ we define the set $B_{b}$ as the union of the $b$ outermost layers of $\Gamma_{q}.$ We now fix some $b\in[\lfloor q/2\rfloor],$ and observe that at most $4b(q-b)$ of the vertices of $\Gamma_{q}$ are in the outermost $b$ layers. This means that at least $f(q,b,r)\coloneqq\frac{q^{2}}{r}-4b(q-b)$ vertices of $Ζ$ are in the central $(q-2b)\times(q-2b)$-subgrid of $\Gamma_{m}.$
We choose the margin $b_{q,r}$ as the maximum integer $b$ such that $4]{f(q,b,r)}\rfloor\geq 2b.$
This implies that there are at least $f(q,b_{q,r},r)$ vertices from $Z$ in the $(q-2b_{q,r})\times(q-2b_{q,r})$-subgrid of $\Gamma_{q}.$ By Subsection 6.2 we obtain, ${\sf bg}(G,X_{i})\geq{\sf bg}(\Gamma_{q},Z_{i})\geq\lfloor\sqrt[4]{f(q,b_{q,r},r)}\rfloor\geq 2b_{q,r}\in\Omega(\frac{q^{1/2}}{{r}^{1/4}}).$ Therefore, ${\sf bg}(G,X_{1}\cup\cdots\cup X_{r})\in\mathcal{O}(r^{4}\ell^{2}).$
The next observation follows directly from the definition of ${\sf bg}.$
[…]
Let $G$ be a graph, $A\subseteq V(G),$ and $X\subseteq V(G)\setminus A,$ where ${\sf bg}(G-A,X)\leq k.$ Then ${\sf bg}(G,X\cup A)\leq k+|A|.$
[…]
For every graph $G,$ every $k\in\mathbb{N},$ and every surface $\Sigma$ with Euler-genus at most $g,$ if $A\subseteq V(G),$ $\delta$ is a $\Sigma$-decomposition of $G-A$ of width at most $w$ and breadth at most $b$, and ${\displaystyle X\coloneqq A\cup\bigcup\{\Sigma(c)\mid\mbox{\small$c$ is a vortex of $\delta$}\}},$ then ${\sf bg}(G,X)\in\mathcal{O}(b^{4}\cdot(b\cdot g\cdot w)^{2}+|A|).$
Proof.
Let $c_{1},\ldots,c_{b}$ be the vortices of $\delta.$ Also, let ${(\Sigma(c_{i}),\Omega_{i})}$ be the vortex society of $c_{i},$ $i\in[b].$ Because of Subsection 6.2, we may assume that $A=\emptyset$ and prove that ${\sf bg}(G,X)\in\mathcal{O}(b^{4}\cdot(b\cdot g\cdot w)^{2}).$ Because of Subsection 6.2, it is enough to prove that for every $\hat{c}\in\{c_{1},\ldots,c_{b}\},$ if $S=V(\Sigma(\hat{c})),$ then ${\sf bg}(G,S)\in\mathcal{O}(bgw).$
Let $z=\lfloor{\sf bg}(G,S)/{c^{\prime}}\rfloor$ where $c^{\prime}$ is the constant of Subsection 6.1.
According to Subsection 6.1 $G$ contains a hairy annotated $z$-wall $(\hat{W},S).$ Moreover, ${\sf tw}(\hat{W})\geq z.$
As a first step, we enhance $G$ by adding, for each vortex $c_{i}\in\{c_{1},\ldots,c_{b}\},$ a cycle $C_{i}$ joining every pair of consecutive vertices in $\Omega_{i}.$ We further enhance $G$ by adding a new vertex $a,$ called apex vertex, and make it adjacent to all the vertices of $S.$ We also update $\delta$ by setting $\Sigma(c_{i})\coloneqq\Sigma(c_{i})\cup C_{i},$ for every $i\in[b],$ and we further enhance $\Sigma(\hat{c})$ by adding $a$ and making it adjacent to all vertices of $S.$
Notice that the above enhancements cannot increase the width of the vortices of $\delta$ by more than three. We now define the pair $(W^{+},S)$ where $W^{+}$ is obtained if we first consider the graph ${W}\cup\boldsymbol{\boldsymbol{\bigcup}}\{C_{i}\mid i\in[b]\}$ and then add the vertex $a$ and all the edges between $a$ and $S.$ Clearly, $W^{+}$ may have vertices outside $\hat{c},$ however all vertices in $\{a\}\cup S$ are vertices of $\Sigma(\hat{c}).$ We now consider the graphs
  • $H_{1}=W^{+}\setminus\bigcup\{\Sigma(c_{i})\setminus V(C_{i})\mid i\in[b]\},$ and
  • $H_{2}=H_{1}\cup\bigcup\{\Sigma(c_{i})\mid i\in[b]\}.$
Observe that $H_{1}$ is a subgraph of $W^{+}$ and $W^{+}$ is a subgraph of $H_{2}.$ By construction, $H_{1}$ is a graph that is embedded in $\Sigma$ without crossings. Also keep in mind that ${\sf tw}(W^{+})\geq{\sf tw}(W)\geq z.$
Recall that every branch vertex of $W$ is at branch distance at most three from $a$. This fact combined with the construction of $H_{1}$ implies that every face or vertex of $H_{1}$ is at radial distance $\mathcal{O}(b)$ from the face of $H_{1}$ whose boundary of $C_{\hat{c}}.$ We obtain that the radius of $H_{1}$ is $\mathcal{O}(b)$ and this, combined with Subsection 6.1 implies that ${\sf tw}(H_{1})\in\mathcal{O}(bg).$
Our next step is to use the fact that ${\sf tw}(H_{1})\in\mathcal{O}(bg)$ in order to prove that ${\sf tw}(H_{2})\in\mathcal{O}(bgw).$ For this, let $(T,\beta)$ be a tree-decomposition of $H_{1}$ and let $L^{i}=\langle X_{1}^{i},X_{2}^{i},\dots,X_{n_{i}}^{i},x_{1}^{i},x_{2}^{i},\dots,x_{n_{i}}^{i}\rangle$ be a linear decomposition of $\Sigma(c_{i}),$ assuming that $x_{1}^{i},x_{2}^{i},\dots,x_{n_{i}}^{i}$ are the vertices of $\Omega_{i}$ ordered according to $\Omega_{i}.$ We now build a tree-decomposition $(T,\beta^{\prime})$ of $H_{2}$ where for each $t\in V(T),$ we define $\beta^{\prime}(t)$ to be the union of the set $\beta(t)$ and every set $X_{j}^{i},$ for which $x_{j}^{i}$ is a vertex of $\beta(t)\cap\boldsymbol{\boldsymbol{\bigcup}}\{\{x_{1}^{i},x_{2}^{i},\dots,x_{n_{i}}^{i}\}\mid i\in[b]\}.$ As each $X_{j}^{i}$ has $\mathcal{O}(w)$ vertices, it follows that $(T,\beta^{\prime})$ has width $\mathcal{O}(bgw),$ as required. We conclude that ${\sf tw}(H_{2})\in\mathcal{O}(bgw).$
Finally, recall that $W^{+}$ is a subgraph of $H_{2},$ therefore $z\leq{\sf tw}(W^{+})\leq{\sf tw}(H_{2})\in\mathcal{O}(bgw).$ In other words, ${\sf bg}(G,S)=c^{\prime}z^{2}\in\mathcal{O}(bgw)$ as required. ∎
Let $α$ be an integer such that for every graph $G,$ ${\sf tw}(G)\in\mathcal{O}({\sf bg}(G))^{\alpha}.$ Such a constant $α$ exists by Subsection 3.1. Then Subsection 6.2 implies the following.
[…]
For every graph $G,$ every $k\in\mathbb{N},$ and every surface $\Sigma$ with Euler-genus at most $g,$ if $A\subseteq V(G),$ $\delta$ is a $\Sigma$-decomposition of $G-A$ with width $\leq w$ and breadth $\leq b$ and ${\displaystyle X=A\cup\bigcup\{\Sigma(c)\mid\mbox{\small$c$ is a vortex of $\delta$}\}},$ then ${\sf tw}(G,X)\in\mathcal{O}((b^{4}\cdot(b\cdot g\cdot w)^{2}+|A|)^{α}),$ where $α$ is the constant of Subsection 3.1.
7 The lower bound
Using the concept of clique-sum closures, we can easily generate new graph parameters. For this, given a graph parameter ${\sf p}\colon\mathcal{G}_{\text{all}}\to\mathbb{N}$ we define the graph parameter ${\sf p}^{*}$ where for every graph $G,$ ${\sf p}^{*}(G)$ is the minimum $k$ such that $G$ is in the clique-sum closure of the class of graphs where the value of ${\sf p}$ is at most $k.$ Notice that ${\sf p}^{*}=({\sf p}^{*})^{*}.$ It is easy to see that if size is the parameter returning the number of vertices of a graph minus one, then ${\sf tw}=\textsf{size}^{*}.$
[…]
If ${\sf p}$ is a minor-monotone graph parameter, then ${\sf p}^{*}$ is also minor-monotone.
Given a graph parameter ${\sf p}\colon\mathcal{G}_{\text{all}}\to\mathbb{N}$ and a minor-closed graph class $\mathcal{G}$ we define the modulator composition of ${\sf p}$ and $\mathcal{G}$, as the parameter ${\sf p}_{\mathcal{G}}:\mathcal{G}_{\text{all}}\to\mathbb{N}$, so that
${\sf p}_{\mathcal{G}}(G)$$\coloneqq$ $\min\{k\mid\mbox{$\mbox{there exists an $X\subseteq V(G)$ such that\leavevmode\nobreak\ }{\sf p}(G,X)\leq k\mbox{\leavevmode\nobreak\ and\leavevmode\nobreak\ }G-X\in\mathcal{G}$}\}.$ (12)
Notice that, given that ${\sf p}$ is minor-monotone and $\mathcal{G}$ is minor closed, we have that ${\sf p}_{\mathcal{G}}$ is minor-monotone as well. Using (12), we define the “annotated modulation to a surface”-version of ${\sf p}$ as the graph parameter ${\sf p}_{\Sigma}$ where ${\sf p}_{\Sigma}\coloneqq{\sf p}_{\mathcal{E}_{\Sigma}}$. Notice that, because of Subsection 3.1, ${\sf tw}\sim_{\mathsf{P}}{\sf bg}.$ By the above definitions, one can easily deduce that ${\sf tw}_{\Sigma}\sim_{\mathsf{P}}{\sf bg}_{\Sigma}$ and moreover that
${\sf tw}_{\Sigma}^{*}$ $\sim_{\mathsf{P}}$ ${\sf bg}_{\Sigma}^{*}$ (13)
In (13), we agree that ${\sf tw}_{\Sigma}^{*}$ is a shortcut of $({\sf tw}_{\Sigma})^{*}$ and that ${\sf bg}_{\Sigma}^{*}$ is a shortcut of $({\sf bg}_{\Sigma})^{*}$. The equivalence in (13) will permit us to use ${\sf bg}_{\Sigma}$ instead of ${\sf tw}_{\Sigma},$ to prove our lower bounds.
In this section we consider for fixed values of $h$ and $c,$ the Dyck-grid $\mathscr{D}_{k}^{(h,c)}$ and a surface $\Sigma$ where $\mathscr{D}_{k}^{(h,c)}$ cannot be embedded (that is a surface $\Sigma$ not containing $\Sigma^{(h,c)}$). Our target is to prove that $\mathscr{D}_{k}^{(h,c)}$ cannot be the clique-sum closure of graphs with low-treewidth modulators to $\Sigma.$ In other words, we prove a (polynomial) lower bound to ${\sf tw}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})$ or, alternatively, because of (13), a lower bound to ${\sf bg}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})$ (in terms of $k$).
Our proof has two steps. In Subsection 7.1 we prove that ${\sf bg}_{\Sigma}(\mathscr{D}_{k}^{(h,c)})\in\Omega(k^{1/5}).$ Based on this result, in Subsection 7.2 we prove that ${\sf bg}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})\in\Omega(k^{1/480}).$ In our proofs we do not make any particular effort to optimize the parametric dependences. While we keep them being polynomial, it seems that much better polynomial bounds might be achieved with a more fine-grained analysis.
7.1 Modulators of Dyck-grids
In our proof we work with a simpler version of $\mathscr{D}^{(h,c)},$ where the annulus grid is suppressed. The new parametric graph is denoted by $\tilde{\mathscr{D}}^{(h,c)}=\langle\tilde{\mathscr{D}}_{k}^{(h,c)}\rangle_{k\in\mathbb{N}},$ where the graph $\tilde{\mathscr{D}}_{k}^{(h,c)}$ is obtained from $\mathscr{D}_{k}^{(h,c)}$ by removing the set $S$ of $4k^{2}$ vertices corresponding to the annulus grid $\mathscr{A}_{k}$ and, in the resulting graph, among the $2k$ neighbors of the vertices that have been removed, we join by edges the $k$ pairs that belong to the same cycle of $\mathscr{D}_{k}^{(h,c)}$ (see Figure 2 for an example). Clearly $\tilde{\mathscr{D}}_{k}^{(h,c)}\leq\mathscr{D}_{k}^{(h,c)}$ and it is easy to see that ${\mathscr{D}}_{k}^{(h,c)}\leq\tilde{\mathscr{D}}_{2k}^{(h,c)}.$ This implies the following observation.
[…]
For every $(h,c)\in\mathbb{N}\times[0,2]\setminus\{(0,0)\},$ $\mathscr{D}^{(h,c)}\sim_{{\mathsf{L}}}\tilde{\mathscr{D}}^{(h,c)}.$
[…]
Figure 14: The graph $\tilde{\mathscr{D}}_{8}^{1,2}$ and a half-integral packing of four subdivisions of $\tilde{\mathscr{D}}_{2}^{1,2},$ colored by different colors.
Let $G$ and $H$ be two graphs. A half-integral packing of $H$ in $G$ is a set $\mathcal{P}=\{H_{1},\ldots,H_{r}\}$ of subgraphs of $G$ such that
  • for every $i\in[r],$ $H\leq H_{i}$ and
  • no vertex of $G$ belongs to more than two graphs in $\mathcal{P}.$
We will use the following observation (we omit the tedious proof that becomes apparent when considering the half-integral packing of $4$ copies of $\mathscr{D}_{2}^{(h,c)}$ in $\mathscr{D}_{8}^{(h,c)}$ depicted in Figure 14).
[…]
Let $(h,c)\in\mathbb{N}\times[0,2]$ and let $x,y\in\mathbb{N}_{\geq 1}.$ Then $\tilde{\mathscr{D}}_{xy}^{(h,c)}$ contains a half-integral packing of $x$ copies of $\tilde{\mathscr{D}}_{y}^{(h,c)}.$
Let $h\in\mathbb{N}$ and $c\in[0,2]$ be integers where $(h,c)$ is different from $(0,0).$ We define $\mathbb{S}_{h,c}$ as the set containing every surface where it is not possible to embed all graphs in $\mathscr{D}^{(h,c)}.$ Notice that $\mathbb{S}_{h,c}$ contains every surface that does not contain $\Sigma^{(h,c)}.$ Due to Subsection 1.2, if $(h,c)\neq(0,0),$ then $\mathbb{S}_{h,c}$ is non-empty.
[…]
Let $(h,c)\in\mathbb{N}\times[0,2]\setminus\{(0,0)\}$ and let $\Sigma\in\mathbb{S}_{h,c}.$ Then ${\sf bg}_{\Sigma}(\mathscr{D}_{k}^{(h,c)})\in\Omega_{h,c}(k^{1/5}).$
Proof.
Notice that by Subsection 7.1 and the minor-monotonicity of ${\sf bg}_{\Sigma}$, it is enough to prove that $\mathsf{bg}_{\Sigma}(\tilde{\mathscr{D}}_{k}^{(h,c)})\in\Omega_{h,c}(k^{1/5}).$ Again by Subsection 7.1 we know that there is no surface in $\mathbb{S}_{h,c}$ where all graphs in $\tilde{\mathscr{D}}^{(h,c)}$ can be embedded. Let $k_{0}$ be minimum such that $\tilde{\mathscr{D}}^{(h,c)}_{k_{0}}$ is not embeddable in any of the (finitely many) surfaces in $\mathbb{S}_{h,c}.$ Clearly, $k_{0}$ depends exclusively on the choice of $h,c.$ Let $X$ be an modulator of $\mathscr{D}_{k}^{(h,c)}$ to the embeddability in $\Sigma.$
Our target is to prove that ${\sf bg}(\tilde{\mathscr{D}}^{(h,c)}_{k},X)\in\Omega_{h,c}(k^{1/5}).$
For each $b\in[0,k],$ we define $\mathcal{C}_{b}$ as the set of the $b$ outermost cycles of $\tilde{\mathscr{D}}_{k}^{(h,c)}$, and we define $B_{b}=V(\boldsymbol{\boldsymbol{\bigcup}}\mathcal{C}_{b})$ and $X_{b}=X\setminus B_{b}.$ The value of $b$ will be determined later.
Figure 15: A visualization of the first subcase of Case 1 in the proof of Subsection 7.1, where $k^{\prime}=10,$ $b=2,$ $m=12,$ and $r=5.$ The $b$ outermost cycles are colored green. We depict the partition of $G^{*}$ to four pairwise disjoint $(2k^{\prime}\times k^{\prime})$-grids, and the $(m\times m)$-grid $\Gamma_{m}$ as a minor in it. The perimeter of the central $(m-2b)\times(m-2b)$-subgrid is depicted red.
By Subsection 7.1, there is a set $\mathcal{P}_{b}$ of $\lfloor\frac{k-b}{k_{0}}\rfloor$ subgraphs of $\tilde{\mathscr{D}}^{(h,c)}_{k}$ that is a half-integral packing of $\tilde{\mathscr{D}}^{(h,c)}_{k_{0}}$ in $\tilde{\mathscr{D}}^{(h,c)}_{k}$ and such that the graphs in $\mathcal{P}_{b}$ do not contain vertices in $B_{b}.$ Notice that $X_{b}$ should contain at least one vertex from every graph in $\mathcal{P}_{b}$ and no vertex of $X_{b}$ can belong to more than two graphs in $\mathcal{P}_{b}.$ This implies that $|X_{b}|\geq\lfloor\frac{k-b}{2k_{0}}\rfloor.$ Observe now that the vertex set of $\tilde{\mathscr{D}}^{(h,c)}_{k}$ can be partitioned into $h+c$ vertex sets $V_{1},\ldots,V_{h+c}$ such that, for each $i\in[h+c],$ $\tilde{\mathscr{D}}^{(h,c)}_{k}$ contains an annotated graph $(G_{i},V_{i})$ as a $V_{i}$-minor where ${G}_{i}$ is the copy of either $\mathscr{H}_{k}$ or $\mathscr{C}_{k}$ and $V_{i}=V({G}_{i})$ (this partition is indicated in Figure 2). As before, we define $\mathcal{C}_{b}^{i}$ to be the set of the $b$ outermost cycles of $G_{i}$, and we define $B_{b}^{i}\coloneqq V(\boldsymbol{\boldsymbol{\bigcup}}\mathcal{C}_{b}^{i})$ and $X_{b}^{i}=X\setminus B_{b}^{i}.$ We now choose $i$ such that $|X_{b}^{i}|$ is maximized and this means that $|X_{b}^{i}|\geq\lfloor\frac{k-b}{2k_{0}(h+c)}\rfloor.$ We distinguish two cases, depending on whether ${G}_{i}=\mathscr{H}_{k}$ or ${G}_{i}=\mathscr{C}_{k}.$
Case 1. ${G}_{i}=\mathscr{H}_{k}.$ Let $k^{\prime}=\lfloor k/2\rfloor.$ Let $\overline{\mathcal{C}}_{b}^{i}$ be the cycles of $\mathscr{H}_{k}$ that are not in $\mathcal{C}_{b}^{i}$. As a first step, among the $k-b$ cycles of $\overline{\mathcal{C}}_{b}^{i}$ we consider the set $\mathcal{C}$ of $k^{\prime}-b$ such cycles whose union meets the biggest possible number of vertices in $X_{b}^{i}.$ Observe that the cycles of $\mathcal{C}$ meet at least $|X_{b}^{i}|(\frac{1}{2}-\frac{b}{2k-2b})$ vertices of $X_{b}^{i}.$ Let $X_{b}^{*i}=X_{b}^{i}\cap V(\boldsymbol{\boldsymbol{\bigcup}}\mathcal{C})$ and keep in mind that $|X_{b}^{*i}|\geq\lfloor(\frac{1}{2}-\frac{b}{2k-2b})\frac{k-b}{2k_{0}(h+c)}\rfloor\geq\lfloor\frac{k-2b}{4k_{0}(h+c)}-\frac{b}{4k_{0}(h+c)}\rfloor=\lfloor\frac{k-3b}{4k_{0}(h+c)}\rfloor.$
We now consider the minor $G^{*}_{i}$ of ${G}_{i}$ that is obtained if we contract every edge that is not in a handle or crosscap path, has endpoints in different cycles of $\overline{\mathcal{C}}_{b}^{i}$ and does do not have both endpoints in $V(\boldsymbol{\boldsymbol{\bigcup}}\mathcal{C})$ (we assume that, during contractions, vertices in $X^{*i}_{b}$ prevail). Notice that $G_{i}^{*}$ is a minor of $G$ containing the vertices of $X_{b}^{*i}$ (that are also vertices of $G_{i}$).
Notice now that $G^{*}_{i}$ contains a $(8k^{\prime}\times k^{\prime})$-cylindrical grid as a spanning subgraph that contains $k^{\prime}$ cycles and $8k^{\prime}$ paths that we call rails. Anti-diametrical pairs of rails can be seen as subpaths of paths of length $2k^{\prime}-1,$ we call them crossing paths, whose endpoints are in the outer cycle of $G^{*}.$ Notice that, the vertex set of $G^{*}$ is the union of two disjoint $2k^{\prime}\times 2k^{\prime}$-grids, say $\Gamma^{1}$ and $\Gamma^{2}.$ We call a crossing path of $G^{*}$ vertical if it belongs to $\Gamma^{1}$ and horizontal if it belongs to $\Gamma^{2}.$
Also, $G^{*}$ contains four pairwise disjoint $(2k^{\prime}\times k^{\prime})$-grids. We call them north, south, west, and east grid. (In Figures 15 and 16 each these grids is bicolored by red/blue frames.)
We now assume that the south grid has the majority of the vertices of $X_{b}^{*}.$ Let this vertex set be $\hat{X}_{b}^{*}$ and observe that $|\hat{X}_{b}^{*}|\geq\lfloor|{X}_{b}^{*}|/4\rfloor\geq\lfloor\frac{k-3b}{16k_{0}(h+c)}\rfloor.$ Next we observe that there is a subset $S_{b}$ of $\hat{X}_{b}^{*}$ and a set $\mathcal{Q}$ of $2b$ consecutive vertical crossing paths such that $|S_{b}|\geq\lfloor\frac{|\hat{X}_{b}^{*}|}{2b}\rfloor-1$ and $S_{b}\cap\boldsymbol{\boldsymbol{\bigcup}}\mathcal{Q}=\emptyset.$ W.l.o.g. we assume that the majority of the vertices of $S_{b}$ is “on the right” of the vertical crossing paths in $\mathcal{Q}$, and we denote these vertices by $S_{b}^{\prime}.$ Clearly $|S_{b}^{\prime}|\geq\lfloor|S_{b}|/2\rfloor\geq\lfloor\frac{k-3b}{64bk_{0}(h+c)}\rfloor-1.$
Figure 16: A visualization of the second subcase of Case 1 in the proof of Subsection 7.1, where $k^{\prime}=10,$ $b=2,$ $m=12,$ and $r=13.$ The $b$ outermost cycles are colored green. We depict the partition of $G^{*}$ to four pairwise disjoint $(2k^{\prime}\times k^{\prime})$-grids, and the $(m\times m)$-grid $\Gamma_{m}$ as a minor in it. The perimeter of the central $(m-2b)\times(m-2b)$-subgrid is depicted red.
Let $r$ be the number of consecutive rails in the south grid starting from the second half of the paths in $\mathcal{Q}$ (notice that $r\geq b+1$ as at least one rail should be intersected by $S_{b}^{\prime}$). In Figure 15 and Figure 16 we have that $r=5$ and $r=13$ respectively. Let $m=k^{\prime}+b.$ We will consider a subdivision of a $(m\times m)$-grid $\Gamma_{m}$ by considering rails of the south grid and, possibly, some of the east one. For this, we distinguish two subcases as follows:
  • If $r<k^{\prime}$ then we complete the $r$ rails with the $k^{\prime}-r+b$ lower rails in the east grid and construct an $(m\times m)$-grid $\Gamma_{m}$ (see Figure 15). Set $S_{b}^{\prime\prime}\coloneqq S_{b}^{\prime}.$
  • If $r\geq k,$ then among all $((k^{\prime}-b)\times k)$-grids that can be made using $k^{\prime}-b$ rails of the south grid and on the right of $\mathcal{Q},$ consider those that intersect the biggest number of vertices in $S_{b}^{\prime}.$ Also complete the $(m\times m)$-grid $\Gamma_{m}$ by adding $b$ rails from the rightmost grid (see Figure 16). Let $S_{b}^{\prime\prime}$ be the vertices of $S_{b}^{\prime}$ that belong to the $k^{\prime}-b$ chosen rails and observe that $|S_{b}^{\prime\prime}|\geq\lfloor|S_{b}^{\prime}|/2\rfloor.$
Notice that, in any case $|S_{b}^{\prime\prime}|\geq\lfloor|S_{b}^{\prime}|/2\rfloor\geq f(b,k)\coloneqq\lfloor\frac{k-3b}{128bk_{0}(h+c)}\rfloor-1$. In other words, the central $(m-2b)\times(m-2b)$-subgrid of this $(m\times m)$-grid contains at least $f(b,k)$ vertices whose precursors contain vertices from $X_{b}.$
Case 2. ${G}_{i}=\mathscr{C}_{k}.$ As we did before, we refer to the $4k$ paths of $\mathscr{C}_{k}$ as rails. Also, anti-diametrical pairs of rails are seen as subpaths of $2k$ paths of length $12k-$ 1 and, as before, we call them crossing paths (see Figure 17). Notice that the vertex set of $\mathscr{C}_{k}$ is the union of two disjoint $(2k\times k)$-grids, say $\Gamma^{1}$ and $\Gamma^{2}.$ We call a crossing path of $\mathscr{C}_{k}$ vertical if it belongs to $\Gamma^{1}$ and horizontal if it belongs to $\Gamma^{2}.$
Observe that there is a subset $S_{b}$ of $X_{b}^{i}$ and a set $\mathcal{Q}$ of $2b$ consecutive crossing paths such that $|S_{b}|\geq\lfloor\frac{|X_{b}^{i}|}{2b}\rfloor-1$ and $S_{b}\cap\boldsymbol{\boldsymbol{\bigcup}}\mathcal{Q}=\emptyset.$ We have that $|S_{b}|\geq f(b,k)\coloneqq\lfloor\frac{k-3b}{4bk_{0}(h+c)}\rfloor-1.$
Let $m=2k.$ Observe that $\mathscr{C}_{k}$ contain as a spanning subgraph an $(m\times m)$-grid $\Gamma_{m}.$ To see this, it is enough to remove all the edges between the vertices of the $b$-th and the vertices of the $(b+1)$-th path of $\mathcal{Q}$ from $\mathscr{C}_{k}$.
Figure 17: A visualization of Case 2 in the proof of Subsection 7.1, where $k=9,$ $b=2,$ and $m=18.$ The $b$ outermost cycles are colored green. We depict the partition of ${G}_{i}=\mathscr{C}_{k}$ to four pairwise disjoint $(k\times k)$-grids and the $(m\times m)$-grid as a spanning subgraph of $\mathscr{C}_{k}.$ The perimeter of the central $(m-2b)\times(m-2b)$-subgrid is depicted red.
Notice that the subgraph of $\Gamma_{m}$ induced by its $b$ outermost vertices is exactly $\Gamma_{m}\setminus(B_{b}\cup\boldsymbol{\boldsymbol{\bigcup}}\mathcal{Q}).$ Moreover, the central $(m-2b)\times(m-2b)$ subgrid of $\Gamma_{m}$ contains at least $f(b,k)$ vertices from $X_{b}.$
In both of the two cases above, we choose the margin $b_{k}$ as the maximum integer $b$ such that $4]{f(b,k)}\rfloor\geq 2b.$ This implies that |$4]{|S_{b_{k}}|}\rfloor\geq 2b_{k},$ therefore, by Subsection 6.2, ${\sf bg}({\tilde{\mathscr{D}}^{(h,c)}_{k}},X)\geq\lfloor\sqrt[4]{|S_{b_{k}}|}\rfloor\geq 2b_{k}.$ Observe now that in both cases $b_{k}\in Θ_{h,c}(k^{1/5}),$ therefore, by Subsection 6.2 ${\sf bg}({\tilde{\mathscr{D}}^{(h,c)}_{k}},X)\in\Omega_{h,c}(k^{1/5}).$
7.2 Lower bounds under the presence of clique-sums
It now remains to prove that, given that ${\sf bg}_{\Sigma}(\hat{\mathscr{D}}_{k}^{(h,c)})$ is big, then also ${\sf bg}_{\Sigma}^{*}(\hat{\mathscr{D}}_{k}^{(h,c)})$ is also big.
A partially triangulated $(k\times k)$-grid is any graph that can be obtained from a $(k\times r)$-grid (we call it the underlying grid) by adding edges in a way that the resulting graph remains planar. A dusted partially triangulated $(k\times k)$-grid is the disjoint union of a partially triangulated $(k\times k)$-grid and a set of isolated vertices.
The following lemma follows easily from the results of [31] (see also [18]).
[…]
There is a universal constant $c$ such that every graph $G$ were ${\sf bg}(G)\geq c\cdot\big{(}\big{(}{\sf hw}(G)\big{)}^{48}+r\big{)}$ contains a vertex set $A$ of size $c\cdot\big{(}{\sf hw}(G)\big{)}^{24}$ such that $G-A$ can be contracted to a dusted partially triangulated $(r\times r)$-grid.
For $k\in\mathbb{N}_{{\geq 1}},$ a graph is a crossed $k$-grid if it can be obtained as follows: first take a $(k+2)\times(k+2)$ grid, then remove all outermost layer edges (edges with both endpoints of degree $<4$), subdivide each of the remaining edges at least once and then take the line graph.
[…]
If $G$ has a tree-decomposition where all torsos have Hadwiger number at most $h,$ then also $G$ has Hadwiger number at most $h.$ In short, ${\sf hw}={\sf hw}^{*}.$
[…]
There is some universal constant $c$ such that if a graph $G$ contains a crossed $t$-grid, for some $t\in\mathbb{N}_{\geq 1},$ then $t\leq c\cdot({\sf hw}(H))^{2}.$
Proof.
If, in a crossed $2t(t-1)$-grid, we contract all edges that do not belong to $4$-cliques we obtain a graph that contains the grid-like graph $H^{1}_{t(t-1)}$ defined by Kawarabayashi, Thomas, and Wollan in [31] as a subgraph. According to [31, Lemma 3.2.], $H^{1}_{t(t-1)}$ contains $K_{t}$ as a minor. ∎
[…]
Let ${\cal G}$ be a graph class of Hadwiger number at most $η$. Then, for every graph $G$, ${\sf hw}(G)\leq({\sf bg}_{\mathcal{G}}(G))^{2}+η+1$.
Figure 18: A crossed $6$-grid. Subdivision vertices are depicted in yellow.
Proof.
Suppose that ${\sf hw}(G)=t.$
Let $X$ be a modulator of $G$ to $\mathcal{G}$. Notice that $G$ can be identified, via contractions, to the clique $K_{t}.$ Let $X^{\prime}$ be the heir of $X$ during these contractions. Notice that $X^{\prime}$ remains an modulator of $K_{t}$ to $\mathcal{G},$ therefore $|X^{\prime}|\geq t-η.$ Notice that ${\sf bg}(G,X)\geq{\sf bg}(K_{k},X^{\prime})$ and that ${\sf bg}(K_{t},X^{\prime})\geq{\sf bg}(K_{t}[X^{\prime}])\geq\lfloor|X^{\prime}|^{1/2}\rfloor\geq\lfloor(t-η)^{1/2}\rfloor\geq(t-η-1)^{1/2}.$ This implies that ${\sf bg}_{\mathcal{G}}(G)\geq(t-η-1)^{1/2}$, and the lemma follows. ∎
A new variant of the Dyck-grid.
For the purposes of our proofs, we will consider an “enhanced version” of $\tilde{\mathscr{D}}_{k}^{(h,c)},$ namely $\hat{\mathscr{D}}^{(h,c)}=\langle\hat{\mathscr{D}}_{k}^{(h,c)}\rangle$ which is obtained from $\tilde{\mathscr{D}}_{k}^{(h,c)}$ as follows: Consider a set of $k(h+c)$ pairwise disjoint paths on $4$ vertices of the outermost cycle. We take them so that the paths are parts of the the perimeters of the (pairwise disjoint) $2h$ $(2k\times k)$-grids and $c$ $(2k\times 2k)$-grids corresponding to the handles and the crosscaps of $\tilde{\mathscr{D}}_{k}^{(h,c)}$ (there is a unique way these paths can be considered). Then, for each of these paths, introduce a new vertex and make it adjacent with its four vertices. We call the $k(h+c)$ new vertices in this construction satellite vertices (see Figure 19).
We say that a parametric graph $\mathscr{H}=\langle\mathscr{H}_{k}\rangle_{k\in\mathbb{N}}$ has the $f$-tightly connected for some non-decreasing function $f:\mathbb{N}\to\mathbb{N}$, if for every separation $(B_{1},B_{2})$ of $\mathscr{H}_{k}$ of order $q<k$ such that both $G[B_{1}\setminus B_{2}]$ and $G[B_{2}\setminus B_{1}]$ are connected, it holds that one of $B_{1},$ $B_{2}$ has at most $f(q)$ vertices.
[…]
Let ${\sf p}$ be some minor-monotone graph parameter and let $h:\mathbb{N}\to\mathbb{N}$ such that for every graph $G$, ${\sf hw}(G)\leq h({\sf p}(G))$. Let also $\mathscr{H}=\langle\mathscr{H}_{k}\rangle_{k\in\mathbb{N}}$ be a parametric graph that is $g$-tightly connected, for some $g:\mathbb{N}\to\mathbb{N}$. Let $r\in\mathbb{N}$ and let ${(T,\beta)}$ be a tree decomposition of $G$ where, for each torso $G_{\bar{t}}$, it holds that ${\sf p}(G_{\bar{t}})\leq r$. If $2(g(h(r)))<|V(\mathscr{H}_{k})|$ and $h(r)<k$, then $\mathscr{H}_{k}$ admits a tree decomposition $(\hat{T},\hat{\beta})$ where $\hat{T}$ is a star with center $t$, where ${\sf p}(G_{t})\leq r$ ($G_{t}$ is the torso of $(\hat{T},\hat{\beta})$ on $t$) and where for every $e=tt^{\prime}\in E(\hat{T})$, $\mathscr{H}_{k}[\hat{β}(t^{\prime})\setminus\hat{β}(t)]$ is a connected graph and $|\hat{β}(t^{\prime})|\leq g(h(r))$.
Proof.
For simplicity we set $G=\mathscr{H}_{k}$. Let $e=tt^{\prime}\in E(T)$ and let $T_{t}$ and $T_{t^{\prime}}$ be the two connected components of $T-e,$ where $t\in V(T_{t})$ and $t^{\prime}\in V(T_{t^{\prime}}).$ We define $B_{t}\coloneqq\bigcup_{s\in V(T_{t})}β(s)$, $B_{t^{\prime}}\coloneqq\bigcup_{s\in V(T_{t^{\prime}})}β(s)$, and $A_{t,t^{\prime}}\coloneqq B_{t}\cap B_{t^{\prime}},$ i.e., $A_{t,t^{\prime}}$ is the common adhesion of $t$ and $t^{\prime}.$ We next modify $(T,β)$ as follows.
Claim. We can assume that, for every $e=tt^{\prime},$ every graph in $\{G[B_{t}\setminus B_{t^{\prime}}],G[B_{t^{\prime}}\setminus B_{t}]\}$ is void or connected.
Proof of Claim: Suppose that this is not the case and assume that for some edge $e=tt^{\prime}\in E(T),$ $G^{\prime}=G[B_{t}\setminus B_{t^{\prime}}]$ is not connected and has $r$ components $H_{1},\ldots,H_{r}.$ Let $Y$ be the component of $T-tt^{\prime}$ containing $t.$ We replace this component in $T$ with $r$ copies of $Y,$ say $Y_{1},\ldots,Y_{r},$ and for each one of them make the leaf $t^{i}$ corresponding to $t$ adjacent to $t^{\prime}.$ The new tree is denoted by $\hat{T}.$ We define $\hat{β}\colon V(\hat{T})\to 2^{V(G)}$ such that, if $s\not\in V(Y),$ then $\hat{β}(s)\coloneqq β(s),$ and if $s$ belongs to $Y_{i},i\in[r],$ then we set $\hat{β}(s)\coloneqq β(s)\cap V(H_{i}).$ It is easy to see that $(\hat{T},\hat{β})$ is a tree-decomposition of $G.$ Notice that each torso $G_{\hat{T}}^{*}$ of $(\hat{T},\hat{β})$ is the subgraph of some torso $G_{t}$ of $(T,β).$ This means that if ${\sf p}(G_{t})\leq r$ holds for each torso $G_{t}$ of $(T,β),$ then the same holds for the torsos of $(\hat{T},\hat{β})$. By repeating the above improvements as long as there are edges violating the statement of the claim, we finally have the required tree-decomposition.
Let $t\in V(T)$ and $t^{\prime}\in N_{T}(t)$. The set $A_{t,t^{\prime}}$ induces a clique in $G_{t}$, therefore, $|A_{t,t^{\prime}}|\leq{\sf hw}(G_{t})\leq h({\sf p}(G_{t}))\leq h(r)<k$. As $G$ is $g$-tightly connected, it holds hat $\min\{|B_{t}|,|B_{t^{\prime}}|\}\leq g(|A_{t,t^{\prime}}|)\leq g(h(r)),$ therefore $\max\{|B_{t}|,|B_{t^{\prime}}|\}\geq|G|-g(h(r))>g(h(r))\geq\min\{|B_{t}|,|B_{t^{\prime}}|\}.$ This permits us to orient every edge $tt^{\prime}$ towards $t,$ if $|B_{t}|>|B_{t^{\prime}}|$ and towards $t^{\prime},$ if $|B_{t^{\prime}}|>|B_{t}|.$ This orientation of $T$ implies that there is a unique node $t$ where all incident edges are oriented towards it. This means that, for every $t^{\prime}\in N_{T}(t)$, we have that $|B_{t^{\prime}}|\leq g(h(r)).$
We now build the claimed tree decomposition $(\hat{T},\hat{\beta})$ where $\hat{T}$ is the subtree of $T$ that is induced by $t$ an its neighbors $t^{\prime}$ in $T$ where $B_{t^{\prime}}\setminus A_{t,t}\neq\emptyset$. Finally, for each $t^{\prime}\in V(\hat{T}\setminus\{t\})$, we set $\hat{β}(t^{\prime})=B_{t^{\prime}}$. ∎
Figure 19: The graph $\hat{\mathscr{D}}_{8}^{1,2}$. The satellite vertices are depicted as squares. The magenta and the blue vertices correspond to the vertices in $A$ and $Y$ respectively, in the proof of Subsection 7.2.
[…]
Proof.
Let $G\coloneqq\hat{\mathscr{D}}_{k}^{(h,c)}.$ Let $(B_{1},B_{2})$ be a separation of $G$ of order $q<k.$ Let $A=B_{1}\cap B_{2},$ $D_{1}=B_{1}\setminus A,$ and $D_{2}=B_{2}\setminus A.$ Also we know that both $G[D_{1}]$ and $G[D_{2}]$ are connected. Let $Q$ be the spanning $(4k(h+c)\times k)$-annulus grid of $G.$ Clearly $Q$ contains $k$ cycles. Also it contains $4k(h+c)$ paths, each on $k$ vertices, which we call tracks. We also use $R$ for the set of the satellites of $\hat{\mathscr{D}}_{k}^{(h,c)}$. We enhance the tracks by extending each of them to the unique satellite vertex that is adjacent to one of its endpoints. Let $Y$ be the union of all the $\geq k-q$ cycles that are not met by $A$ and of all $\geq 4k(h+c)-4q$ tracks that are not met by $A$ (observe that if a vertex of $A$ belongs to $R$, then it meets four tracks). See Figure 19 for a visualization of $A$ and $Y$ in the graph $\hat{\mathscr{D}}_{8}^{1,2}$.
As every track in $Y$ has a common endpoint to every cycle in $Y,$ we obtain that $Y$ is connected, therefore $V(Y)$ is either a subset of $D_{1}$ or a subset of $D_{2}.$ W.l.o.g., we assume that $V(Y)\subseteq D_{1}.$ We next prove that $|B_{2}|\leq(2q+1)^{2}.$
Let $x$ be a vertex of $D_{2}$ and let $P_{x}$ be some path of $G[B_{2}]$ starting from $x$ and finishing to some vertex of $A$ and with all internal vertices in $D_{2}.$ This path cannot meet more than $q=|A|$ cycles of $Q$ because each such cycle contains some vertex of $S\subseteq D_{1}.$ Similarly, $P_{x}$ cannot meet more than $q=|A|$ tracks as each such track contains some vertex of $Y\subseteq D_{1}.$ This implies that every vertex of $G[B_{2}]$ should be at distance at most $q$ from $x.$ It is now easy to verify that, in $G,$ the vertices at distance at most $q$ from some $x\in A$ is upper bounded by $(2q+1)^{2}.$ As all vertices of $B_{2}$ are accessible from $x$ at this distance in $G[B_{2}],$ we conclude that $|B_{2}|\leq(2q+1)^{2}.$
[…]
Proof.
We need to introduce some constants (all depending on $\Delta$, $m$, $l$, and $r$). From now on, $c$ and $c^{\prime}$ are the universal constants of Subsection 7.2 and Subsection 7.2 respectively. We set
$\hat{h}$ $\coloneqq\max\{2l+1,c^{\prime}m^{2}+3\}\in\mathcal{O}(m^{2}),$
$d$ $\coloneqq(c^{\prime}m^{2}+3)^{4}\cdot(1+2l)^{2}+c\cdot m^{24}\in\mathcal{O}(m^{8}l^{2}+m^{24}),$
$t$ $\coloneqq d\cdot(1+4l)^{2}+\Delta c\cdot m^{24}\in\mathcal{O}(m^{8}l^{4}+\Delta m^{24}l^{2}),$
$a$ $\coloneqq\lceil{(t+1)^{1/2}}\rceil\cdot(r+1)+2l+2\hat{h}\in\mathcal{O}(r^{1/2}\cdot(m^{8}l^{4}+\Delta m^{24}l^{2})),\mbox{\leavevmode\nobreak\ and}$
$a^{\prime}$ $\coloneqq c\cdot(m^{48}+a)\in\mathcal{O}(m^{48}+r^{1/2}\cdot(m^{8}l^{4}+\Delta m^{24}l^{2}))\subseteq\text{$\mathsf{poly}$}(m+l+r+\Delta).$
Let $H$ be an $(X\cup B)$-minor of $G^{\mathsf{c}}$ such that ${\sf bg}(H)={\sf bg}(G^{\mathsf{c}},X\cup B)>a^{\prime}.$ We claim that it is enough to prove that ${\sf bg}(H-B)>r$. Indeed, as $H-B$ is an $X\setminus B$-minor of $G_{t}$, this implies ${\sf bg}(G_{t},X\setminus B)>r$ which implies ${\sf bg}(G_{t},X)\geq r$. We next prove that ${\sf bg}(H-B)>r$.
By Subsection 7.2, we have that $H$ contains a vertex set $\hat{A}$ where $|\hat{A}|\leq c\cdot m^{24}$ and such that $H-\hat{A}$ can be contracted to a dusted partially triangulated $(a\times a)$-grid $D_{a}.$ Keep in mind that the fact that $D_{a}$ is a contraction of $H-\hat{A}$ means that if two vertices of $D_{a}$ are at distance $\geq\delta$ in $D_{a}$, then their precursors in $H-\hat{A}$ are also at distance $\geq\delta$.
The fact that $H$ is an $X\cup B$-minor of $G^{\mathsf{c}}$ implies that the set of precursors of a vertex $v$ of $D_{a}$ in $H,$ and also in $G^{\mathsf{c}},$ contains some vertex of $X\cup B.$
Let $q=a-2\hat{h}$ and consider the central $(q\times q)$-grid $D_{q}$ of $D_{a}$ We say that a vertex $v$ of $D_{q}$ is marked by some vertex $t^{\prime}\in N\coloneqq V(T)\setminus\{t\},$ if its precursor in $H$, and also in $G^{\mathsf{c}}$, contains some vertex of $C\coloneqq B\setminus β(t)$, Notice that $v$ may be marked by more than one vertex $t^{\prime}$ of $N.$ We also define, for each $t^{\prime}\in N,$ we set $A_{t^{\prime}}=β(t)\cap β(t^{\prime})$.
As an important ingredient of the proof, we prove that there cannot be many marked vertices.
[…]
$D_{q}$ has at most $d$ marked vertices.
Proof of claim.
Let $\hat{M}$ be the set of marked vertices of $D_{q}$ and suppose to the contrary that
$|\hat{M}|>d=(c^{\prime}m^{2}+3)^{4}\cdot(1+4l)^{2}+c\cdot m^{24}.$
Notice that for at least $|\hat{M}|-c\cdot m^{24}$ of the vertices in $\hat{M},$ the set $C_{t^{\prime}}\coloneqq β(t^{\prime})\setminus β(t)$ that contains their precursor can be chosen so that it does not intersect $\hat{A}$ and we call them strongly marked. Let $M$ be the set of strongly marked vertices of $D_{q}$ and keep in mind that $|{M}|>(c^{\prime}m^{2}+3)^{4}\cdot(1+4l)^{2}.$ We say that two strongly marked vertices $v_{1}$ and $v_{2}$ are independent if there is no $t^{\prime}\in N$ such that both $v_{1}$ and $v_{2}$ are marked by $t^{\prime}.$ We know that each $C_{t^{\prime}}$ has at most $|β(t^{\prime})|\leq l$ vertices. Therefore, if $v_{1}$ (resp. $v_{2}$) is strongly marked by $t^{\prime}_{1}$ (resp. $t^{\prime}_{2}$) and $\mathsf{dist}_{D_{a}}(v_{1},v_{2})\geq 2l+1$ then $C_{t_{1}^{\prime}}\cap C_{t_{2}^{\prime}}=\emptyset$ which, in turn, implies that $v_{1}$ and $v_{2}$ are independent.
We now construct a set $M^{\prime}\subseteq M$ of pairwise-independent strongly marked vertices of $D_{q}$ by greedily picking from $M$ a strongly marked vertex $x$ and then excluding as possible choices all vertices of $M$ that are at distance $<2l+1$ from $x$ in $D_{a}$.
Recall that $2l+1\leq\hat{h},$ therefore, a vertex in the partially triangulated grid $D_{q}$ has at most $(1+4l)^{2}$ vertices at distance $\leq 2l$ in $D_{a}.$ We obtain that $|M^{\prime}|\geq|M|/(1+4l)^{2}\geq(c^{\prime}m^{2}+3)^{4}.$ Let $\ell\coloneqq c^{\prime}m^{2}+3$ and recall that $\ell\leq\hat{h}.$
We have just found a set of $\ell^{4}$ vertices of $M^{\prime}$ that belong to the central $(a-2\ell)\times(a-2\ell)$-grid of $D_{a}.$ By Subsection 6.2, $D_{a}$ contains an $(\ell\times\ell)$-grid $\Gamma_{\ell}$ as an $M^{\prime}$-minor. This also implies that $H,$ and also $G^{\mathsf{c}},$ contains $\Gamma_{\ell}$ as a minor. For every vertex $u$ of $\Gamma_{\ell}$ we know that the set of its precursors in $H$, and also in $G^{\mathsf{c}}$, contains some vertex $z_{u}$ of some set $C_{t^{\prime}_{u}}$ that does not intersect $\hat{A}.$ As there may be many possible choices of $z_{u},$ from now on, we arbitrarily fix one so to uniquely determine $t^{\prime}_{u}$ and $C_{t^{\prime}_{u}}$ as well.
By the way we defined $M^{\prime},$ it holds that, for distinct $u$’s in $\Gamma_{\ell},$ the corresponding $C_{t^{\prime}_{u}}$’s are pairwise disjoint and moreover, the corresponding $t^{\prime}_{u}$’s as well as the corresponding $z_{u}$’s are also pairwise distinct. Let $u\in V(\Gamma_{\ell}).$ As $G$, and therefore also $G^{\mathsf{c}}$, is 4-connected, we have that $|A_{t^{\prime}_{u}}|\geq 4$. Moreover, in $G^{\mathsf{c}}$, there are four internally disjoint paths in $G[β(t^{\prime}_{u})]$ starting from $z_{u}\in C_{t^{\prime}_{u}},$ finishing at any possible choice of four distinct vertices of $A_{t^{\prime}_{u}}$ and such that the internal vertices of these paths all belong to $C_{t^{\prime}_{u}}.$
Figure 20: A visualization of the grids $D_{a}$ and $D_{q}$ in the proof of Subsection 7.2. In red we depict the vertices $z_{t^{\prime}_{u}}$ of the of the $(\ell\times\ell)$-grid $\Gamma_{\ell}$ (the corners are squares). In blue we depict the subdivided paths of $\Gamma_{\ell}$ that are also paths of $G^{\mathsf{c}}.$ The corresponding $C_{t^{\prime}_{u}}$’s (resp. $A_{t^{\prime}}$’s) are in yellow (resp. green).
The above imply that the $4$-connected graph $G^{\mathsf{c}}$ contains as a subgraph a subdivision $L$ of the $(\ell\times\ell)$-grid where the vertices of degree four are all vertices of $Z\coloneqq\{z_{t^{\prime}_{u}}\mid u\in V(\Gamma_{\ell})\}$ (see Figure 20). Let $x\coloneqq z_{t^{\prime}_{u}}\in Z$ and assume that, in the subdivided grid $L,$ $x$ is the intersection of the “horizontal” path $H_{x}$ and the “vertical” path $V_{x}.$
We traverse $H_{x}$ from left to right and we define $x^{\textrm{left}}$ to be the first vertex of $A_{t^{\prime}_{u}}$ appearing on $H_{x}$ and $x^{\textrm{right}}$ to be the last vertex of $A_{t^{\prime}_{u}}$ appearing on $H_{x}.$ Similarly, we traverse $V_{x}$ from up to down and we define $x^{\textrm{up}}$ to be the first vertex of $A_{t^{\prime}_{u}}$ appearing on $V_{x}$ and $x^{\textrm{down}}$ to be the last vertex of $A_{t^{\prime}_{u}}$ appearing on $V_{x}.$ We now construct a subgraph $L^{\prime}$ of $L$ by removing, for each $x=z_{t^{\prime}_{u}}\in Z,$ all internal vertices of the path between $x^{\textrm{left}}$ and $x^{\textrm{right}}$ and all internal vertices of path between $x^{\textrm{up}}$ and $x^{\textrm{down}}.$
[…]
Figure 21: A visualization of the last part of the proof of Subsection 7.2. The blue lines indicate the packing of the $t+1$ pairwise vertex-disjoint $(r+1,r+1)$-grids in $D_{q^{\prime}}$
We set $q^{\prime}\coloneqq q-2l\geq a-2\hat{h}-2l=\lceil{(r+1)^{1/2}}\rceil\cdot(t+1)$, and we denote by $D_{q^{\prime}}$ the central $(q^{\prime}\times q^{\prime})$-grid $D_{q^{\prime}}$ of $D_{q}.$ Let ${B}^{(q^{\prime})}$ be the vertices of $V(D_{q^{\prime}})$ whose set of precursors in $G^{\mathsf{c}}$ contains a vertex of some $β(t^{\prime}),t^{\prime}\in N$ where $β(t^{\prime})\cap\hat{A}=\emptyset$.
Observe that, because $|β(t^{\prime})|\leq l$ and $G[B_{t^{\prime}}]$ is connected, every vertex $u\in B^{(q^{\prime})}$ should be at distance at most $l$ from some of the $\leq d$ marked vertices of $D_{q}.$ As a consequence, $|B^{(q^{\prime})}|\leq d\cdot(1+2l)^{2}.$ Let now $\tilde{B}^{(q^{\prime})}$ be the vertices of $V(D_{q^{\prime}})$ whose precursors in $H$ contain a vertex of some $B_{t^{\prime}}$ that, moreover, contains some vertex of $\hat{A}.$ As all vertices in $\hat{A}$ have degree at most $\Delta$, each of them may be adjacent with at most $\Delta$ vertices of some of the $B_{t^{\prime}}$’s. This means that $|\tilde{B}^{(q^{\prime})}|\leq\Delta\cdot|\hat{A}|\leq\Delta c\cdot m^{24}.$ We stress that ${B}^{(q^{\prime})}$ and $\tilde{B}^{(q^{\prime})}$ are not necessarily disjoint, however we just proved that $|{B}^{(q^{\prime})}\cup\tilde{B}^{(q^{\prime})}|\leq d\cdot(1+2l)^{2}+\Delta c\cdot m^{24}=t.$
Notice that, by their definition, the vertices in $V(D_{q^{\prime}})\setminus({B}^{(q^{\prime})}\cup\tilde{B}^{(q^{\prime})})$ have precursors in $H$ that are all vertices of $H-B.$
As $q^{\prime}\geq\lceil{(t+1)^{1/2}}\rceil\cdot(r+1),$ we may consider a packing of $t+1$ pairwise vertex-disjoint $((r+1)\times(r+1))$-grids of $D_{q^{\prime}}$. Clearly, for some, say $\Gamma_{r},$ of these $((r+1)\times(r+1))$-grids, it holds that the precursors in $H$ of its vertices are all vertices outside $B$ (see Figure 21). This implies that ${\sf bg}(H-B)\geq r+1$, as required. ∎
We are now ready to prove the main result of this section, that is the following.
[…]
Let $(h,c)\in\mathbb{N}\times[0,2]\setminus\{(0,0)\}$ and let $\Sigma\in\mathbb{S}_{h,c}.$ Then ${\sf bg}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})\in\Omega_{h,c}(k^{1/480}).$
Proof.
Let $η=\max\{{\sf hw}(G)\mid G\in\mathcal{E}_{\Sigma}\}.$ Notice that $\tilde{\mathscr{D}}_{k}^{(h,c)}\leq\hat{\mathscr{D}}_{k}^{(h,c)}\leq\tilde{\mathscr{D}}_{k+1}^{(h,c)},$ therefore $\tilde{\mathscr{D}}_{k}^{(h,c)}\sim_{{\mathsf{L}}}\hat{\mathscr{D}}_{k}^{(h,c)}.$ By Subsection 7.1, we obtain that ${\mathscr{D}}_{k}^{(h,c)}\sim_{{\mathsf{L}}}\hat{\mathscr{D}}_{k}^{(h,c)}.$ By Subsection 7.1 and the fact that ${\mathscr{D}}_{k}^{(h,c)}\sim_{{\mathsf{L}}}\hat{\mathscr{D}}_{k}^{(h,c)},$ it is enough to prove that
${\sf bg}_{\Sigma}(\hat{\mathscr{D}}_{k}^{(h,c)})\in\mathcal{O}_{h,c}\big{(}{\sf bg}_{\Sigma}^{*}(\hat{\mathscr{D}}_{k}^{(h,c)})^{96}\big{)}.$ (14)
We prefer to work with $\hat{\mathscr{D}}_{k}^{(h,c)}$ (depicted in Figure 19), instead of $\tilde{\mathscr{D}}_{k+1}^{(h,c)}$ or ${\mathscr{D}}_{k+1}^{(h,c)}$, because it is $4$-connected.
For simplicity, we set $G\coloneqq\hat{\mathscr{D}}_{k}^{(h,c)}$, and we assume that ${\sf bg}_{\Sigma}^{*}(\hat{\mathscr{D}}_{k}^{(h,c)})\leq r,$ therefore there exists a tree-decomposition ${(T,\beta)}$ of $G$ such that, for every $t\in V(T)$,
$\max\{{{\sf bg}_{\Sigma}(G_{t})\mid t\in V(T)}\}$$\leq$ $r.$ (15)
Our objective is to prove that ${\sf bg}_{\Sigma}(G)\in\mathcal{O}_{h,c}(r^{96})$. We set
$m$ $=$ $r^{2}+η+1$ (16)
and we may assume that $2(2m+1)^{2}<4k^{2}$, otherwise ${\sf bg}_{\Sigma}(G)\leq|V(G)|=4k^{2}(h+c)\leq 2(2m+1)^{2}(h+c)$$\in\mathcal{O}_{h,c}(r^{4})\subseteq\mathcal{O}_{h,c}(r^{96})$, and we are done. This in turn implies that
$2(2m+1)^{2}$ $<$ $4k^{2}({h}+c),\mbox{\leavevmode\nobreak\ and}$ (17)
$m$ $<$ $k$ (18)
Notice that the assumptions of Subsection 7.2 hold for ${\sf p}={\sf bg}_{\Sigma}$ (for the function $h(x)=x^{2}+η+1$, because of Subsection 7.2) and for $\mathscr{H}=\hat{\mathscr{D}}^{(h,c)}$ for the function $g(x)=(2x+1)^{2}$ (because of Subsection 7.2). Moreover, given (15) and (16), the conditions $2(g(h(r)))<|V(\mathscr{H}_{k})|$ and $h(r)<k$ are satisfied, because of (17) and (18) respectively. Therefore, $G$ has a tree decomposition $(\hat{T},\hat{\beta})$ where $\hat{T}$ is a star with center $t$. Let $G_{t}$ be the torso of $(\hat{T},\hat{\beta})$ on $t$. From Subsection 7.2, it holds that ${\sf bg}_{\Sigma}(G_{t})\leq r$, therefore there is some $X\subseteq V(G_{t})$, such that
${\sf bg}_{\Sigma}(G_{t},X)\leq r\mbox{\leavevmode\nobreak\ and\leavevmode\nobreak\ }$ (19)
$G_{t}-X\in\mathcal{E}_{\Sigma}$ (20)
Moreover, for every $e=tt^{\prime}\in E(\hat{T})$, $\mathscr{H}_{k}[\hat{β}(t^{\prime})\setminus\hat{β}(t)]$ is a connected graph and $|\hat{β}(t^{\prime})|\leq g(h(r))=(2m+1)^{2}$.
For all $t^{\prime}\in N\coloneqq N_{\hat{T}}(t),$ we define $A_{t^{\prime}}=\hat{β}(t)\cap\hat{β}(t^{\prime})$, $B_{t^{\prime}}=\hat{β}(t^{\prime})$, and $C_{t^{\prime}}=B_{t^{\prime}}\setminus A_{t^{\prime}}$ and keep in mind that each $C_{t^{\prime}}$ induces in $G$ a connected graph where
$|C_{t^{\prime}}|\leq|B_{t^{\prime}}|\leq(2m+1)^{2}.$ (21)
Let $G^{\mathsf{c}}\coloneqq\bigcup_{t\in V(T)}G_{t},$ $B\coloneqq\bigcup_{t^{\prime}\in N}B_{t^{\prime}},$ and $C\coloneqq\bigcup_{t^{\prime}\in N}C_{t^{\prime}}.$ By Subsection 7.2 and Subsection 7.2, we have
${\sf hw}(G^{\mathsf{c}})$$\leq$ $m.$ (22)
[…]
For this notice that Subsection 7.2 applies for $\Delta=4$, $(T,\beta)=(\hat{T},\hat{\beta})$, and $l=(2m+1)^{2}\in\mathcal{O}(m^{2})$. As $m\in\mathcal{O}_{h,c}(r^{2})$, Subsection 7.2 implies that ${\sf bg}(G^{c},X\cup B)\in\mathcal{O}_{h,c}(r^{96})$. As we proved (14), this completes the proof of the lemma. ∎
We are now in position to prove the lower bound in Theorem 1.2.
[…]
For every non-empty finite and closed set of surfaces $\mathbb{S},$ $\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}\preceq\mathbb{S}\mbox{-}\mathsf{{\sf tw}}.$
Proof.
Recall that, $\mathsf{sobs}(\{\Sigma^{\varnothing}\})=\{\Sigma^{(0,0)}\}$ and, by definition, $\{\Sigma^{\varnothing}\}\mbox{-}{\sf tw}\sim{\sf tw}$ and $\{\Sigma^{(0,0)}\}\mbox{-}{\sf bg}\sim{\sf bg}.$ Therefore, we may examine the case where $\mathsf{sobs}(\mathbb{S})$ contains two surfaces $\Sigma^{(h_{i},c_{i})},i\in[2],$ different from the sphere: one orientable (assuming $h_{1}>0$ and $c_{1}=0$) and one non-orientable (assuming $h_{1}>h_{2}\geq 0$ and $c_{2}>0$). Recall also that, by the definition of $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}$ in (6) and ${\sf tw}_{\Sigma}^{*}$ in (12), it holds that for every graph $G,$
$\mathbb{S}\mbox{-}\mathsf{{\sf tw}}(G)$ $=$ $\min\{{\sf tw}_{\Sigma}^{*}\mid\Sigma\in\mathbb{S}\}.$ (23)
Let $G$ be a graph where $\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}(G)\geq k.$ From the definition of $\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}$ in (7), this means that $G$ contains the graph $\mathscr{D}_{k}^{\Sigma^{(h_{i},c_{i}})}$ as a minor for some $i\in[2].$ Notice that the fact that $\Sigma^{(h_{i},c_{i})}\in\mathsf{sobs}(\mathbb{S})$ implies that $\mathbb{S}\subseteq\mathbb{S}_{h_{i},c_{i}}$. Indeed, $\mathbb{S}\subseteq\mathbb{S}_{h_{i},c_{i}}$ follows by Subsection 1.2 and the definition of $\mathbb{S}_{h,c}$ (that is the set containing every surface where it is not possible to embed all graphs in $\mathscr{D}^{(h,c)}$).
By Subsection 7.2, for every $\Sigma\in\mathbb{S}_{h,c}$ it holds that ${\sf tw}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})\in\Omega(k^{1/480}),$ therefore for every $\Sigma\in\mathbb{S}$ it holds that ${\sf tw}_{\Sigma}^{*}(\mathscr{D}_{k}^{(h,c)})\in\Omega_{c,h}(k^{1/480}),$ which, by (23) and Section 7, implies that $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}(G)\in\Omega(k^{1/480})$ and we are done.∎
8 Putting everything together
We define a more refined version of ${\mathbb{S}}\mbox{-}\mathsf{{\sf hw}},$ namely ${\mathbb{S}}\mbox{-}\mathsf{\mathsf{rhw}},$ where:
${\mathbb{S}}\mbox{-}\mathsf{\mathsf{rhw}}(G)$ is the minimum $k$ for which $G$ has a tree-decomposition $(T,\beta)$ of adhesion at most $k$ such that for every $d\in V(T),$
  • either $|\beta(d)|\leq k,$ or
  • there exist a set $A_{d}\subseteq\beta(d)$ of size at most $k$ and a surface $\Sigma\in\mathbb{S}$ such that $\mathsf{Torso}_{G}(d)-A_{d}$ has a $\Sigma$-decomposition $\delta$ of width at most $k$ and breadth at most $k,$ and
[…]
    • i.
      $|(\beta(d)\cap\beta(d^{\prime}))\setminus X_{d}|\leq 3$ and
    • ii.
      $\mathsf{Torso}_{G}(d)-X_{d}$ is a minor of $G.$
Notice that the main difference of the above definition with the one of $\mathbb{S}\mbox{-}{\sf hw}$ in (10) are the extra conditions i. and ii. Condition i. says that the “surface part” of each adhesion should not have more than 3 vertices and Condition ii. says that the “surface part” part of each torso is a minor of the graph. We stress that these two conditions are important for algorithmic applications of the MGST. Our next result implies that no matter if we include or remove these two conditions, we define asymptotically equivalent parameters.
We now give the following most extended version of Theorem 1.5.
Theorem 8.1.
For every finite and closed set of surfaces $\mathbb{S},$ it holds that the parameters ${\mathbb{S}}\mbox{-}\mathsf{{\sf tw}},$ ${\mathbb{S}}\mbox{-}\mathsf{{\sf hw}},$ ${\mathbb{S}}\mbox{-}\mathsf{\mathsf{rhw}},$ and ${\mathsf{sobs}(\mathbb{S})}\mbox{-}{\sf bg}$ are asymptotically equivalent, i.e., $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}\sim{\mathbb{S}\mbox{-}\mathsf{{\sf hw}}}\sim{\mathbb{S}}\mbox{-}\mathsf{\mathsf{rhw}}\sim{\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}.$
Proof.
Recall that $\mathsf{sobs}(\emptyset)=\{\Sigma^{\varnothing}\}$ and, by definition, all values of -$\emptyset\mbox{-}\mathsf{{\sf tw}},$$\emptyset\mbox{-}\mathsf{{\sf hw}},$$\emptyset\mbox{-}\mathsf{\mathsf{rhw}},$ and $\{\Sigma^{\varnothing}\}\mbox{-}{\sf bg}$ are equal to $\infty$.∞ . Therefore, we may assume that $\mathbb{S}$ is non-empty.
We prove (in order) ${\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}\preceq\mathbb{S}\mbox{-}\mathsf{{\sf tw}}\preceq\mathbb{S}\mbox{-}\mathsf{{\sf hw}}\preceq\mathbb{S}\mbox{-}\mathsf{\mathsf{rhw}}\preceq{\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}.$
Proof of ${\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}\preceq\mathbb{S}\mbox{-}\mathsf{{\sf tw}}.$ This is Subsection 7.2.
Proof of $\mathbb{S}\mbox{-}\mathsf{{\sf tw}}\preceq\mathbb{S}\mbox{-}\mathsf{\text{${\sf hw}$}}.$ This follows from Subsection 6.2, the definition of ${\mathbb{S}}\mbox{-}{\sf tw}$ in (6), and the definition of ${\mathbb{S}}\mbox{-}{\sf hw}$ in (10).
Proof of $\mathbb{S}\mbox{-}\mathsf{{\sf hw}}\preceq\mathbb{S}\mbox{-}\mathsf{\mathsf{rhw}}.$ This follows directly because the definition of $\mathbb{S}\mbox{-}\mathsf{{\sf hw}}$ is the definition of $\mathbb{S}\mbox{-}\mathsf{\mathsf{rhw}}$ without the additional conditions i. and ii.
Proof of $\mathbb{S}\mbox{-}\mathsf{\mathsf{rhw}}\preceq{\mathsf{sobs}(\mathbb{S})\mbox{-}{\sf bg}}.$ This is Theorem 5.2. ∎
Let us now present the general statement of the main structural result of the Graph Minors Series.
[…]
There exists a function $f:\mathbb{N}\to\mathbb{N}$ such that for every graph $G,$ if ${\sf hw}(G)\leq k,$ then $G$ has a tree-decomposition ${(T,\beta)}$ such that for all ${t\in V(T)},$ there exist a set $A_{t}\subseteq\beta(t)$ of size at most $f(k)$ such that there exists a surface of Euler-genus at most $f(k)$ and a $\Sigma$-decomposition $\delta$ of $\mathsf{Torso}_{G}(d)-A_{d}$ with breadth at most $f(k)$ and width at most $f(k).$
A more direct way to state the above theorem is to say that ${\sf hw}^{\prime\prime}\preceq{\sf hw},$ where ${\sf hw}^{\prime\prime}$ is as defined in (9) (Subsection 1.3).
Let us now finish this section by going back in the parameter ${\sf hw}^{\prime}$ defined in the introduction ((2) in Subsection 1.1). Given Section 8, it appears that Subsection 6.2 is the only essential ingredient for proving the asymptotic equivalence of ${\sf hw}^{\prime}$ with the Hadwiger number ${\sf hw}.$
Theorem 8.2.
${\sf hw}\sim{\sf hw}^{\prime}\sim{\sf hw}^{\prime\prime}.$
Proof.
As we already mentioned, ${\sf hw}^{\prime\prime}\preceq{\sf hw}$ is a restatement of Section 8. The inequality ${\sf hw}^{\prime}\preceq{\sf hw}^{\prime\prime}$ follows by seeing (using Subsection 6.2) the set $A_{d}$ along with the vertices of the vortices of $\delta$ as a low treewidth modulator. Finally, ${\sf hw}\preceq{\sf hw}^{\prime}$ follows because, for every $X\subseteq V(K_{k}),$ it holds that ${\sf bg}(G,X)\geq\lfloor\sqrt{|X|}\rfloor\in\Omega(\sqrt{k})$ and that ${\sf eg}(K_{k-|X|})\in\Omega(\sqrt{k-|X|}).$
9 Discussion and open problems
In this paper we defined a hierarchy of parameters that bijectively reflects the containment-relation hierarchy of all two-dimensional surfaces without boundaries. This permitted us to reveal the (asymptotic) structure of graphs excluding graphs of bounded Euler-genus. Also, we introduced a compact way to describe this structure using low-treewidth modulators.
Recall that the parameters we consider in this paper are defined via low-treewidth modulators. What if we do not have any modulation at all? Such a parameter would quantify the clique-sum closures of graphs that have either bounded size or fixed Euler-genus. To derive a finite set of universal obstructions for this is a interesting, however highly non-trivial, challenge.
Acknowledgements:
We are thankful to Laure Morelle, Christophe Paul, Evangelos Protopapas, and Giannos Stamoulis, for their helpful comments on the containments and the presentation of this paper.